### v2

# Orbital orders under magnetic fields in cubic PrIr2Zn20

Abstract
We study orbital orders in quadrupolar heavy-fermion compound PrIr2Zn20 under magnetic fields on the basis of the Landau theory. Assuming Eg orbital (electric quadrupolar) orders in the cubic symmetry with the ordering wavenumber at the L points in the face-centered cubic lattice Brillouin zone as observed experimentally, we construct the Landau free energy and analyze it. We find that the unidentified high-temperature ordered phase under the magnetic field $\bm{H}\parallel[001]$ reported earlier can be interpreted as the consequence of the internal rotation in the orbital moments of f electrons at the Pr site. We discuss the phase diagram for various magnetic-field directions and also possible double-$\bm{q}$ quadrupolar orders in this system.
pacs:
Valid PACS appear here
I Introduction
Strongly correlated electron systems with orbital moments have attracted significant attention in recent years. Tokura and Nagaosa (2000) Various multipolar orders appear in d- and f-electron-based materials and have been intensively studiedMydosh and Oppeneer (2014); Kuramoto et al. (2009); Chen et al. (2010). Among them, electric quadrupole orders are the most typical and extensively investigatedShiina et al. (1997); Onimaru et al. (2005); Chubukov et al. (2016). In addition, spin nematic orders have been also studied as well.Andreev and Grishchuk (1984); Läuchli et al. (2006); Tsunetsugu and Arikawa (2006); Jiang et al. (2023) The nature of orbital degrees of freedom leads to spatially anisotropic interactions, making the symmetry in orbital space generally anisotropic. Each type of orbital order has unique properties, and understanding these properties remains a key issue in condensed matter theory and experiment.
[…]
Among the 1-2-20 compounds, PrIr2Zn20 is a prototypical material, exhibiting quadrupole ordering below 1 K and undergoing a phase transition to a superconducting state at 0.1 KOnimaru et al. (2011); Ishii et al. (2011); Onimaru et al. (2016). The early neutron scattering experimentIwasa et al. (2017) under an applied magnetic field $\bm{H}\parallel[\bar{1}10]$ suggested an ordering vector of $\bm{q}=(\pi,\pi,\pi)$, corresponding to the L point at the first Brillouin zone boundary. Here, we have set the lattice constant to unity. The analysis of induced magnetic moments indicates an orbital order of the $x^{2}-y^{2}$ type for $\bm{H}\parallel[\bar{1}10]$. Recent reinvestigations of the temperature–magnetic field phase diagram reveal strong anisotropy and highlight the existence of the high-temperature phaseKittaka et al. (2024) reported in earlier studiesOnimaru et al. (2011) across a wide range of magnetic field directions from [001] to [111], while it is absent for $\bm{H}\parallel[110]$.
In this paper, we analyze a phenomenological Landau model for PrIr2Zn20, incorporating $E_{g}$ quadrupole moments at the L points. We elucidate the nature of $E_{g}$ orbital ordering under magnetic fields and its dependence on field direction. We examine the order parameter suggested by the neutron scattering experiments for $\bm{H}\parallel[\bar{1}10]$, construct phase diagrams under applied magnetic fields, and compare them with experimental results. The ordering vectors at the L points reveal unique multiple-$\bm{q}$ orders, which differs from the previous studies about quadrupole orders at the X point ordering vectors in the face-center cubic lattice.Tsunetsugu et al. (2021); Hattori et al. (2023)
This paper is organized as follows. In Sec. II, we construct the Landau free energy for the $E_{g}$ quadrupole moments at wave vectors $\bm{q}=(\pi,\pi,\pi)$ and $(0,0,0)$. In Sec. III, we present the results of the Landau analysis for single-$q$ orbital orders, outlining the fundamental properties of the system. We then discuss multiple-$q$ orbital orders and the stability of various ordered phases in Sec. IV. In Sec. V, we compare our results with experimental data for PrIr2Zn20 and related materials. Finally, Sec. VI provides a summary of this paper.
II Model
We start to introduce the exchange interactions between the Eg orbital moments in Sec. II.1 and the Landau free energy for the order parameters at the L points in Sec. II.2. Mode coupling terms related to the order parameter at the L points and other degrees of freedom are discussed in Sec. II.5
Figure 1: Fourier transform of the exchange interaction for $J_{1}=1$ and (a) $(J_{2},J_{3},J_{4})=(0.0,0.0,0.0)$, (b) $(0.0,0.0,1.0)$, and (c) $(0.5,0.3,0.4)$. (d) The first Brillouin zone of the fcc lattice. The four L points are indicated by filled red circles and other symmetric points are indicated by filled black circles.
II.1 Exchange interaction
Let us denote Eg quadrupole moments at the position $\bm{r}=(x,y,z)$ in the diamond structure as $\bm{\phi}_{\bm{r}}^{\alpha}=(u_{\bm{r}}^{\alpha},v_{\bm{r}}^{\alpha})$, where $\alpha=$A or B is the sublattice index. We take one of the A sublattice site at the origin $\bm{r}=(0,0,0)$ and its pair B sublattice site in the unit cell at $\bm{r}+(1,1,1)/4$. $\bm{\phi}_{\bm{r}}^{\alpha}$ form a two-dimensional vector field as a classical variable. Here, $(u,v)$ represents $(u,v)\sim\{2z^{2}-x^{2}-y^{2},\sqrt{3}(x^{2}-y^{2})\}$, where $\bm{r}=(x,y,z)$ is the real space coordinate. The Fourier transform in the wavenumber ${\bm{k}}$ space is defined as
${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}=\begin{bmatrix}u^{\alpha}_{{\bm{k}}}\\ v^{\alpha}_{{\bm{k}}}\end{bmatrix}=\frac{1}{\sqrt{N}}\sum_{\bm{r}}e^{-i{\bm{k}}\cdot\bm{r}}\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}.$ (1)
Here, the summation for ${\bm{r}}$ runs over the A-sublattice sites in the diamond structure and $N$ is the number of the unit cell.
The exchange Hamiltonian in the diamond structure is in general written as
$H=\sum_{{\bm{k}}}[u^{A}_{-{\bm{k}}},v^{A}_{-{\bm{k}}},u^{B}_{-{\bm{k}}},v^{B}_{-{\bm{k}}}]\begin{bmatrix}H^{AA}_{{\bm{k}}}&H^{AB}_{{\bm{k}}}\\ {H^{AB}_{{\bm{k}}}}^{\dagger}&H^{AA}_{{\bm{k}}}\end{bmatrix}\begin{bmatrix}u^{A}_{{\bm{k}}}\\ v^{A}_{{\bm{k}}}\\ u^{B}_{{\bm{k}}}\\ v^{B}_{{\bm{k}}}\end{bmatrix},$ (2)
where $H^{\alpha\beta}_{{\bm{k}}}$ is a $2\times 2$ matrix. We focus on the orders with the ordering wavenumber at the L points: ${\bm{k}}_{1}\equiv(\pi,\pi,\pi)$, ${\bm{k}}_{2}\equiv(-\pi,\pi,\pi)$, ${\bm{k}}_{3}\equiv(-\pi,-\pi,\pi)$, and ${\bm{k}}_{4}\equiv(\pi,-\pi,\pi)$. Since the exchange interactions at the L points are isotropic due to the symmetry, we retain such exchange interactions while neglect anisotropic ones. Although the latter exchange interactions are generally present and can influence the ordering vectors and the excitation energy of quadrupole wave, it is no effect on the transition temperature for the order at the L points. Thus, a model without anisotropic interactions can be a good approximation as a minimal model. In this section, we consider exchange interactions up to the fourth neighbor interactions, while we will simplify the model in the later sections. Hereafter, the lattice constant $a$ is set to $a=1$.
The next nearest-neighbor and the fourth-neighbor bonds are those between the A-A (B-B) sublattices. They are
$H^{AA}_{{\bm{k}}}$ $=[J_{2}\gamma_{2}({\bm{k}})+J_{4}\gamma_{4}({\bm{k}})]\hat{\sigma}_{0}$ (3)
Here, $\hat{\sigma}_{0}={\rm diag}(1,1)$ and $\hat{\sigma}^{x,z}$ is the Pauli matrix and we have followed the phase convention in Ref. [Hattori2014]. $J_{i}$’s are isotropic coupling constants. The expressions for $\gamma$’s are listed in Table 1.
The nearest-neighbor and the third-neighbor bonds are those between the A-B sublattices. For the nearest-neighbor interaction, there is no anisotropic term owing to the symmetry. The form is thus given as
$H^{AB}_{{\bm{k}}}$ $=\frac{1}{2}[J_{1}\gamma_{1}({\bm{k}})+J_{3}\gamma_{3}({\bm{k}})]\hat{\sigma}_{0}.$ (4)
We note that $\gamma_{1}$ and $\gamma_{3}$ are complex. From the neutron experimental data, the ordering patterns are consistent with those expected for $J_{1}>0$.
For describing the symmetry breaking at the L point modulation within our model, it is important to check which interactions are important. As will be explained later, four special wavenumbers at the $\Gamma$ point and the X points $\{(2\pi,0,0),(0,2\pi,0),(0,0,2\pi)\}$ are important other than the L point. First, among the same sub-lattice interactions, we note that the next-nearest interaction $\gamma_{2}$ vanishes at the four L points, while the fourth neighbor one $\gamma_{4}$ is finite as shown in Table 1. This means that positive $J_{4}$ favors the order at the L points while does not those at the $\Gamma$ and X points. For the inter sublattice ones, the nearest-neighbor interaction $\gamma_{1}$ and the third-neighbor one $\gamma_{3}$ are finite, while both vanish at the X points. In this sense, the effects of $J_{1}$ and $J_{3}$ are similar. Diagonalizing the matrix in Eq. (2), one obtains the eigenvalues $J({\bm{k}})$. Figure 1 shows $J({\bm{k}})$ for typical parameter sets with $J_{1}=1$ as a unit of energy. For the nearest-neighbor model with $J_{4}=0$ in Fig. 1(a), $J({\bm{k}})$ is smallest at the $\Gamma$ point, leading to a Néel type antifero quadrupole order. When $J_{4}$ is sufficiently large or $J_{2}$ is positive and not too large, the $J({\bm{k}})$ has minima at the L points as shown in Figs. 1(b) and 1(c), and these cases are situations we analyze in this paper.
Table 1: exchange interaction coefficients. The values of $\gamma({\bm{k}})$’s at ${\bm{k}}=(0,0,0),(0,0,2\pi)$, and $(\pi,\pi,\pi)$ are also shown. $\eta_{{\bm{k}}}\equiv e^{i(k_{x}+k_{y}+k_{z})/4}$, $\zeta^{\pm}_{{\bm{k}}}\equiv\pm 2(\cos\frac{k_{x}}{2}+\cos\frac{k_{y}}{2}+\cos\frac{k_{z}}{2})-3$. The mean-field second-order transition temperatures $T_{c}$’s are $T_{c}=2|J_{1}-3J_{3}|+6J_{4}$ for the L points, $T_{c}=4|J_{1}+3J_{3}|-12J_{2}-6J_{4}$ for the $\Gamma$ points, and $T_{c}=4J_{2}-6J_{4}$ for the X points.
notation expression $\Gamma(0,0,0)$ X(0,0,$2\pi$) L($\pi,\pi,\pi$)
$\gamma_{2}({\bm{k}})$ $2\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{z}}{2}\cos\frac{k_{x}}{2}\right)$ $6$ $-2$ 00
$\gamma_{4}({\bm{k}})$ $\cos k_{x}+\cos k_{y}+\cos k_{z}$ $3$ $3$ $-3$
$\gamma_{1}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\right)$ $4$ 00 $-2$
$\gamma_{3}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}\zeta^{+}_{{\bm{k}}}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\zeta^{-}_{{\bm{k}}}\right)$ $12$ 00 $6$
II.2 Landau Expansion
We now focus on the order parameter at the four equivalent L points. Note that ${\bm{k}}_{1}+{\bm{k}}_{2}+{\bm{k}}_{3}+{\bm{k}}_{4}\equiv{\bf 0}$ and that the fields at the L points are real since ${\bm{k}}_{n}\equiv-{\bm{k}}_{n}$.
The Landau’s free energy $F$ for the quadrupole moments is written as
$F$ $=\sum_{{\bm{k}}}[{\boldsymbol{\phi}}^{A}_{-{\bm{k}}},{\boldsymbol{\phi}}^{B}_{-{\bm{k}}}]\left[\frac{T}{2}+\hat{J}({\bm{k}})\right]\begin{bmatrix}{\boldsymbol{\phi}}^{A}_{{\bm{k}}}\\ {\boldsymbol{\phi}}^{B}_{{\bm{k}}}\end{bmatrix}$
$+\sum_{\alpha}\Big{[}-b\sum_{{\bm{k}},{\bm{k}}^{\prime},{\bm{k}}^{\prime\prime}}u^{\alpha}_{{\bm{k}}}\Big{(}u^{\alpha}_{{\bm{k}}^{\prime}}u^{\alpha}_{{\bm{k}}^{\prime\prime}}-3v^{\alpha}_{{\bm{k}}^{\prime}}v^{\alpha}_{{\bm{k}}^{\prime\prime}}\Big{)}\delta_{{\bm{k}}+{\bm{k}}^{\prime}+{\bm{k}}^{\prime\prime},{\bm{G}}}$
$+\frac{c}{4}\sum_{{\bm{k}},{\bm{k}}^{\prime},{\bm{k}}^{\prime\prime},{\bm{k}}^{\prime\prime\prime}}({\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}}\cdot{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime}})({\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime\prime}})\delta_{{\bm{k}}+{\bm{k}}^{\prime}+{\bm{k}}^{\prime\prime}+{\bm{k}}^{\prime\prime\prime},{\bm{G}}}\Big{]}.$ (5)
Here, $\hat{J}({\bm{k}})$ is the $4\times 4$ matrix in Eq. (2) with its eigenvalue $J({\bm{k}})$ and $T$ is the temperature and $\bm{G}$ is a reciprocal lattice vector. We have taken into account local cubic and fourth-order terms only, which arise from local anisotropy such as CEF potential.
II.3 L point free energy
When squeeze out the L point components ${\boldsymbol{\phi}}^{\alpha}_{n}=(u_{n}^{\alpha},v_{n}^{\alpha})\equiv{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}(n=1,2,3,4)$ with the short handed notation, one finds
$F_{\rm L}$ $=\frac{a_{\rm L}}{2}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}+J_{\rm L}\sum_{n}{\boldsymbol{\phi}}^{A}_{n}\cdot{\boldsymbol{\phi}}^{B}_{n}+\frac{c}{4}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{4}$
$+\frac{c^{\prime}}{4}\sum_{\alpha}\sum_{m<n}|{\boldsymbol{\phi}}^{\alpha}_{m}|^{2}|{\boldsymbol{\phi}}^{\alpha}_{n}|^{2}+\frac{c^{\prime\prime}}{4}\sum_{\alpha}\sum_{m<n}({\boldsymbol{\phi}}_{m}^{\alpha}\cdot{\boldsymbol{\phi}}_{n}^{\alpha})^{2}$
$+\frac{c^{\prime\prime\prime}}{4}\sum_{\alpha}\sum_{m<n}\sum_{m^{\prime}<n^{\prime}}({\boldsymbol{\phi}}^{\alpha}_{m}\cdot{\boldsymbol{\phi}}^{\alpha}_{n})({\boldsymbol{\phi}}^{\alpha}_{m^{\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{n^{\prime}}).$ (6)
Note that $b$ terms are absent owing to the wavnumber conservation and $(m,n,m^{\prime},n^{\prime})=(1,2,3,4)$, $(1,3,2,4)$, and $(1,4,2,3)$ in the last term in the right-hand side. The coupling constants are denoted as
$a_{\rm L}$ $\equiv T-6J_{4},\quad J_{\rm L}\equiv-2J_{1}+6J_{3},$ (7)
$c^{\prime}=2c,\quad c^{\prime\prime}=4c,\quad c^{\prime\prime\prime}=8c.$ (8)
Note that ${\bm{k}}_{n}\equiv-{\bm{k}}_{n}$ and thus the ${\boldsymbol{\phi}}_{n}\cdot{\boldsymbol{\phi}}_{n}$ pair in the summation in $\sum_{{\bm{k}}}$ appears once. Since each of the fourth order terms are allowed by the cubic symmetry, we denote them as independent parameters.
For discussing a single-$\bm{q}$ order at the L points, $c^{\prime}$, $c^{\prime\prime}$, and $c^{\prime\prime\prime}$ terms are irrelevant and the fourth-order free energy is simply given by
$F_{L}^{1(4)}=\frac{c}{4}\sum_{\alpha}|{\boldsymbol{\phi}}^{\alpha}|^{4}.$ (9)
When multiple-$\bm{q}$ states are considered, the coupling constants $c^{\prime}$, $c^{\prime\prime}$, and $c^{\prime\prime\prime}$ matter as will be discussed in Sec. V.
II.4 $\Gamma$ point free energy
Under finite magnetic fields, the uniform quadrupole moment ${\boldsymbol{\phi}}_{0}^{\alpha}\equiv{\boldsymbol{\phi}}_{{\bm{k}}=\bf 0}^{\alpha}$ is induced via the effective field $\tilde{\bm{h}}$ in the second order perturbation of the magnetic field $\bm{H}$. The free energy up to the fourth-order terms for the $\Gamma$ point modes are
$F_{\Gamma}$ $=\sum_{\alpha}\Big{(}-\tilde{\bm{h}}\cdot{\boldsymbol{\phi}}_{0}^{\alpha}+\frac{a_{\Gamma 0}}{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}\Big{)}+J_{\Gamma 0}{\boldsymbol{\phi}}_{0}^{\rm A}\cdot{\boldsymbol{\phi}}_{0}^{\rm B},$
$+\sum_{\alpha}\Big{\{}-bu_{0}^{\alpha}\Big{[}\big{(}u_{0}^{\alpha}\big{)}^{2}-3\big{(}v_{0}^{\alpha}\big{)}^{2}\Big{]}+\frac{c}{4}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{4}\Big{\}},$ (10)
where
$a_{\Gamma 0}$ $=T+12J_{2}+6J_{4},\quad J_{\Gamma 0}=4J_{1}+12J_{3},$ (11)
and the effective field is given by
$\tilde{\bm{h}}=\begin{bmatrix}\tilde{h}_{u}\\ \tilde{h}_{v}\end{bmatrix}\equiv g\begin{bmatrix}2H_{z}^{2}-H_{x}^{2}-H_{y}^{2}\\ \sqrt{3}(H_{x}^{2}-H_{y}^{2})\end{bmatrix},$ (12)
and $g=(g_{J}\mu_{\rm B})^{2}[7/(3E_{4})-1/E_{5}]$ given by the excited $\Gamma_{4,5}$ CEF energy $E_{4,5}$.Hattori and Tsunetsugu (2014) $\mu_{\rm B}$ is the Bohr magneton and $g_{J}$ is the Lande’s $g$ factor for the multiplet with the angular momentum $J$; $g_{J}=4/5$ for the PrIr2Zn20 with $J=4$.
II.5 Mode coupling terms
We now discuss mode coupling terms consisting of the fields at the L points and those at $\Gamma$ points. The lowest order contributions come from the cubic terms as
$\delta F_{3}=$ $-3\lambda_{b}\sum_{\alpha,n}\Big{\{}u_{0}^{\alpha}\Big{[}\big{(}u_{n}^{\alpha}\big{)}^{2}-\big{(}v_{n}^{\alpha}\big{)}^{2}\Big{]}-v_{0}^{\alpha}(2u_{n}^{\alpha}v_{n}^{\alpha})\Big{\}}$
$=-3\lambda_{b}\sum_{\alpha,n}\phi_{0}^{\alpha}\big{(}\phi_{n}^{\alpha}\big{)}^{2}\cos(\theta^{\alpha}_{0}+2\theta^{\alpha}_{n}),$ (13)
Here we have parameterized
${\boldsymbol{\phi}}^{\alpha}_{\bf 0}$ $=\phi^{\alpha}_{0}\begin{bmatrix}\cos\theta^{\alpha}_{0}\\ \sin\theta^{\alpha}_{0}\end{bmatrix},\quad{\boldsymbol{\phi}}^{\alpha}_{n}=\phi^{\alpha}_{n}\begin{bmatrix}\cos\theta^{\alpha}_{n}\\ \sin\theta^{\alpha}_{n}\end{bmatrix},$ (14)
and the coupling constant $\lambda_{b}$ as an independent parameter. If $\lambda_{b}$ is estimated by Eq. (5), $\lambda_{b}=b$. The next leading order terms are the fourth order and given by
$\delta F_{4}=\sum_{\alpha,n}\left[\frac{\lambda^{(1)}_{c}}{2}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}+\lambda_{c}^{(2)}({\boldsymbol{\phi}}_{n}^{\alpha}\cdot{\boldsymbol{\phi}}_{0}^{\alpha})^{2}\right].$ (15)
Again, the coupling constant $\lambda_{c}^{(1,2)}$ has been introduced as an independent parameter. If $\lambda_{c}^{(1,2)}$ is estimated by Eq. (5), $\lambda^{(1)}_{c}=\lambda_{c}^{(2)}=c$.
In principle, ${\boldsymbol{\phi}}_{0}^{\alpha}$ can couple with ${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}$ in the form of $\sim{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$, ${\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$. In this paper, we retain the order parameter ${\boldsymbol{\phi}}_{n}^{\alpha}$ and the induced moment ${\boldsymbol{\phi}}_{0}^{\alpha}$ under magnetic fields since the other fields are not important for the phase transition. Some renormalization effects arising from the field at the X points will be discussed in Sec. IV and Appendix B.
III Results: single-q orders
In this section, we carry out minimization of the free energy $F_{L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}$ [Eqs. (6), (10), (13), and (13)] with the single-$\bm{q}$ assumptions. The purpose here is to understand how mode coupling effects affect the order parameter configuration. The results in this section will become bases for understanding the full phase diagram shown in Sec. IV where we sill take into account multiple-$\bm{q}$ configurations. Since we are interested in the phase transition owing to order parameter at the L point, we assume the second-order coefficients of the free energy at the other wavenumber are always positive. We first discuss the candidates of the stable configurations by analyzing the third and fourth order terms in $F_{\rm L}$. Then, we will show the numerical results of full temperature-field phase diagram.
III.1 Qualitative analysis
First, we notice that the inter-sublattice correlations are governed by $J_{\rm L}=-2J_{1}+6J_{3}$ for ${\boldsymbol{\phi}}_{n}^{\alpha}$ in Eq. (6). To reproduce the quadrupole configuration propposed by the neutron scattering experimentsIwasa et al. (2017), one needs $J_{\rm L}<0$. Thus, it is legitimate to set
${\boldsymbol{\phi}}_{\rm L}\equiv{\boldsymbol{\phi}}_{n}^{A}={\boldsymbol{\phi}}_{n}^{B}=\begin{bmatrix}u_{\rm L}\\ v_{\rm L}\end{bmatrix}=\phi_{\rm L}\begin{bmatrix}\cos\theta_{\rm L}\\ \sin\theta_{\rm L}\end{bmatrix}.$ (16)
Here, we have chosen one of the L points at ${\bm{k}}_{n}$ as a representative one since the four L points are equivalent even when $\bm{h}$ is finite. For the modes at the $\Gamma$ point, which are gapped and just induced by magnetic fields or by the square of the order parameter at the L point, we set
${\boldsymbol{\phi}}_{0}\equiv{\boldsymbol{\phi}}_{0}^{A}={\boldsymbol{\phi}}_{0}^{B}=\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}=\phi_{0}\begin{bmatrix}\cos\theta_{0}\\ \sin\theta_{0}\end{bmatrix}.$ (17)
Second, we note that the single-$\bm{q}$ free energy $F_{L}$ for the L point order parameter is isotropic. This means there is no anisotropy at the transition temperature $T_{c}=6J_{4}+2|J_{1}-3J_{3}|$ at ${\bm{H}}={\bf 0}$. This isotropic nature, however, is destroyed by the mode coupling to the fields at the $\Gamma$ points through $\delta F_{3}$ and $\delta F_{4}$. When $T\lesssim T_{c}$, the induced ${\boldsymbol{\phi}}_{0}$ is sufficiently small, and thus the third and fourth order terms in $F_{\Gamma}$ can be irrelevant. In this situation the relative directions between ${\boldsymbol{\phi}}_{n}$ and ${\boldsymbol{\phi}}_{0}$ is determined by $\delta F_{3}$ and $\delta F_{4}$ as
$\cos(\theta_{0}+2\theta_{\rm L})=1\ {\rm and}\ \cos(\theta_{0}-\theta_{\rm L})=0.$ (18)
This leads to
$\theta_{0}=\frac{(2j-1)\pi}{3},\quad\theta_{\rm L}=\frac{(2j-1)\pi}{3}\pm\frac{\pi}{2},$ (19)
with $j=1,2,3$. These six domains are degenerate at ${\bm{H}}={\bf 0}\ (\tilde{\bm{h}}={\bf 0})$. See Fig. 2(a). We should comment that the three directions of ${\boldsymbol{\phi}}_{0}$ do not coincide with those favored by the $b(>0)$ term in $F_{\Gamma}:-b\phi_{0}^{3}\cos(3\theta_{0})$. Since ${\boldsymbol{\phi}}_{0}$ is not the primary order parameter, the third-order term is expected to play a minor role and the configurations (19) are robust at least $\tilde{\bm{h}}={\bf 0}$.
Figure 2: Schematic single-$\bm{q}$ order parameter configurations for (a) ${\bm{H}=0}$, (b) small ${\bm{H}}$ and low $T$, (c) small ${\bm{H}}$ and high $T$, and (d) large ${\bm{H}}$. (e) large ${\bm{H}}$. Blue (red) arrows indicate ${\boldsymbol{\phi}}_{0}$(${\boldsymbol{\phi}}_{\rm L}$). Filled red circles represent the directions $\theta_{L}$ given by Eq. (19) and open blue circles do $\theta_{\rm 0}$. A configuration for large ${\bm{H}}$ in the phase II in (c) is shown in (e) and that for ${\bm{H}}\parallel[110]$ in (f).
Now let us consider the cases for finite magnetic fields. For high symmetric directions $\bm{H}=\tfrac{H}{\sqrt{2}}(\sin\theta,\sin\theta,\sqrt{2}\cos\theta)$ with $20\leq\theta\leq\pi/$ 2, the effective field reads as $\tilde{h}_{u}=gH^{2}(3\cos^{2}\theta-1)$ and $\tilde{h}_{v}=0$. Note that $\tilde{\bm{h}}={\bf 0}$ at $\bm{H}\parallel[111]$ and $\tilde{h}_{u}>0$ for $\theta<\tan^{-1}\sqrt{2}\equiv\theta_{111}$, while $\tilde{h}_{u}<0$ for $\theta>\theta_{111}$. The effective field induces $u_{0}$, which leads to two “transition temperatures” $\tilde{T}_{c}^{u}$ for the $u_{\rm L}$ order and $\tilde{T}_{c}^{v}$ for $v_{\rm L}$. They are given by
$\tilde{T}_{c}^{u}$ $=T_{c}+6\lambda_{b}u_{0}-3\lambda_{c}u_{0}^{2},\ \tilde{T}_{c}^{v}=T_{c}-6\lambda_{b}u_{0}-\lambda_{c}u_{0}^{2},$ (20)
where we have set $\lambda_{c}^{(1)}=\lambda_{c}^{(2)}=\lambda_{c}$ for simplicity. Note that $\tilde{T}_{c}^{u,v}$ is a hypothetical transition temperature from the disordered state. The difference between the two is $\Delta\tilde{T}_{c}=\tilde{T}_{c}^{u}-\tilde{T}_{c}^{v}=2\lambda_{c}(\phi_{0}^{*}-u_{0})u_{0}$ with $\phi_{0}^{*}\equiv 6\lambda_{b}/\lambda_{c}$. For $u_{0}>0$, this means $\tilde{T}_{c}^{u}>\tilde{T}_{c}^{v}$ for $u_{0}<\phi_{0}^{*}$, i.e., for low fields, while $\tilde{T}_{c}^{v}>\tilde{T}_{c}^{u}$ for the higher field with $u_{0}>\phi_{0}^{*}$. In contrast, $\Delta\tilde{T}_{c}<0$ for $u_{0}<0$.
As is clear from the expression of $\Delta\tilde{T}_{c}$, it is expected that there is no additional phase transition within the single-$\bm{q}$ ${\boldsymbol{\phi}}_{n}$ sector for $u_{0}<0$, i.e., $\theta_{0}>\theta_{111}$. In the following, we will concentrate on the cases for $u_{0}>0$ ($\theta_{0}<\theta_{111}$). For $u_{0}>0$, the linear coupling between ${\boldsymbol{\phi}}_{0}$ and $\tilde{h}_{u}$ can alter the stable configuration from the above six domains (19) since their energy gain is given by the second order in ${\boldsymbol{\phi}}_{0}$. Since $\Delta\tilde{T}_{c}>0$, the phase transition from the disordered phase occurs in the $u_{\rm L}$ component: $\theta_{n}=0$ or $\pi$. Here, the two choices 00 and $\pi$ represent a simple translation and we will consider $\theta_{\rm L}=0$ below. For small $\tilde{h}_{u}$, $\delta F_{3}$ is more important than $\delta F_{4}$. This means $\theta_{0}=0$ satisfying $02\theta_{\rm L}+\theta_{0}=$ 0 [Fig. 2(b)]. Such a collinear state possesses the free energy $F_{\rm col}$ and the fluctuations around this is given as
$\delta F_{\rm col}$ $\simeq\lambda_{c}\phi_{0}\phi_{\rm L}^{2}\begin{bmatrix}\theta_{0}&\theta_{\rm L}\end{bmatrix}\hat{A}\begin{bmatrix}\theta_{0}\\ \theta_{\rm L}\end{bmatrix},$ (21)
$\hat{A}$ $=\begin{bmatrix}(\lambda_{c}\phi_{\rm L}^{2})^{-1}(\tilde{h}_{u}+9b\phi_{0}^{2})+\frac{1}{2}\phi_{0}^{*}-2\phi_{0}&\phi_{0}^{*}+2\phi_{0}\\ \phi_{0}^{*}+2\phi_{0}&2\phi_{0}^{*}-2\phi_{0}\end{bmatrix}.$ (22)
Indeed, the instability to finite $\theta_{{\rm L},0}$ occurs when one of the eigenvalue of the matrix $\hat{A}$ in Eq. (22) becomes zero. This takes place when
$\tilde{h}_{u}^{c1}=27\lambda_{b}\frac{\phi_{0}}{\phi_{0}^{*}-\phi_{0}}\phi_{\rm L}^{2}-9b\phi_{0}^{2}.$ (23)
Note that the factor $\phi_{0}^{*}-\phi_{0}$ of the denominator coincides with the factor in $\Delta T_{c}$. The configuration with $\theta_{0}\neq 0$ and $\theta_{\rm L}\neq 0$ gradually crossovers toward one for $\bm{H}=0$. See Fig. 2(c). When $\tilde{\bm{h}}$ is sufficiently large, $u_{0}$ grows and $\delta F_{4}$ plays an important role. Thus, ${\boldsymbol{\phi}}_{\rm L}$ is locked at $\theta_{\rm L}=\pm\pi/2$ as shown in Fig. 2(d). This is because $\delta F_{4}>\delta F_{3}$ for large $\tilde{h}_{u}$. By the similar analysis to Eqs. (21)-(23), the boundary determined by the field below which $|\theta_{\rm L}|<\pi/2$ [Fig. 2(e)] is given as
$\tilde{h}_{u}^{c2}=27\lambda_{b}\frac{\phi_{0}}{\phi_{0}-\phi_{0}^{*}}\phi_{\rm L}^{2}-9b\phi_{0}^{2}.$ (24)
For sufficiently large $\lambda_{b}$, the $\theta_{\rm L}=\pm\pi/2$ phase never appears because of $\tilde{T}_{c}^{v}<0$. Of course, for extremely large $\tilde{\bm{h}}$, the L point order vanishes trivially.
Finally, we briefly comment about the cases for $u_{0}<0$ ($\theta>\theta_{111}$) under finite magnetic fields. As we have noted, there is no phase transition once a quadrupole moment with $\theta_{\rm L}=\pm\pi/2$ is ordered as shown in Fig. 2(f). This is because $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$ can minimize $\delta F_{3}$, $\delta F_{4}$, and $F_{\Gamma}$, simultaneously.
III.2 Phase diagram
Now we show the results of numerical minimization of the single-$\bm{q}$ free energy (6), (10), (13), and (15). We will show the result for ${\bm{H}}\parallel[001]$, i.e., $\tilde{h}_{u}>0$ and then for ${\bm{H}}\parallel[110]$, $\tilde{h}_{u}<0$. To understand the qualitative nature of the phase diagram, it is sufficient to discuss the cases for $J_{2}=J_{3}=0$.
Figures 3(a)–3(c) show the $T$-$H$ phase diagram for $\bm{H}\parallel[001]$, where the color represents the angle for the L point moment $\theta_{\rm L}$: (a) $b=0.1$, (b) $b=0.05$ , and (c) $b=0.01$ with $c=0.5$ fixed. As discussed in the previous section, the phase III is stabilized under high magnetic fields. This phase, however, is suppressed when $b$ becomes larger. We note that the collinear phase (phase I) appears for high $T$ and intermediate $H$. The reason why the phase I cannot appear at the lower $T$ is understood by seeing Eq. (23); there is no solution of the critical field $\tilde{h}_{u}^{c1}$ when $\phi_{\rm L}^{2}$, which is approximately $H$ independent, is large.
Figure 3: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, $J_{2}=J_{3}=0.0$, $J_{4}=1.0$, and $c=0.5$. (a) $b=0.1$, (b) $b=0.05$, and (c) $b=0.01$. Color map represents the L point order parameter angle $\theta_{\rm L}$.
For $\bm{H}\parallel[110]$, the situation becomes quite simpler. Figure 4(a) shows the $T$-$H$ phase diagram for $\bm{H}\parallel[110]$. As discussed in the last part in the previous section, there is only one stable phase (phase IV), where $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$.
Figures 4(b) and 4(c) show the real space quadrupole configurations with the ordering vector $\bm{k}_{1}$ for $\bm{H}\parallel[001]$ and $\bm{H}\parallel[110]$, respectively. For both cases, the quadrupole orders with $\bm{\phi}_{A}=\bm{\phi}_{B}$ are ferroic on a (111) plane, while antiferroic between the neighboring (111) plane as shown in different colors in Fig. 4. When the ordering vector is $\bm{k}_{n}(n=2,3,4)$, the quadrupole configurations are ferroic on the plane perpendicular to ${\bm{k}}_{n}$ with $\bm{\phi}_{A}=-\bm{\phi}_{B}$.
Figure 4: (a) $T$$H$ phase diagram for $\bm{H}\parallel[110]$, $J_{1}=-1$, $J_{2}=J_{3}=0$, $J_{4}=1$, $c=0.5$, and $b=0.05$. Color map represents the L point order parameter angle $\theta_{\rm L}$. (b) $\theta$$H$ phase diagram from [001]($\theta=0$) to [110] ($\theta=\pi/2$).
IV Results: multiple-q orders
In this section, we consider multiple-$\bm{q}$ orders for the quadrupole moments at the L points. We will first introduce a simple symmetric multiple-$\bm{q}$ states with the same amplitudes for all the $\bm{q}_{n}$ in Sec. IV.1. Then, we will show numerical results, where we relax the constraint about the amplitudes of the order parameters.
IV.1 Free energy for symmetric states
In this section, we consider double-$\bm{q}\ (\mathcal{N}=2)$ and quadruple-$\bm{q}\ (\mathcal{N}=4)$ configurations, whose real space moments are given by
${\boldsymbol{\phi}}^{\alpha}(\bm{r})=\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}=\frac{\Phi_{\alpha}}{\sqrt{\mathcal{N}}}\sum_{n=1}^{\mathcal{N}}e^{i{\bm{k}}_{n}\cdot\bm{r}}\begin{bmatrix}\cos\theta_{n}^{\alpha}\\ \sin\theta_{n}^{\alpha}\\ \end{bmatrix},$ (25)
where $\bm{r}$ is labeled by the A sublattice position as before and we have assumed that the magnitude for each ${\bm{k}}_{n}$ is the same for simplicity (symmetric states). In this normalization $(\mathcal{N})$, the quadratic part is unchanged from the single-$\bm{q}$ expression. Thus, types of order realized below the second-order transition at $T=T_{c}$ are determined by the fourth-order terms.
For double-$\bm{q}$ states, the fourth-order terms are
$F_{\rm L}^{2(4)}=\sum_{\alpha}\Phi_{\alpha}^{4}\left[\frac{c}{8}+\frac{c^{\prime}}{16}+\frac{c^{\prime\prime}}{16}(c^{\alpha}_{12})^{2}\right].$ (26)
Here, $c^{\alpha}_{mn}\equiv\cos(\theta^{\alpha}_{m}-\theta^{\alpha}_{n})$. This clearly indicates that $\theta_{1}^{\alpha}-\theta_{2}^{\alpha}=\pm\pi/2$ for $c^{\prime\prime}>0$ (orthogonal double-$\bm{q}$), while $\theta_{1}^{\alpha}-\theta_{2}^{\alpha}=0$ or $\pi$ for $c^{\prime\prime}<0$ (parallel double-$\bm{q}$). As explained in Appendix A, the double-$\bm{q}$ configurations consist of four ferroic chains running along the direction of ${\bm{k}}_{3}+{\bm{k}}_{4}$ when only the A-sublattice sites are concerned.
For quadruple-$\bm{q}$ states,
$F_{\rm L}^{4(4)}=$ $\sum_{\alpha}\Phi_{\alpha}^{4}\Bigg{[}\frac{c}{16}+\frac{3c^{\prime}}{32}+\frac{c^{\prime\prime}}{64}\sum_{m<n}(c^{\alpha}_{mn})^{2}$
$+$ $\frac{c^{\prime\prime\prime}}{64}(c_{12}^{\alpha}c_{34}^{\alpha}+c_{13}^{\alpha}c_{24}^{\alpha}+c_{14}^{\alpha}c_{23}^{\alpha})\Bigg{]}.$ (27)
Note that the $c^{\prime\prime\prime}$ term can be negative, e.g., when $\theta_{1}=\theta_{2}=\theta_{3}=\theta_{4}+\pi$. This means that when $c^{\prime\prime\prime}$ is large, such quadruple-$\bm{q}$ orders are favored. The lowest free energy under the constraint (8) is $F_{\rm L}^{4(4)}=\tfrac{c}{4}\sum_{\alpha}\Phi_{\alpha}^{4}$ with $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(\theta,\theta,\theta^{\prime},\theta^{\prime}+\pi)$, where $\theta$ and $\theta^{\prime}$ are arbitrary and those with any two $\theta_{n}$’s permutations are equivalent ones. The unit cell for the quadrupole-$\bm{q}$ configurations contains eight A-sublattice sites and thus sixteen sites in total. See also Fig. 9 in Appendix A.
In Eqs. (26) and (27), $\theta_{n}$ dependence appears only in $c^{\prime\prime}$ and $c^{\prime\prime\prime}$ terms. Figure 6 shows the phase diagram at $T=T_{c}$ on the $c^{\prime\prime\prime}$$c^{\prime\prime}$ plane for (a) $c^{\prime}>2c$ and (b) $c^{\prime}<2c$ with $c>0$. For large $c^{\prime}$, multiple-$\bm{q}$ states are not realized as is evident from the expressions (26) and (27). See Fig. 6(a). The conditions given by Eq. (8) just correspond to the phase boundary for $c^{\prime\prime},c^{\prime\prime\prime}>0$ and $c^{\prime}=2c$, where all the single-$\bm{q}$, double-$\bm{q}$, and quadruple-$\bm{q}$ are degenerate. Note that away from the phase boundaries between quadruple-$\bm{q}$ and double- or single-$\bm{q}$ phases, the direction of the $\bm{\phi}_{n}$ in the quadruple-$\bm{q}$ states are pinned to $(\theta_{1}^{\alpha},\theta_{2}^{\alpha},\theta_{3}^{\alpha},\theta_{4}^{\alpha})$$=$(0,0,0,0)$0$(0,0,0,0)( 0 , 0 , 0 , 0 ), $(0,0,\pi,\pi)$, $(0,\pi,\pi,0)$, $(0,\pi,0,\pi)$ for $c^{\prime\prime\prime}<0$, while $(\theta_{1}^{\alpha},\theta_{2}^{\alpha},\theta_{3}^{\alpha},\theta_{4}^{\alpha})$$=$$(0,0,0,\pi)$, $(0,0,\pi,0)$, $(0,\pi,0,0)$, $(\pi,0,0,0)$ for $c^{\prime\prime\prime}>0$. Here, global sign changes $\bm{\phi}^{\alpha}\to-\bm{\phi}^{\alpha}$: $\theta_{n}^{\alpha}\to\theta_{n}^{\alpha}+\pi$, have not been included since they are also realized by trivial translations.
Figure 5: Phase diagram at $T=T_{c}$ in the $c^{\prime\prime\prime}$$c^{\prime\prime}$ plane. (a) $c^{\prime}>2c$, where the phase boundary between the single-$\bm{q}$(1q) and quadruple-$\bm{q}$(4q) is given by $c^{\prime\prime}=|c^{\prime\prime\prime}|/2-(c^{\prime}-2c)$. (b) $c^{\prime}<2c$, where the phase boundary between the double-$\bm{q}$(2q) and quadruple-$\bm{q}$(4q) is given by $c^{\prime\prime}=|c^{\prime\prime\prime}|/2+(2c-c^{\prime})/3$.
So far, we have treated the fourth-order coefficients $c$’s as a phenomenological parameters. Here, we discuss the renormalization owing to modes other than $\bm{\phi}_{0}$ and $\bm{\phi}_{n}$. In the lowest orders, the cubic terms in Eq. (5) are the most important. Among them, there are special terms which connect two L-point modes at $\bm{q}_{m}$ and $\bm{q}_{n}$ ($m\neq n$) via $\sim\phi_{\bm{k}}\phi_{m}\phi_{n}$, where ${\bm{k}}$ is one of the wavenumbers at the X points $\bm{q}_{\rm X}=(2\pi,0,0),\bm{q}_{\rm Y}=(0,2\pi,0)$, or $\bm{q}_{\rm Z}=(0,0,2\pi)$. The combination of $({\bm{k}},\bm{q}_{m},\bm{q}_{n})$ is restricted by the Umklapp condition ${\bm{k}}+\bm{q}_{m}+\bm{q}_{n}$ and there are two sets of $(m,n)$ for a given ${\bm{k}}$. As shown in Appendix B, these couplings lead to suppression of the coefficients $c^{\prime}$ and $c^{\prime\prime\prime}$. See Eq. (38). Furthermore, nontrivial couplings between the $\Gamma$ and the L points fields appear after integrating out the X point fields as $\sim\phi_{0}\phi_{m}^{2}\phi_{n}^{2}$ and $\sim\phi_{0}\phi_{1}\phi_{2}\phi_{3}\phi_{4}$. They play an important role under magnetic fields $\bm{H}$ since $\phi_{0}$ is induced by $\bm{H}$, and thus, these terms renormalize $c$’s with varying their magnitude depending on $\bm{H}$.
The suppression of $c^{\prime}$ and $c^{\prime\prime\prime}$ due to the renormalization with retaining $c^{\prime\prime}>0$ suggests that the orthogonal double-$\bm{q}$ states are more stable. However, the coupling between the fields at the $\Gamma$ and the L points possesses an energy gain for the single-$\bm{q}$ orders and double-$\bm{q}$ ones with $\phi_{m}\neq\phi_{n}$. This is because there is a finite $\phi_{0}$ induced in the single-$\bm{q}$ states, while there is no gain for the double-$\bm{q}$ states with $\phi_{n}=\phi_{m}$ with $\theta_{n}-\theta_{m}=\pm\pi/2$. See the coupling in Eq. (13). As we will see in Sec. IV.2, the nontrivial terms generated by the integration of the X points fields choose some particular $\theta$ among degenerate orthogonal double-$\bm{q}$ configurations.
IV.2 Phase diagram
Now, we discuss $T$$H$ phase diagram ($H\equiv|\bm{H}|$) taking into account multiple-$\bm{q}$ orders in addition to the single-$\bm{q}$ ones. We numerically minimize the free energy $F_{\rm L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}+\delta F$, where $\delta F$ is due to the integration of the X point fields as discussed in Appendix B and the others are given in Eqs. (6), (10), (13), and (15). Figures 6 and 8 show phase diagrams for $\bm{H}\parallel[001]$, i.e., $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\geq 0)$, and double-$\bm{q}$ configurations in the quadrupole space and real space. Figure 6(a) shows the phase diagram for $J_{1}=1$, $c=0.5$, and $b=0.1$. The exchange coupling constants are the same as used in Fig. 3: $(J_{2},J_{3},J_{4})=(0.0,0.0,1.0)$. Except for the low-field region, the phase diagram is similar to those shown in Fig. 3, showing the single-$\bm{q}$ orders I, II, and III. An orthogonal double-$\bm{q}$ configuration I2q with $\bm{\phi}_{{\bm{k}}_{1}}=(\phi_{1},0)$ and $\bm{\phi}_{{\bm{k}}_{2}}=(0,\phi_{2})$ is stabilized for finite magnetic fields at high $T$ region below the phase I. Here, we take a domain with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. The configuration of I2q state in the quadrupole space is shown in Fig. 8(a) and its real-space one is illustrated in Fig. 8(b). At ${\bm{H}}=0$, the orthogonal double-$\bm{q}$ states with $\phi_{1}=\phi_{2}$ discussed in Sec. IV.1 are realized and degenerate in the angle $\theta=0$. As ${\bm{H}}$ increases, $\phi_{1}$ becomes larger than $\phi_{2}$. Since the phase I is the single-$\bm{q}$ order with $v_{\rm L}=0$, the symmetry breaking from the phase I to the phase I2q is caused by the emergence of $\bm{\phi}_{2}=(0,\phi_{2})$ through the second-order transition. In this sense, $\phi_{1}>\phi_{2}$ is quite natural. In the free energy, this is due to the $\delta F_{3}$, which is minimized when the L point field is parallel to $\bm{\phi}_{0}=(u_{0},0)$, while maximized when it is perpendicular. As for the phase boundary between the phase I2q and the phase II, it is of first-order, while the other phase boundaries are second-order.
At lower $T$, the induced $\Gamma$ point moment $\bm{\phi}_{0}$ favors the configuration III2q shown in Figs. 8(c) and 8(f), where $\bm{\phi}_{1}=\tfrac{1}{\sqrt{2}}\phi_{1}(1,1)$ and $\bm{\phi}_{2}=\tfrac{1}{\sqrt{2}}\phi_{2}(-1,1)$ with $\phi_{1}=\phi_{2}$. This configuration is stabilized by the nontrivial coupling discussed in the last part of Sec. IV.1. See also the last term in Eq. (39). In Fig. 6(a), an intermediate phase II2q toward the phase III2q appears at low temperatures. The quadrupole configurations in the phase II2q are “rotated” from those in the phase I2q with keeping $\bm{\phi}_{1}\perp\bm{\phi}_{2}$ as shown in Fig. 8(b). For smaller $b=0.05$ [Fig. 6(b)], the phase I, II, and the double-$\bm{q}$ orders shrink, while the phase III expands. The overall phase diagram, however, is unchanged.
As an example for more general cases, we show the phase diagram for $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ in Fig. 6(c). For this parameter set, the second-order coefficient $a_{\rm X}$ of the X point fields almost vanish at $T=0$ as shown in Fig. 1(c), the renormalization effects due to the X point fields grows rapidly at low temperatures. This causes the strong stabilization of the phase III2q in Fig. 6(c), showing reentrant behavior as a function of $H$. Interestingly, there is a thin region of the phase II2q. The phase boundary between the phase II2q and II is of the second order. Similarly, that between the small region of phase I2q and III is also second order. We note that the boundaries indicated by the dashed lines are not symmetry breaking phase boundary since the symmetry of the three double-$\bm{q}$ phases are the same as is evident from the real space configurations in Figs. 8(d)–8(f). See also the discussion in Appendix C.
For $\bm{H}\parallel[110]$, a phase I2q appears at low field region as shown in Fig. 7. In contrast to the phase I2q for $\bm{H}\parallel[001]$, the quadrupole configuration is $\bm{\phi}_{1}=(0,\phi_{1})$ and $\bm{\phi}_{2}=(\phi_{2},0)$ with $\phi_{1}>\phi_{2}$. This is natural since the neighboring single-$\bm{q}$ phase is the phase IV. The sharp increase of the critical field between I2q and IV phases are due to the fact that $a_{\rm X}$ is close to 0 in (c).
Figure 6: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, and $c=0.5$. (a) $(b,J_{2},J_{3},J_{4})=(0.1,0.0,0.0,1.0)$, (b) $(0.05,0.0,0.0,1.0)$, and (c) $(0.03,0.5,0.3,0.4)$. Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, where double-$\bm{q}$ orders with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$ are assumed. For single-$\bm{q}$ orders, we use the convention of $\phi_{2}=0$. Thick lines represent phase boundaries, where “1st” represents a line of first-order phase transition. Dashed lines are boundaries among double-$\bm{q}$ configurations. Low $T$ regime is not shown since the Landau expansion is not valid in general.
Figure 7: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, and $c=0.5$. (a) $(b,J_{2},J_{3},J_{4})=(0.1,0.0,0.0,1.0)$, (b) $(0.05,0.0,0.0,1.0)$, and (c) $(0.03,0.5,0.3,0.4)$, which correspond to those in Fig. 7(a), 7(b), and 7(c), respectively. Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, similarly to Fig. 6
Figure 8: Schematic orthogonal double-$\bm{q}$ order parameter configurations for (a) I2q: $(\theta_{1},\theta_{2})=(0,\pi/2)$ with $\phi_{1}>\phi_{2}$, (b) II2q: $(\theta_{1},\theta_{2})=(\theta,\theta+\pi/2)$ with $\phi_{1}>\phi_{2}$, and (c) III2q: $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$ and $\phi_{1}=\phi_{2}$. Arrows indicate the order parameters $\bm{\phi}_{1}$ (L1), $\bm{\phi}_{2}$ (L2), and $\bm{\phi}_{0}$ $(\Gamma)$. Four small circles represent quadrupole moments at the A sublattice real-space positions $\bm{r}=(0,0,0),(1,0,0),(\tfrac{1}{2},\frac{1}{2},0)$, and $(\tfrac{1}{2},\frac{1}{2},1)$. Real space L point quadrupole configurations for (d) $(\theta_{1},\theta_{2})=(0,\pi/2)$ and $\phi_{1}=\sqrt{3}\phi_{2}$ in the phase I2q, (e) $(\theta_{1},\theta_{2})=(\pi/24,13\pi/24)$ and $\phi_{1}=1.2\phi_{2}$ in the phase II2q, and (f) $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$, and $\phi_{1}=\phi_{2}$ in the phase III2q. Note that only L point quadrupole components without those at the $\Gamma$ point are shown. Skyblue and gray arrows represent the directions of ${\bm{k}}_{1}$ and $-{\bm{k}}_{2}$. For guide to eyes, some of the nearest-neighbor bonds are connected by white bonds. Circles are the A sublattice atoms in the diamond structure and their colors represent the sublattice (red, green, blue, white). For clarity, the B sublattice atoms are not drawn. Colored bonds running [01¯$1]$ 1 ] directions connects the A sublattice sites with identical quadrupole moments.
V Discussions
V.1 Comparison with the experimental data of PrIr2Zn20
PrIr2Zn20 shows several phase transitions under magnetic field at low temperatures as reported in the early study.Onimaru et al. (2011) However, the order parameters have not been identified except for the $x^{2}-y^{2}$ type quadrupole order with the ordering vector at the L point for $\bm{H}\parallel[110]$ by the neutron scattering experimentIwasa et al. (2017). The analysis of this study is fully based on this fact. Namely, we have assumed that the leading order parameter is the quadrupole moment at the L points. Constructing the symmetry-allowed free energy consisting of them and the related degrees of freedom, we have analyzed the phase diagram for $\bm{H}\parallel[001]$ in addition to [110]. Our analysis demonstrates that a $2z^{2}-x^{2}-y^{2}$ configuration with the ordering vector at the L point is stable at high-$T$ and up to intermediate magnetic field strength for $\bm{H}\parallel[001]$. This suggests that the $2z^{2}-x^{2}-y^{2}$ order can be a promising candidate for the high-$T$ phase found for $\bm{H}\parallel[001]$ in the experimental worksOnimaru et al. (2011); Kittaka et al. (2024). As for the order parameter at zero field, the single-$\bm{q}$ configuration is not stable while the double-$\bm{q}$ I2q is realized with $\phi_{1}=\phi_{2}$. For finite ${\bm{H}}\parallel[001]$, this double-$\bm{q}$ configuration is modified and $\bm{\phi}_{1}=(\phi_{1},0)$ and $\bm{\phi}_{2}=(0,\phi_{2})$ with $\phi_{1}>\phi_{2}$ is favored. $\phi_{2}$ increases as lowering $T$ and there appear phase II2q and III2q by simultaneously rotating $\bm{\phi}_{1}$ and $\bm{\phi}_{2}$. As the cubic anisotropy $b$ increases, the influence of the renormalized fourth-order coefficients becomes larger and favors the phase II2q and III2q.
As magnetic field increases, the double-$\bm{q}$ order is destabilized and single-$\bm{q}$ orders take place as discussed in Sec. III. Thus, as increasing $\bm{H}\parallel[001]$, the double-$\bm{q}$, the single-$\bm{q}$ phase II with both $u_{\rm L}$ and $v_{\rm L}$ finite, and then the phase III with only $v_{\rm L}$ are realized. The trace of the phase boundaries between them have not been clearly observed in the experiments. However, there are several anomalies in the ultrasonic experimentsIshii et al. (2011). We hope the relation between our phase diagram and these anomalies can be clarified in future experimental studies. In addition to this, there is ambiguity about the stability of the phase III, whether the phase III extends toward the higher field as shown in Fig. 3(c) or occupies only small area [Fig. 3(a)]. This can be controlled by the cubic coupling $b$. For $\bm{H}\parallel[110]$, there is no high-T phase in contrast to the case for $\bm{H}\parallel[001]$ in the experiments.Onimaru et al. (2011); Kittaka et al. (2024) This point is consistent with our results. Other anomalies observed in the ultrasound experimentsIshii et al. (2011) similar to those for $\bm{H}\parallel[001]$ are also interesting to examine in more detail in the future works. Since the anomalies are tiny and thus the comparison between the experiments and the present theory is not straightforward, while there is a phase boundary between the phase I2q and IV in Fig. 7.
We also note that the phase diagram as a function of magnetic field from $\bm{H}\parallel[001]$ to $\bm{H}\parallel[110]$ is described in a function of angle $\Theta$, where $\bm{H}=\tfrac{1}{2}H(\sin\Theta,\sin\Theta,2\cos\Theta)$. In Eq. (12), the effective field to the quadrupole moments is written as $\tilde{\bm{h}}=gH^{2}(3\cos^{2}\Theta-1,0)$. Thus, the critical field of quadrupole orders for the angle $\Theta$, $H_{c}(\Theta)$ is expected to be
$H_{c}(\Theta)=\frac{\sqrt{2}H_{c}(0)}{\sqrt{3\cos^{2}\Theta-1}}.$ (28)
Such behavior can be one of strong evidences for quadrupole orders. We hope the future experimental studies on the phase diagram at finite fields clarify this point.
V.2 Future perspectives
We have shown that single- and double-$\bm{q}$ quadrupole orders are realized in our study as a phenomenological model for PrIr2Zn20. The orthogonal double-$\bm{q}$ orders are favored by the effective fourth-order terms due to the X point fields, while the quadruple ones are not. Although we are interested in the orbital orders at the L points in this paper, quadruple-$\bm{q}$ orders at incommensurate ordering vectors along [111] directions has been studied in cubic magnets, showing hedgehog type topological magnetic structure.Okumura et al. (2020, 2022) In principle, such exotic phases are also possible for the orbital model discussed in this paper when the exchange coupling favors incommensurate ordering vectors by introducing interaction frustration or anisotropic interaction ignored in this studyHattori et al. (2024).
For quantitative comparison between the theoretical and experimental results, it is also needed to incorporate the actual crystal electric field states in the real materials. They vary under magnetic fields and thus the parameters in the free energy can change. For theoretical or computational aspects, it is also interesting to take into account the fluctuation effects by using e.g., classical Monte Carlo simulation employed in the previous studies on Pr-based materials.Hattori and Tsunetsugu (2016); Freyer et al. (2018, 2020) The thermal fluctuation may alter the mean-field phase diagram and to check their stability is important issue as one of our future problems.
VI Summary
We have analyzed quadrupole orders with its modulation characterized by the wavenumber at the L points. Their Landau free energy is constructed under magnetic fields $\bm{H}$. We find that the phase diagrams show strong anisotropy. We have clarified that $x^{2}-y^{2}$ type quadrupole order is stabilized along $\bm{H}\parallel[110]$, which is consistent with that proposed in the neutron scattering experiment. We also propose that the unidentified high-temperature phase for $\bm{H}\parallel[001]$ in the experiments is a collinear $2z^{2}-x^{2}-y^{2}$ order and the orthogonal double-$\bm{q}$ configurations are stable at low fields. These results offers further experimental examination of the order parameters in PrIr2Zn20 and related compounds.
Acknowledgment
The authors thank T. Onimaru T. Ishitobi, and S. Kittaka for fruitful discussion. This work was supported by JSPS KAKENHI (Grant Nos. JP23K20824 and JP23H04869).
Appendix A Multiple-q configuration
In this Appendix, we show the real-space modulations for the double-$\bm{q}$ and quadruple-$\bm{q}$ orders since they are rather complicated. The real-space configurations for the double- and quadrupole-$\bm{q}$ orders are given by Eq. (25). Here, we concentrate on the sublattice independent orders $\bm{\phi}(\bm{r})$ with $\Phi^{\alpha}=\Phi$ and $\theta^{\alpha}_{n}=\theta_{n}$ since only such orders appear in this paper. Thus, one rewrites Eq. (25) as
$\bm{\phi}(\bm{r})=\mathbb{E}_{\mathcal{N}}(\bm{r})\cdot\begin{pmatrix}\bm{\phi}_{1}\\ \bm{\phi}_{2}\\ \cdots\end{pmatrix},\ \bm{\phi}_{n}=\frac{\Phi_{n}}{\sqrt{\mathcal{N}}}\begin{pmatrix}\cos\theta_{n}\\ \sin\theta_{n}\end{pmatrix},$ (29)
where $n=1,2,\cdots,\mathcal{N}$ and $\mathbb{E}_{\mathcal{N}}$ is an ${\mathcal{N}}$ dimensional vector
$\mathbb{E}_{2}(\bm{r})$ $=(e^{i{\bm{k}}_{1}\cdot\bm{r}},e^{i{\bm{k}}_{2}\cdot\bm{r}}),$ (30)
$\mathbb{E}_{4}(\bm{r})$ $=(e^{i{\bm{k}}_{1}\cdot\bm{r}},e^{i{\bm{k}}_{2}\cdot\bm{r}},e^{i{\bm{k}}_{3}\cdot\bm{r}},e^{i{\bm{k}}_{4}\cdot\bm{r}}).$ (31)
Here, we have assumed, without loss of generality, double-$\bm{q}$ orders with ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. Since $\bm{r}$ is the position in the fcc lattice, there are four choices for $\mathbb{E}_{2}(\bm{r}):(\pm 1,\pm 1)$, and eight choices for $\mathbb{E}_{4}(\bm{r}):\pm(1,1,1,1),\ \pm(1,1,-1,-1),\ \pm(1,-1,1,-1)$, and ±$\pm(-1,1,1,-1)$. In Fig. 9, $\mathbb{E}_{\mathcal{N}}(\bm{r})$’s are illustrated and the colors distinguish the vectors $\mathbb{E}_{2,4}$. Note that only the A-sublattice positions in the diamond structure are shown. The unit cell for the double-$\bm{q}$ orders contains four A sublattice sites, each of which forms ferroic chain along the direction parallel to ${\bm{k}}_{3}+{\bm{k}}_{4}\parallel(0,-1,1)$ as shown in Fig. 9(a). The primitive translation vector along this direction is unchanged, while those perpendicular to it are modified to $\frac{1}{2}(2,1,1)$ and $\frac{1}{2}(2,-1,-1)$. For quadruple-$\bm{q}$ orders, the primitive translational vectors are as twice large as those for the original fcc lattice: (1,1,0), (1,0,1), and (0,1,1). The unit cell contains eight sites for the quadruple-$\bm{q}$ orders and sixteen when the B-sublattice is included.
Figure 9: (a) Double-$\bm{q}$ configuration $\mathbb{E}_{2}$ at the A sublattice. Tubes connect the same $\mathbb{E}_{2}$’s. Thin lines indicates the NN bonds of diamond structure. Arrows indicates the direction of ${\bm{k}}_{1}$ and $-{\bm{k}}_{2}$. (b) Quadruple-$\bm{q}$ configuration $\mathbb{E}_{4}$. (c) Symmetric quadruple-$\bm{q}$ configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and (d) for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$. For (d), there are four kinds of ferroic chains. Two of them are similar to the double-$\bm{q}$ configuration for $x=2n$($n:$ integers) planes with the $2z^{2}-x^{2}-y^{2}$ orbital, while the other two chains for $x=2n+1$ plane with the $x^{2}-y^{2}$ orbital run along ${\bm{k}}_{1}+{\bm{k}}_{2}$ direction.
Since several double-$\bm{q}$ configurations are shown in Fig. 8, we here examine the real space configuration of the symmetric quadruple-$\bm{q}$ orders discussed in Sec. IV.1, although they are not stabilized in the $T$$H$ phase diagram in Fig. 6. It is evident that the configurations with “collinear” types such as $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,\pi)$ are partially ordered states. For example, the quadrupole moments vanish at the sites where $\sum_{n}e^{i{\bm{k}}_{n}\cdot\bm{r}}=0$ for the former as shown in Fig. 9(c). As an example of “non-collinear” case, Fig. 9(d) illustrates the configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$.
Appendix B Renormalization effects due to the fields at the X points
In this Appendix, we consider the influence of the modes $\bm{Q}_{\ell}^{\alpha}=(Q_{\ell,z}^{\alpha},Q_{\ell,x}^{\alpha})$ ($\ell=$ X, Y, Z) at the X points: $\bm{q}_{\rm X}=(2\pi,0,0),\bm{q}_{\rm Y}=(0,2\pi,0)$, and $\bm{q}_{\rm Z}=(0,0,2\pi)$. They are more important than other modes at general wavenumbers since, for example, $\bm{q}_{1}+\bm{q}_{2}+\bm{q}_{\rm X}\equiv\bf{0}$ leads to cubic mode coupling with the two different fields at the L points.
The free energy for the X points fields are sublattice diagonal and $F_{\rm X}=\sum_{\alpha={\rm A,B}}F_{\rm X}^{\alpha}$, where
$F^{\alpha}_{\rm X}=$ $\frac{1}{2}\sum_{\ell={\rm X,Y,Z}}\bm{Q}^{\alpha}_{\ell}\begin{bmatrix}a_{\rm X}-6bu_{0}^{\alpha}&6bv_{0}^{\alpha}\\ 6bv_{0}^{\alpha}&a_{\rm X}+6bu_{0}^{\alpha}\end{bmatrix}\bm{Q}^{\alpha}_{\ell}$
$-6b\sum_{\ell={\rm X,Y,Z}}\sum_{i=1}^{2}\Big{[}Q_{\ell,z}^{\alpha}C_{\ell i}^{\alpha}-Q_{\ell,x}^{\alpha}S_{\ell i}^{\alpha}\Big{]}+\cdots.$ (32)
Here, the fourth-order terms and the cubic term for three X points fields are neglected since the modes at the X points are not critical. $a_{\rm X}$ and $C$’s and $S$’s consisting of two L points variables are given as
$a_{\rm X}\equiv T-4J_{2}+6J_{4},$ (33)
$\begin{split}&C_{{\rm X}1}^{\alpha}=u_{1}^{\alpha}u_{2}^{\alpha}-v_{1}^{\alpha}v_{2}^{\alpha},\quad C_{{\rm X}2}^{\alpha}=u_{3}^{\alpha}u_{4}^{\alpha}-v_{3}^{\alpha}v_{4}^{\alpha},\\ &S_{{\rm X}1}^{\alpha}=u_{1}^{\alpha}v_{2}^{\alpha}+v_{2}^{\alpha}u_{1}^{\alpha},\quad S_{{\rm X}2}^{\alpha}=u_{3}^{\alpha}v_{4}^{\alpha}+v_{3}^{\alpha}u_{4}^{\alpha},\\ &C_{{\rm Y}1}^{\alpha}=u_{1}^{\alpha}u_{4}^{\alpha}-v_{1}^{\alpha}v_{4}^{\alpha},\quad C_{{\rm Y}2}^{\alpha}=u_{2}^{\alpha}u_{3}^{\alpha}-v_{2}^{\alpha}v_{3}^{\alpha},\\ &S_{{\rm Y}1}^{\alpha}=u_{1}^{\alpha}v_{4}^{\alpha}+v_{1}^{\alpha}u_{4}^{\alpha},\quad S_{{\rm Y}2}^{\alpha}=u_{2}^{\alpha}v_{3}^{\alpha}+v_{2}^{\alpha}u_{3}^{\alpha},\\ &C_{{\rm Z}1}^{\alpha}=u_{1}^{\alpha}u_{3}^{\alpha}-v_{1}^{\alpha}v_{3}^{\alpha},\quad C_{{\rm Z}2}^{\alpha}=u_{2}^{\alpha}u_{4}^{\alpha}-v_{2}^{\alpha}v_{4}^{\alpha},\\ &S_{{\rm Z}1}^{\alpha}=u_{1}^{\alpha}v_{3}^{\alpha}+v_{1}^{\alpha}u_{3}^{\alpha},\quad S_{{\rm Z}2}^{\alpha}=u_{2}^{\alpha}v_{4}^{\alpha}+v_{2}^{\alpha}u_{4}^{\alpha}.\end{split}$ (34)
By minimizing $F_{\rm X}$ with respect to the X points fields, we obtain their stationary values as
$Q_{\ell,z}^{\alpha}=\frac{6b}{a_{\rm X}}\sum_{i=1}^{2}C^{\alpha}_{{\ell}i},\quad Q^{\alpha}_{\ell,x}=-\frac{6b}{a_{\rm X}}\sum_{i=1}^{2}S^{\alpha}_{{\ell}i}.$ (35)
Substituting them into Eq. (32), we obtain up to the linear order in $\phi_{0}$: $\delta F=\sum_{\alpha={\rm A,B}}\delta F^{\alpha}_{0}+\delta F^{\alpha}_{1}+\cdots$ as
$\delta F^{\alpha}_{0}=$ $-\frac{18b^{2}}{a_{\rm X}}\sum_{m>n}(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}$
$-\frac{36b^{2}}{a_{\rm X}}(c_{12}^{\alpha}c_{34}^{\alpha}+c_{13}^{\alpha}c_{24}^{\alpha}+c_{14}^{\alpha}c_{23}^{\alpha})\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4},$ (36)
and
$\delta F_{1}^{\alpha}=$ $-\frac{108b^{3}}{a_{\rm X}^{2}}\Bigg{\{}\sum_{m>n}\Big{[}u_{0}^{\alpha}\cos(2\theta^{\alpha}_{mn})+v_{0}^{\alpha}\sin(2\theta^{\alpha}_{mn})\Big{]}(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}$
$+6\Big{(}u_{0}^{\alpha}\cos\bar{\theta}^{\alpha}+v_{0}^{\alpha}\sin\bar{\theta}^{\alpha}\Big{)}\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4}\Bigg{\}},$ (37)
where $\theta_{mn}\equiv\theta_{m}+\theta_{n}$ and $\bar{\theta}^{\alpha}\equiv\theta_{1}^{\alpha}+\theta_{2}^{\alpha}+\theta_{3}^{\alpha}+\theta_{4}^{\alpha}$. Equation (36) renormalizes $c^{\prime}$ and $c^{\prime\prime\prime}$ as
$c^{\prime}\to c^{\prime}-\frac{72b^{2}}{a_{\rm X}},\quad c^{\prime\prime\prime}\to c^{\prime\prime\prime}-\frac{144b^{2}}{a_{\rm X}}.$ (38)
The terms appearing in $\delta F_{1}^{\alpha}$ are emergent fifth-order coupling constants between the mode at the $\Gamma$ and the L points. Since we have kept terms up to the fourth orders in the original free energy in Eq. (6), cases for large $108b^{3}/a_{\rm X}$ are beyond the applicable range of the approximation.
Appendix C Free energy for the orthogonal double-q configuration
[…]
$F=$ $\sum_{n=1,2}\Big{[}(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big{]}+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
$-6\lambda_{b}\phi_{0}(\phi_{1}^{2}-\phi_{2}^{2})\cos(2\theta+\theta_{0})$
$+2\lambda_{c}^{(2)}\phi_{0}^{2}\Big{[}\phi_{1}^{2}\cos^{2}(\theta-\theta_{0})+\phi_{2}^{2}\sin^{2}(\theta-\theta_{0})\Big{]}$
$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos(4\theta-\theta_{0})+F_{\Gamma},$ (39)
where $c^{\prime}=2c-72b^{2}/a_{\rm X},\ \lambda_{b}=b,\ \lambda_{c}^{(1)}=\lambda_{c}^{(2)}=c$, $F_{\Gamma}$ is given by Eq. (10). Let us briefly discuss possible stationary angle $\theta$ under magnetic fields. For simplicity, we assume $\theta_{0}=0$, since $\tilde{h}$ induces $u_{0}$ under both $\bm{H}\parallel[001]$ and $[110]$. We treat $\phi_{0,1,2}$ as given parameters and the stationary condition of $F$ with respect to $\cos(2\theta)$ leads to
$\cos(2\theta)=\frac{6\lambda_{b}-\lambda_{c}^{(2)}\phi_{0}}{216b^{3}/a_{\rm X}^{3}}\left(\frac{\phi_{1}^{2}-\phi_{2}^{2}}{\phi_{1}^{2}\phi_{2}^{2}}\right).$ (40)
This means that $\phi_{1}/\phi_{2}\to 1$ for $\theta\to\pi/4$ unless $\phi_{0}$ is tuned to $6\lambda_{b}/\lambda_{c}^{(2)}$, which is the phase III2q seen in the main text. The free energy for the phase III2q is
$F_{{\rm III}_{2q}}=$ $2\big{[}a_{\rm L}+J_{\rm L}+(\lambda_{c}^{(1)}+\lambda_{c}^{(2)})\phi_{0}^{2}\big{]}\phi_{1}^{2}$
$+\left(c+\frac{c^{\prime}}{2}-\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\cos\theta_{0}\right)\phi_{1}^{4}+F_{\Gamma}.$ (41)
One can also see that the deviation $\delta\phi=\phi_{1}-\phi_{2}\neq 0$ from the phase III2q accompanies changes in the angle as $\theta=\pi/4+\epsilon$. The quadratic form of $\delta\phi$ and $\epsilon$ describe the stability of the phase III2q and when one of the eigenvalues vanishes a second-order transition takes place toward the phase II2q.
For $\theta=0$, i.e., the phase I2q, the free energy is given by
$F_{\rm{I}_{2q}}=$ $\sum_{n=1,2}\Big{[}(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big{]}+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
$-6\lambda_{b}\phi_{0}(\phi_{1}^{2}-\phi_{2}^{2})\cos\theta_{0}$
$+2\lambda_{c}^{(2)}\phi_{0}^{2}\Big{[}\phi_{1}^{2}\cos^{2}\theta_{0}+\phi_{2}^{2}\sin^{2}\theta_{0}\Big{]}$
$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos\theta_{0}+F_{\Gamma}.$ (42)
Rotating $\theta$ leads to the transition to the phase II2q. Assuming $|\theta|\ll 1$ and $\theta_{0}=0$ in Eq. (39), we obtain the quadratic part of $F$ in $\theta$ as
$F=$ $\Big{[}12\lambda_{b}(\phi_{1}^{2}-\phi_{2}^{2})-\lambda_{c}^{(2)}\phi_{0}(\phi_{1}^{2}-2\phi_{2}^{2})-\frac{1728b^{3}}{a_{\rm X}^{2}}\phi^{2}_{1}\phi^{2}_{2}\Big{]}\phi_{0}\theta^{2}.$ (43)
Thus, a second-order transition to the phase II2q takes place when []delimited-[][$\cdots$][ ⋯ ] in Eq. (43) vanishes.

### v3

@@
 We study orbital orders in quadrupolar heavy-fermion compound PrIr2Zn20 under magnetic fields on the basis of the Landau theory.
-Assuming Eg orbital (electric quadrupolar) orders in the cubic symmetry with the ordering wavenumber at the L points in the face-centered cubic lattice Brillouin zone as observed experimentally, we construct the Landau free energy and analyze it.
-We find that the unidentified high-temperature ordered phase under the magnetic field $\bm{H}\parallel[001]$ reported earlier can be interpreted as the consequence of the internal rotation in the orbital moments of f electrons at the Pr site.
-We discuss the phase diagram for various magnetic-field directions and also possible double-$\bm{q}$ quadrupolar orders in this system.
+Assuming $E_{g}$ orbital (electric quadrupolar) orders in the cubic symmetry with the ordering wavenumber at the L points in the face-centered cubic lattice Brillouin zone as observed experimentally, we construct the Landau free energy and analyze it.
+We find that the unidentified high-temperature ordered phase under the magnetic field $\bm{H}\parallel[001]$ reported earlier can be interpreted as a consequence of the internal rotation in the orbital moments of f electrons at the Pr site.
+We also discuss the phase diagram for other magnetic-field directions and possible double-$\bm{q}$ quadrupolar orders in this system.
 pacs:
@@
 II Model
-We start to introduce the exchange interactions between the Eg orbital moments in Sec.
+We start to introduce the exchange interactions between the $E_{g}$ orbital moments in Sec.
 II.1 and the Landau free energy for the order parameters at the L points in Sec.
@@
 II.1 Exchange interaction
-Let us denote Eg quadrupole moments at the position $\bm{r}=(x,y,z)$ in the diamond structure as $\bm{\phi}_{\bm{r}}^{\alpha}=(u_{\bm{r}}^{\alpha},v_{\bm{r}}^{\alpha})$, where $\alpha=$A or B is the sublattice index.
-We take one of the A sublattice site at the origin $\bm{r}=(0,0,0)$ and its pair B sublattice site in the unit cell at $\bm{r}+(1,1,1)/4$.
+Let us denote $E_{g}$ quadrupole moments at the position $\bm{r}=(x,y,z)$ in the diamond structure as $\bm{\phi}_{\bm{r}}^{\alpha}=(u_{\bm{r}}^{\alpha},v_{\bm{r}}^{\alpha})$, where $\alpha=A$ or $B$ is the sublattice index.
+We take one of the $A$ sublattice site at the origin $\bm{r}=(x,y,z)=(0,0,0)$ and its pair $B$ sublattice site in the unit cell at $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$, where we set the lattice constant $a=1$.
 $\bm{\phi}_{\bm{r}}^{\alpha}$ form a two-dimensional vector field as a classical variable.
-Here, $(u,v)$ represents $(u,v)\sim\{2z^{2}-x^{2}-y^{2},\sqrt{3}(x^{2}-y^{2})\}$, where $\bm{r}=(x,y,z)$ is the real space coordinate.
+The two components are $u$ $\sim$ $2z^{2}$--$x^{2}$--$y^{2}$ and $v$ $\sim$ $\sqrt{3}$($x^{2}$--$y^{2}$).
 The Fourier transform in the wavenumber ${\bm{k}}$ space is defined as
 ${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}=\begin{bmatrix}u^{\alpha}_{{\bm{k}}}\\ v^{\alpha}_{{\bm{k}}}\end{bmatrix}=\frac{1}{\sqrt{N}}\sum_{\bm{r}}e^{-i{\bm{k}}\cdot\bm{r}}\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}.$ (1)
-Here, the summation for ${\bm{r}}$ runs over the A-sublattice sites in the diamond structure and $N$ is the number of the unit cell.
-The exchange Hamiltonian in the diamond structure is in general written as
+Here, the summation for ${\bm{r}}$ runs over the $A$-sublattice sites and $N$ is the number of the unit cell.
+The exchange Hamiltonian is in general written as
 $H=\sum_{{\bm{k}}}[u^{A}_{-{\bm{k}}},v^{A}_{-{\bm{k}}},u^{B}_{-{\bm{k}}},v^{B}_{-{\bm{k}}}]\begin{bmatrix}H^{AA}_{{\bm{k}}}&H^{AB}_{{\bm{k}}}\\ {H^{AB}_{{\bm{k}}}}^{\dagger}&H^{AA}_{{\bm{k}}}\end{bmatrix}\begin{bmatrix}u^{A}_{{\bm{k}}}\\ v^{A}_{{\bm{k}}}\\ u^{B}_{{\bm{k}}}\\ v^{B}_{{\bm{k}}}\end{bmatrix},$ (2)
@@
 Since the exchange interactions at the L points are isotropic due to the symmetry, we retain such exchange interactions while neglect anisotropic ones.
-Although the latter exchange interactions are generally present and can influence the ordering vectors and the excitation energy of quadrupole wave, it is no effect on the transition temperature for the order at the L points.
-Thus, a model without anisotropic interactions can be a good approximation as a minimal model.
+Although the anisotropic ones are generally present and can influence the ordering vectors and the excitation energy of quadrupole wave, it has no effect on the transition temperature for the order at the L points.
+Thus, a model without anisotropic interactions can be a good starting model as a minimal one.
 In this section, we consider exchange interactions up to the fourth neighbor interactions, while we will simplify the model in the later sections.
-Hereafter, the lattice constant $a$ is set to $a=1$.
-The next nearest-neighbor and the fourth-neighbor bonds are those between the A-A (B-B) sublattices. They are
-$H^{AA}_{{\bm{k}}}$ $=[J_{2}\gamma_{2}({\bm{k}})+J_{4}\gamma_{4}({\bm{k}})]\hat{\sigma}_{0}$ (3)
-Here, $\hat{\sigma}_{0}={\rm diag}(1,1)$ and $\hat{\sigma}^{x,z}$ is the Pauli matrix and we have followed the phase convention in Ref.
-[Hattori2014].
+The $2\times 2$ blocks in Eq. (2) are given by
+$H^{AA}_{{\bm{k}}}$ $=[J_{2}\gamma_{2}({\bm{k}})+J_{4}\gamma_{4}({\bm{k}})]\hat{1},$ (3)
+$H^{AB}_{{\bm{k}}}$ $=\frac{1}{2}[J_{1}\gamma_{1}({\bm{k}})+J_{3}\gamma_{3}({\bm{k}})]\hat{1}.$ (4)
+Here, $\hat{1}$ is a 2$\times$2 identity matrix.
 $J_{i}$’s are isotropic coupling constants.
 The expressions for $\gamma$’s are listed in Table 1.
-The nearest-neighbor and the third-neighbor bonds are those between the A-B sublattices.
-For the nearest-neighbor interaction, there is no anisotropic term owing to the symmetry.
-The form is thus given as
-$H^{AB}_{{\bm{k}}}$ $=\frac{1}{2}[J_{1}\gamma_{1}({\bm{k}})+J_{3}\gamma_{3}({\bm{k}})]\hat{\sigma}_{0}.$ (4)
-We note that $\gamma_{1}$ and $\gamma_{3}$ are complex. From the neutron experimental data, the ordering patterns are consistent with those expected for $J_{1}>0$.
+Note that $\gamma_{1}({\bm{k}})$ and $\gamma_{3}({\bm{k}})$ are complex.
+From the neutron experimental data, the ordering patterns are consistent with those expected for $J_{1}>0$ Iwasa et al.
+(2017).
 For describing the symmetry breaking at the L point modulation within our model, it is important to check which interactions are important.
@@
 Table 1: exchange interaction coefficients.
-The values of $\gamma({\bm{k}})$’s at ${\bm{k}}=(0,0,0),(0,0,2\pi)$, and $(\pi,\pi,\pi)$ are also shown.
+The values of $\gamma({\bm{k}})$’s at ${\bm{k}}=(0,0,0),(0,0,2\pi)$, ${\bm{k}}_{1}=(\pi,\pi,\pi)$, and the other L points are also shown.
+Here, we take the unit cell convention of $A$: (0,0,0) and $B$: $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$ and thus, the ${\bm{k}}_{1}$ is apparently inequivalent to the remaining ${\bm{k}}_{2,3,4}$.
 $\eta_{{\bm{k}}}\equiv e^{i(k_{x}+k_{y}+k_{z})/4}$, $\zeta^{\pm}_{{\bm{k}}}\equiv\pm 2(\cos\frac{k_{x}}{2}+\cos\frac{k_{y}}{2}+\cos\frac{k_{z}}{2})-3$.
 The mean-field second-order transition temperatures $T_{c}$’s are $T_{c}=2|J_{1}-3J_{3}|+6J_{4}$ for the L points, $T_{c}=4|J_{1}+3J_{3}|-12J_{2}-6J_{4}$ for the $\Gamma$ points, and $T_{c}=4J_{2}-6J_{4}$ for the X points.
-notation expression $\Gamma(0,0,0)$ X(0,0,$2\pi$) L($\pi,\pi,\pi$)
-$\gamma_{2}({\bm{k}})$ $2\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{z}}{2}\cos\frac{k_{x}}{2}\right)$ $6$ $-2$ 00
-$\gamma_{4}({\bm{k}})$ $\cos k_{x}+\cos k_{y}+\cos k_{z}$ $3$ $3$ $-3$
-$\gamma_{1}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\right)$ $4$ 00 $-2$
-$\gamma_{3}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}\zeta^{+}_{{\bm{k}}}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\zeta^{-}_{{\bm{k}}}\right)$ $12$ 00 $6$
+notation expression $\Gamma(0,0,0)$ X(0,0,$2\pi$) L($\pi,\pi,\pi$) L(others)
+$\gamma_{2}({\bm{k}})$ $2\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{z}}{2}\cos\frac{k_{x}}{2}\right)$ $6$ $-2$ 00 00
+$\gamma_{4}({\bm{k}})$ $\cos k_{x}+\cos k_{y}+\cos k_{z}$ $3$ $3$ $-3$ $-3$
+$\gamma_{1}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\right)$ $4$ 00 $-2$ $2$
+$\gamma_{3}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}\zeta^{+}_{{\bm{k}}}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\zeta^{-}_{{\bm{k}}}\right)$ $12$ 00 $6$ $-6$
 II.2 Landau Expansion
@@
 II.3 L point free energy
-When squeeze out the L point components ${\boldsymbol{\phi}}^{\alpha}_{n}=(u_{n}^{\alpha},v_{n}^{\alpha})\equiv{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}(n=1,2,3,4)$ with the short handed notation, one finds
-$F_{\rm L}$ $=\frac{a_{\rm L}}{2}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}+J_{\rm L}\sum_{n}{\boldsymbol{\phi}}^{A}_{n}\cdot{\boldsymbol{\phi}}^{B}_{n}+\frac{c}{4}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{4}$
+We now concentrate on the L point components ${\boldsymbol{\phi}}^{\alpha}_{n}=(u_{n}^{\alpha},v_{n}^{\alpha})\equiv{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}(n=1,2,3,4)$ with the short handed notation, we find the part of the free energy consisting of only ${\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}$’s is
+$F_{\rm L}$ $=\frac{a_{\rm L}}{2}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}+J_{\rm L}\sum_{n}(-1)^{p_{n}}{\boldsymbol{\phi}}^{A}_{n}\cdot{\boldsymbol{\phi}}^{B}_{n}+\frac{c}{4}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{4}$
 $+\frac{c^{\prime}}{4}\sum_{\alpha}\sum_{m<n}|{\boldsymbol{\phi}}^{\alpha}_{m}|^{2}|{\boldsymbol{\phi}}^{\alpha}_{n}|^{2}+\frac{c^{\prime\prime}}{4}\sum_{\alpha}\sum_{m<n}({\boldsymbol{\phi}}_{m}^{\alpha}\cdot{\boldsymbol{\phi}}_{n}^{\alpha})^{2}$
-$+\frac{c^{\prime\prime\prime}}{4}\sum_{\alpha}\sum_{m<n}\sum_{m^{\prime}<n^{\prime}}({\boldsymbol{\phi}}^{\alpha}_{m}\cdot{\boldsymbol{\phi}}^{\alpha}_{n})({\boldsymbol{\phi}}^{\alpha}_{m^{\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{n^{\prime}}).$ (6)
+$+\frac{c^{\prime\prime\prime}}{4}\sum_{\alpha}\sum_{(m,n,m^{\prime},n^{\prime})}({\boldsymbol{\phi}}^{\alpha}_{m}\cdot{\boldsymbol{\phi}}^{\alpha}_{n})({\boldsymbol{\phi}}^{\alpha}_{m^{\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{n^{\prime}}).$ (6)
 Note that $b$ terms are absent owing to the wavnumber conservation and $(m,n,m^{\prime},n^{\prime})=(1,2,3,4)$, $(1,3,2,4)$, and $(1,4,2,3)$ in the last term in the right-hand side.
+The factor $p_{n}(n=1,2,3,4)$ represents the sign of exchange interaction between the $A$ and $B$ sublattices for four kinds of the L points.
+Since our unit cell consists of the bond parallel to ${\bm{k}}_{1}$, $p_{n}=0$ is distinct from the others $p_{2,3,4}=1$.
 The coupling constants are denoted as
@@
 For discussing a single-$\bm{q}$ order at the L points, $c^{\prime}$, $c^{\prime\prime}$, and $c^{\prime\prime\prime}$ terms are irrelevant and the fourth-order free energy is simply given by
-$F_{L}^{1(4)}=\frac{c}{4}\sum_{\alpha}|{\boldsymbol{\phi}}^{\alpha}|^{4}.$ (9)
+$F_{L}^{1\bm{q}(4)}=\frac{c}{4}\sum_{\alpha}|{\boldsymbol{\phi}}^{\alpha}|^{4}.$ (9)
 When multiple-$\bm{q}$ states are considered, the coupling constants $c^{\prime}$, $c^{\prime\prime}$, and $c^{\prime\prime\prime}$ matter as will be discussed in Sec. V.
 II.4 $\Gamma$ point free energy
-Under finite magnetic fields, the uniform quadrupole moment ${\boldsymbol{\phi}}_{0}^{\alpha}\equiv{\boldsymbol{\phi}}_{{\bm{k}}=\bf 0}^{\alpha}$ is induced via the effective field $\tilde{\bm{h}}$ in the second order perturbation of the magnetic field $\bm{H}$.
+Under finite magnetic fields, the uniform quadrupole moment ${\boldsymbol{\phi}}_{0}^{\alpha}\equiv{\boldsymbol{\phi}}_{{\bm{k}}=\bf 0}^{\alpha}$ is induced via the effective field $\tilde{\bm{h}}$ in the second-order perturbation of the magnetic field $\bm{H}$.
 The free energy up to the fourth-order terms for the $\Gamma$ point modes are
-$F_{\Gamma}$ $=\sum_{\alpha}\Big{(}-\tilde{\bm{h}}\cdot{\boldsymbol{\phi}}_{0}^{\alpha}+\frac{a_{\Gamma 0}}{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}\Big{)}+J_{\Gamma 0}{\boldsymbol{\phi}}_{0}^{\rm A}\cdot{\boldsymbol{\phi}}_{0}^{\rm B},$
+$F_{\Gamma}$ $=\sum_{\alpha}\Big{(}-\tilde{\bm{h}}\cdot{\boldsymbol{\phi}}_{0}^{\alpha}+\frac{a_{\Gamma 0}}{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}\Big{)}+J_{\Gamma 0}{\boldsymbol{\phi}}_{0}^{A}\cdot{\boldsymbol{\phi}}_{0}^{B}$
 $+\sum_{\alpha}\Big{\{}-bu_{0}^{\alpha}\Big{[}\big{(}u_{0}^{\alpha}\big{)}^{2}-3\big{(}v_{0}^{\alpha}\big{)}^{2}\Big{]}+\frac{c}{4}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{4}\Big{\}},$ (10)
 where
-$a_{\Gamma 0}$ $=T+12J_{2}+6J_{4},\quad J_{\Gamma 0}=4J_{1}+12J_{3},$ (11)
-and the effective field is given by
+$a_{\Gamma 0}$ $=T+12J_{2}+6J_{4},\quad J_{\Gamma 0}=4J_{1}+12J_{3}.$ (11)
+The effective field is given by
 $\tilde{\bm{h}}=\begin{bmatrix}\tilde{h}_{u}\\ \tilde{h}_{v}\end{bmatrix}\equiv g\begin{bmatrix}2H_{z}^{2}-H_{x}^{2}-H_{y}^{2}\\ \sqrt{3}(H_{x}^{2}-H_{y}^{2})\end{bmatrix},$ (12)
-and $g=(g_{J}\mu_{\rm B})^{2}[7/(3E_{4})-1/E_{5}]$ given by the excited $\Gamma_{4,5}$ CEF energy $E_{4,5}$.Hattori and Tsunetsugu (2014) $\mu_{\rm B}$ is the Bohr magneton and $g_{J}$ is the Lande’s $g$ factor for the multiplet with the angular momentum $J$; $g_{J}=4/5$ for the PrIr2Zn20 with $J=4$.
+and $g=(g_{J}\mu_{\rm B})^{2}[7/(3E_{4})-1/E_{5}]$ with the excited $\Gamma_{4,5}$ CEF energy $E_{4,5}$.Hattori and Tsunetsugu (2014) $\mu_{\rm B}$ is the Bohr magneton and $g_{J}$ is the Lande’s $g$ factor for the multiplet with the angular momentum $J$; $g_{J}=4/5$ for the PrIr2Zn20 with $J=4$.
 II.5 Mode coupling terms
-We now discuss mode coupling terms consisting of the fields at the L points and those at $\Gamma$ points. The lowest order contributions come from the cubic terms as
+We now discuss mode coupling terms consisting of the fields at the L points and those at the $\Gamma$ points. The lowest-order contributions come from the cubic terms as
 $\delta F_{3}=$ $-3\lambda_{b}\sum_{\alpha,n}\Big{\{}u_{0}^{\alpha}\Big{[}\big{(}u_{n}^{\alpha}\big{)}^{2}-\big{(}v_{n}^{\alpha}\big{)}^{2}\Big{]}-v_{0}^{\alpha}(2u_{n}^{\alpha}v_{n}^{\alpha})\Big{\}}$
-$=-3\lambda_{b}\sum_{\alpha,n}\phi_{0}^{\alpha}\big{(}\phi_{n}^{\alpha}\big{)}^{2}\cos(\theta^{\alpha}_{0}+2\theta^{\alpha}_{n}),$ (13)
-Here we have parameterized
+$=-3\lambda_{b}\sum_{\alpha,n}\phi_{0}^{\alpha}\big{(}\phi_{n}^{\alpha}\big{)}^{2}\cos(\theta^{\alpha}_{0}+2\theta^{\alpha}_{n}).$ (13)
+Here, we have parameterized
 ${\boldsymbol{\phi}}^{\alpha}_{\bf 0}$ $=\phi^{\alpha}_{0}\begin{bmatrix}\cos\theta^{\alpha}_{0}\\ \sin\theta^{\alpha}_{0}\end{bmatrix},\quad{\boldsymbol{\phi}}^{\alpha}_{n}=\phi^{\alpha}_{n}\begin{bmatrix}\cos\theta^{\alpha}_{n}\\ \sin\theta^{\alpha}_{n}\end{bmatrix},$ (14)
-and the coupling constant $\lambda_{b}$ as an independent parameter. If $\lambda_{b}$ is estimated by Eq. (5), $\lambda_{b}=b$. The next leading order terms are the fourth order and given by
+and the coupling constant $\lambda_{b}$ has been introduced as an independent parameter.
+If $\lambda_{b}$ is estimated by Eq.
+(5), $\lambda_{b}=b$.
+The next leading order terms are in the fourth order and given by
 $\delta F_{4}=\sum_{\alpha,n}\left[\frac{\lambda^{(1)}_{c}}{2}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}+\lambda_{c}^{(2)}({\boldsymbol{\phi}}_{n}^{\alpha}\cdot{\boldsymbol{\phi}}_{0}^{\alpha})^{2}\right].$ (15)
-Again, the coupling constant $\lambda_{c}^{(1,2)}$ has been introduced as an independent parameter. If $\lambda_{c}^{(1,2)}$ is estimated by Eq. (5), $\lambda^{(1)}_{c}=\lambda_{c}^{(2)}=c$.
-In principle, ${\boldsymbol{\phi}}_{0}^{\alpha}$ can couple with ${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}$ in the form of $\sim{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$, ${\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$.
-In this paper, we retain the order parameter ${\boldsymbol{\phi}}_{n}^{\alpha}$ and the induced moment ${\boldsymbol{\phi}}_{0}^{\alpha}$ under magnetic fields since the other fields are not important for the phase transition.
-Some renormalization effects arising from the field at the X points will be discussed in Sec.
+Again, the coupling constant $\lambda_{c}^{(1)}$ and $\lambda_{c}^{(2)}$ have been introduced as independent parameters.
+If $\lambda_{c}^{(1,2)}$ is estimated by Eq.
+(5), $\lambda^{(1)}_{c}=\lambda_{c}^{(2)}=c$.
+In principle, ${\boldsymbol{\phi}}_{0}^{\alpha}$ can couple with ${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}$ in the form of $\sim{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$ and ${\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$.
+In this paper, we retain the order parameter ${\boldsymbol{\phi}}_{n}^{\alpha}$ and the induced moment ${\boldsymbol{\phi}}_{0}^{\alpha}$ under magnetic fields since the other fields are not important for the phase transition with the L point fields.
+Some renormalization effects arising from fields at the X points will be discussed in Sec.
 IV and Appendix B.
@@
 In this section, we carry out minimization of the free energy $F_{L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}$ [Eqs.
-(6), (10), (13), and (13)] with the single-$\bm{q}$ assumptions.
+(6), (10), (13), and (15)], assuming the single-$\bm{q}$ orders.
 The purpose here is to understand how mode coupling effects affect the order parameter configuration.
 The results in this section will become bases for understanding the full phase diagram shown in Sec.
-IV where we sill take into account multiple-$\bm{q}$ configurations.
-Since we are interested in the phase transition owing to order parameter at the L point, we assume the second-order coefficients of the free energy at the other wavenumber are always positive.
-We first discuss the candidates of the stable configurations by analyzing the third and fourth order terms in $F_{\rm L}$.
+IV, where we will take into account multiple-$\bm{q}$ configurations.
+Since we are interested in the phase transition owing to the order parameters at the L point, we assume the coefficients of the quadratic terms in the free energy at the other wavenumber are always positive.
+We first discuss the candidates of the stable configurations by analyzing the third- and fourth-order terms in $F_{\rm L}$.
 Then, we will show the numerical results of full temperature-field phase diagram.
@@
 Thus, it is legitimate to set
-${\boldsymbol{\phi}}_{\rm L}\equiv{\boldsymbol{\phi}}_{n}^{A}={\boldsymbol{\phi}}_{n}^{B}=\begin{bmatrix}u_{\rm L}\\ v_{\rm L}\end{bmatrix}=\phi_{\rm L}\begin{bmatrix}\cos\theta_{\rm L}\\ \sin\theta_{\rm L}\end{bmatrix}.$ (16)
-Here, we have chosen one of the L points at ${\bm{k}}_{n}$ as a representative one since the four L points are equivalent even when $\bm{h}$ is finite.
+${\boldsymbol{\phi}}_{\rm L}\equiv{\boldsymbol{\phi}}_{1}^{A}={\boldsymbol{\phi}}_{1}^{B}=\begin{bmatrix}u_{\rm L}\\ v_{\rm L}\end{bmatrix}=\phi_{\rm L}\begin{bmatrix}\cos\theta_{\rm L}\\ \sin\theta_{\rm L}\end{bmatrix}.$ (16)
+Here, we have chosen one of the L points at ${\bm{k}}_{1}$ as a representative one since the other three fields at ${\bm{k}}_{n}(n=2,3,4)$ with ${\boldsymbol{\phi}}_{n}^{A}=-{\boldsymbol{\phi}}_{n}^{B}$ are equivalent and this is also the case even when $\bm{H}$ is finite.
+See Table 1.
 For the modes at the $\Gamma$ point, which are gapped and just induced by magnetic fields or by the square of the order parameter at the L point, we set
 ${\boldsymbol{\phi}}_{0}\equiv{\boldsymbol{\phi}}_{0}^{A}={\boldsymbol{\phi}}_{0}^{B}=\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}=\phi_{0}\begin{bmatrix}\cos\theta_{0}\\ \sin\theta_{0}\end{bmatrix}.$ (17)
-Second, we note that the single-$\bm{q}$ free energy $F_{L}$ for the L point order parameter is isotropic.
-This means there is no anisotropy at the transition temperature $T_{c}=6J_{4}+2|J_{1}-3J_{3}|$ at ${\bm{H}}={\bf 0}$.
-This isotropic nature, however, is destroyed by the mode coupling to the fields at the $\Gamma$ points through $\delta F_{3}$ and $\delta F_{4}$.
-When $T\lesssim T_{c}$, the induced ${\boldsymbol{\phi}}_{0}$ is sufficiently small, and thus the third and fourth order terms in $F_{\Gamma}$ can be irrelevant.
+Second, we note that the single-$\bm{q}$ free energy $F_{\rm L}$ is isotropic.
+This means that the transition temperatures for $u_{\rm L}$ and $v_{\rm L}$ are the same at ${\bm{H}}={\bf 0}$ and $T_{c}=6J_{4}+2|J_{1}-3J_{3}|$ This isotropic nature, however, is destroyed by the mode coupling to the fields at the $\Gamma$ points through $\delta F_{3}$ and $\delta F_{4}$.
+When $T\lesssim T_{c}$, the induced ${\boldsymbol{\phi}}_{0}$ is sufficiently small, and thus, the third- and fourth-order terms in $F_{\Gamma}$ can be irrelevant.
 In this situation the relative directions between ${\boldsymbol{\phi}}_{n}$ and ${\boldsymbol{\phi}}_{0}$ is determined by $\delta F_{3}$ and $\delta F_{4}$ as
@@
 Since ${\boldsymbol{\phi}}_{0}$ is not the primary order parameter, the third-order term is expected to play a minor role and the configurations (19) are robust at least $\tilde{\bm{h}}={\bf 0}$.
-Figure 2: Schematic single-$\bm{q}$ order parameter configurations for (a) ${\bm{H}=0}$, (b) small ${\bm{H}}$ and low $T$, (c) small ${\bm{H}}$ and high $T$, and (d) large ${\bm{H}}$.
-(e) large ${\bm{H}}$.
+Figure 2: Schematic single-$\bm{q}$ order parameter configurations for (a) ${\bm{H}=0}$, (b) small ${\bm{H}}$, (c) small ${\bm{H}}$ and low $T$, and (d) large ${\bm{H}}$.
 Blue (red) arrows indicate ${\boldsymbol{\phi}}_{0}$(${\boldsymbol{\phi}}_{\rm L}$).
-Filled red circles represent the directions $\theta_{L}$ given by Eq.
+Filled red circles represent the directions $\theta_{\rm L}$ given by Eq.
 (19) and open blue circles do $\theta_{\rm 0}$.
-A configuration for large ${\bm{H}}$ in the phase II in (c) is shown in (e) and that for ${\bm{H}}\parallel[110]$ in (f).
+A configuration for large ${\bm{H}}$ in the phase II is shown in (e) and that for ${\bm{H}}\parallel[110]$ in (f).
 Now let us consider the cases for finite magnetic fields.
@@
 This is because $\delta F_{4}>\delta F_{3}$ for large $\tilde{h}_{u}$.
+See Eq.
+(18).
 By the similar analysis to Eqs.
@@
 We note that the collinear phase (phase I) appears for high $T$ and intermediate $H$.
-The reason why the phase I cannot appear at the lower $T$ is understood by seeing Eq.
+The reason why the phase I cannot appear at the lower $T$ is understood by examining Eq.
 (23); there is no solution of the critical field $\tilde{h}_{u}^{c1}$ when $\phi_{\rm L}^{2}$, which is approximately $H$ independent, is large.
@@
 Figure 4(a) shows the $T$-$H$ phase diagram for $\bm{H}\parallel[110]$.
-As discussed in the last part in the previous section, there is only one stable phase (phase IV), where $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$.
+As discussed in the last part in Sec.
+III.1, there is only one stable phase (phase IV), where $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$.
 Figures 4(b) and 4(c) show the real space quadrupole configurations with the ordering vector $\bm{k}_{1}$ for $\bm{H}\parallel[001]$ and $\bm{H}\parallel[110]$, respectively.
-For both cases, the quadrupole orders with $\bm{\phi}_{A}=\bm{\phi}_{B}$ are ferroic on a (111) plane, while antiferroic between the neighboring (111) plane as shown in different colors in Fig.
+For both cases, the quadrupole orders with $\bm{\phi}_{A}=\bm{\phi}_{B}$ are ferroic on the (111) plane, while antiferroic between the neighboring (111) planes as shown in different colors in Fig.
 4.
 When the ordering vector is $\bm{k}_{n}(n=2,3,4)$, the quadrupole configurations are ferroic on the plane perpendicular to ${\bm{k}}_{n}$ with $\bm{\phi}_{A}=-\bm{\phi}_{B}$.
-Figure 4: (a) $T$$H$ phase diagram for $\bm{H}\parallel[110]$, $J_{1}=-1$, $J_{2}=J_{3}=0$, $J_{4}=1$, $c=0.5$, and $b=0.05$.
+Figure 4: (a) $T$$H$ phase diagram for $\bm{H}\parallel[110]$, $J_{1}=1.0$, $J_{2}=J_{3}=0$, $J_{4}=1$, $c=0.5$, and $b=0.05$.
 Color map represents the L point order parameter angle $\theta_{\rm L}$.
-(b) $\theta$$H$ phase diagram from [001]($\theta=0$) to [110] ($\theta=\pi/2$).
+Single-$\bm{q}$ order with ${\bm{k}}={\bm{k}}_{1}$ for (b) the phase I and (c) the phase IV.
 IV Results: multiple-q orders
@@
 IV.1 Free energy for symmetric states
-In this section, we consider double-$\bm{q}\ (\mathcal{N}=2)$ and quadruple-$\bm{q}\ (\mathcal{N}=4)$ configurations, whose real space moments are given by
+We consider double-$\bm{q}\ (\mathcal{N}=2)$ and quadruple-$\bm{q}\ (\mathcal{N}=4)$ configurations, whose real space moments are given by
 ${\boldsymbol{\phi}}^{\alpha}(\bm{r})=\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}=\frac{\Phi_{\alpha}}{\sqrt{\mathcal{N}}}\sum_{n=1}^{\mathcal{N}}e^{i{\bm{k}}_{n}\cdot\bm{r}}\begin{bmatrix}\cos\theta_{n}^{\alpha}\\ \sin\theta_{n}^{\alpha}\\ \end{bmatrix},$ (25)
-where $\bm{r}$ is labeled by the A sublattice position as before and we have assumed that the magnitude for each ${\bm{k}}_{n}$ is the same for simplicity (symmetric states).
+where $\bm{r}$ is labeled by the $A$ sublattice position as before and we have assumed that the magnitude for each ${\bm{k}}_{n}$ is the same $\Phi_{\alpha}/\sqrt{\mathcal{N}}(>0)$ for simplicity (symmetric states).
+For $n=1$ $\bm{\phi}^{A}_{1}=\bm{\phi}^{B}_{1}$, while, for $n=2,3$ and $4$, $\bm{\phi}^{A}_{n}=-\bm{\phi}^{B}_{n}$.
+As mentioned before, this is due to our unit cell convention.
 In this normalization $(\mathcal{N})$, the quadratic part is unchanged from the single-$\bm{q}$ expression.
@@
 For double-$\bm{q}$ states, the fourth-order terms are
-$F_{\rm L}^{2(4)}=\sum_{\alpha}\Phi_{\alpha}^{4}\left[\frac{c}{8}+\frac{c^{\prime}}{16}+\frac{c^{\prime\prime}}{16}(c^{\alpha}_{12})^{2}\right].$ (26)
+$F_{\rm L}^{2\bm{q}(4)}=\sum_{\alpha}\Phi_{\alpha}^{4}\left[\frac{c}{8}+\frac{c^{\prime}}{16}+\frac{c^{\prime\prime}}{16}(c^{\alpha}_{12})^{2}\right].$ (26)
 Here, $c^{\alpha}_{mn}\equiv\cos(\theta^{\alpha}_{m}-\theta^{\alpha}_{n})$.
@@
 For quadruple-$\bm{q}$ states,
-$F_{\rm L}^{4(4)}=$ $\sum_{\alpha}\Phi_{\alpha}^{4}\Bigg{[}\frac{c}{16}+\frac{3c^{\prime}}{32}+\frac{c^{\prime\prime}}{64}\sum_{m<n}(c^{\alpha}_{mn})^{2}$
-$+$ $\frac{c^{\prime\prime\prime}}{64}(c_{12}^{\alpha}c_{34}^{\alpha}+c_{13}^{\alpha}c_{24}^{\alpha}+c_{14}^{\alpha}c_{23}^{\alpha})\Bigg{]}.$ (27)
-Note that the $c^{\prime\prime\prime}$ term can be negative, e.g., when $\theta_{1}=\theta_{2}=\theta_{3}=\theta_{4}+\pi$.
-This means that when $c^{\prime\prime\prime}$ is large, such quadruple-$\bm{q}$ orders are favored.
-The lowest free energy under the constraint (8) is $F_{\rm L}^{4(4)}=\tfrac{c}{4}\sum_{\alpha}\Phi_{\alpha}^{4}$ with $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(\theta,\theta,\theta^{\prime},\theta^{\prime}+\pi)$, where $\theta$ and $\theta^{\prime}$ are arbitrary and those with any two $\theta_{n}$’s permutations are equivalent ones.
-The unit cell for the quadrupole-$\bm{q}$ configurations contains eight A-sublattice sites and thus sixteen sites in total.
+$F_{\rm L}^{4\bm{q}(4)}=$ $\sum_{\alpha}\Phi_{\alpha}^{4}\Bigg{[}\frac{c}{16}+\frac{3c^{\prime}}{32}+\frac{c^{\prime\prime}}{64}\sum_{m<n}(c^{\alpha}_{mn})^{2}\Bigg{]}.$ (27)
+Note that the $c^{\prime\prime\prime}$ terms cancel when summing up the terms for $\alpha=A$ and $B$, since $\phi^{A}_{n}=(-1)^{p_{n}}\phi^{B}_{n}$.
+The minimum value of $F_{\rm L}^{4\bm{q}(4)}$ for $c^{\prime\prime}>0$ is realized with $\sum_{m<n}(c^{\alpha}_{mn})^{2}=2$, where two orthogonal pairs with the relative angle between the two is arbitrary, e.g., $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})=(\theta^{\alpha}+\pi/2,\theta^{\alpha},\varphi^{\alpha}+\pi/2,\varphi^{\alpha})$.
+For $c^{\prime\prime}<0$, the configurations for the minimum free energy is, e.g., $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})=(\theta,\theta,\theta,\theta)$ with $\sum_{m<n}(c^{\alpha}_{mn})^{2}=6$.
+The unit cell for the quadrupole-$\bm{q}$ configurations contains eight $A$-sublattice sites with sixteen sites in total.
 See also Fig.
 9 in Appendix A.
-In Eqs.
-(26) and (27), $\theta_{n}$ dependence appears only in $c^{\prime\prime}$ and $c^{\prime\prime\prime}$ terms.
-Figure 6 shows the phase diagram at $T=T_{c}$ on the $c^{\prime\prime\prime}$$c^{\prime\prime}$ plane for (a) $c^{\prime}>2c$ and (b) $c^{\prime}<2c$ with $c>0$.
-For large $c^{\prime}$, multiple-$\bm{q}$ states are not realized as is evident from the expressions (26) and (27).
-See Fig.
-6(a).
+Figure 6 shows the phase diagram at $T=T_{c}$ on the $c^{\prime\prime}$$c^{\prime}$ plane for $c>0$.
+Note that there is no $c^{\prime\prime\prime}$ dependence.
+For large $c^{\prime}$, multiple-$\bm{q}$ states are not realized as is evident from Eqs.
+(26) and (27).
 The conditions given by Eq.
-(8) just correspond to the phase boundary for $c^{\prime\prime},c^{\prime\prime\prime}>0$ and $c^{\prime}=2c$, where all the single-$\bm{q}$, double-$\bm{q}$, and quadruple-$\bm{q}$ are degenerate.
-Note that away from the phase boundaries between quadruple-$\bm{q}$ and double- or single-$\bm{q}$ phases, the direction of the $\bm{\phi}_{n}$ in the quadruple-$\bm{q}$ states are pinned to $(\theta_{1}^{\alpha},\theta_{2}^{\alpha},\theta_{3}^{\alpha},\theta_{4}^{\alpha})$$=$(0,0,0,0)$0$(0,0,0,0)( 0 , 0 , 0 , 0 ), $(0,0,\pi,\pi)$, $(0,\pi,\pi,0)$, $(0,\pi,0,\pi)$ for $c^{\prime\prime\prime}<0$, while $(\theta_{1}^{\alpha},\theta_{2}^{\alpha},\theta_{3}^{\alpha},\theta_{4}^{\alpha})$$=$$(0,0,0,\pi)$, $(0,0,\pi,0)$, $(0,\pi,0,0)$, $(\pi,0,0,0)$ for $c^{\prime\prime\prime}>0$.
-Here, global sign changes $\bm{\phi}^{\alpha}\to-\bm{\phi}^{\alpha}$: $\theta_{n}^{\alpha}\to\theta_{n}^{\alpha}+\pi$, have not been included since they are also realized by trivial translations.
-Figure 5: Phase diagram at $T=T_{c}$ in the $c^{\prime\prime\prime}$$c^{\prime\prime}$ plane.
-(a) $c^{\prime}>2c$, where the phase boundary between the single-$\bm{q}$(1q) and quadruple-$\bm{q}$(4q) is given by $c^{\prime\prime}=|c^{\prime\prime\prime}|/2-(c^{\prime}-2c)$.
-(b) $c^{\prime}<2c$, where the phase boundary between the double-$\bm{q}$(2q) and quadruple-$\bm{q}$(4q) is given by $c^{\prime\prime}=|c^{\prime\prime\prime}|/2+(2c-c^{\prime})/3$.
+(8) just correspond to the point $(c^{\prime}/c,c^{\prime\prime}/c)=(2,4)$, where the single-$\bm{q}$ and double-$\bm{q}$ are degenerate.
+As long as the local fourth-order terms are dominant, quadrupole-$\bm{q}$ orders are unlikely.
+Thus, the $c^{\prime}$ term is important when one compares the free energy for the single- and double-$\bm{q}$ orders.
+Figure 5: Phase diagram at $T=T_{c}$ in the $c^{\prime}$$c^{\prime\prime}$ plane for $c>0$.
+Filled circle indicate the bare parameter set $c^{\prime}=2c$ and $c^{\prime\prime}=4c$ given by Eq.
+(8).
+Region for $c^{\prime}+c^{\prime\prime}<-2c$ is unphysical since the fourth-order term becomes negative.
+For quadruple-$\bm{q}$ states, their representative configurations for $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})$ are indicated.
 So far, we have treated the fourth-order coefficients $c$’s as a phenomenological parameters.
@@
 Among them, there are special terms which connect two L-point modes at $\bm{q}_{m}$ and $\bm{q}_{n}$ ($m\neq n$) via $\sim\phi_{\bm{k}}\phi_{m}\phi_{n}$, where ${\bm{k}}$ is one of the wavenumbers at the X points $\bm{q}_{\rm X}=(2\pi,0,0),\bm{q}_{\rm Y}=(0,2\pi,0)$, or $\bm{q}_{\rm Z}=(0,0,2\pi)$.
-The combination of $({\bm{k}},\bm{q}_{m},\bm{q}_{n})$ is restricted by the Umklapp condition ${\bm{k}}+\bm{q}_{m}+\bm{q}_{n}$ and there are two sets of $(m,n)$ for a given ${\bm{k}}$.
+The combination of $({\bm{k}},\bm{q}_{m},\bm{q}_{n})$ is restricted by the Umklapp condition ${\bm{k}}+\bm{q}_{m}+\bm{q}_{n}\equiv\bf{0}$ and there are two sets of $(m,n)$ for a given ${\bm{k}}$.
 As shown in Appendix B, these couplings lead to suppression of the coefficients $c^{\prime}$ and $c^{\prime\prime\prime}$.
@@
 (38).
-Furthermore, nontrivial couplings between the $\Gamma$ and the L points fields appear after integrating out the X point fields as $\sim\phi_{0}\phi_{m}^{2}\phi_{n}^{2}$ and $\sim\phi_{0}\phi_{1}\phi_{2}\phi_{3}\phi_{4}$.
+Furthermore, the fifth-order couplings between the $\Gamma$ and the L points fields appear after integrating out the X point fields as $\sim\phi_{0}\phi_{m}^{2}\phi_{n}^{2}$ and $\sim\phi_{0}\phi_{1}\phi_{2}\phi_{3}\phi_{4}$.
 They play an important role under magnetic fields $\bm{H}$ since $\phi_{0}$ is induced by $\bm{H}$, and thus, these terms renormalize $c$’s with varying their magnitude depending on $\bm{H}$.
@@
 However, the coupling between the fields at the $\Gamma$ and the L points possesses an energy gain for the single-$\bm{q}$ orders and double-$\bm{q}$ ones with $\phi_{m}\neq\phi_{n}$.
-This is because there is a finite $\phi_{0}$ induced in the single-$\bm{q}$ states, while there is no gain for the double-$\bm{q}$ states with $\phi_{n}=\phi_{m}$ with $\theta_{n}-\theta_{m}=\pm\pi/2$.
+This is because there is a finite $\phi_{0}$ induced in the single-$\bm{q}$ states, while there is no gain for the double-$\bm{q}$ states with $\phi_{n}=\phi_{m}$ and $\theta_{n}-\theta_{m}=\pm\pi/2$.
 See the coupling in Eq.
@@
 As we will see in Sec.
-IV.2, the nontrivial terms generated by the integration of the X points fields choose some particular $\theta$ among degenerate orthogonal double-$\bm{q}$ configurations.
+IV.2, the fifth-order terms generated by the integration of the X points fields choose some particular $\theta$ among degenerate orthogonal double-$\bm{q}$ configurations.
 IV.2 Phase diagram
@@
 (6), (10), (13), and (15).
-Figures 6 and 8 show phase diagrams for $\bm{H}\parallel[001]$, i.e., $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\geq 0)$, and double-$\bm{q}$ configurations in the quadrupole space and real space.
+For simplify the analysis, we assume ${\boldsymbol{\phi}}_{n}^{B}={\boldsymbol{\phi}}_{n}^{A}(-1)^{p_{n}}$ valid for $J_{\rm L}<0$ and do not take into account quadruple-$\bm{q}$ orders, since their energy is much higher than single- and double-$\bm{q}$ states for the conventional fourth-order terms [Eq.
+(8)] as shown in Fig.
+6.
+Figures 6 and 7 show phase diagrams for $\bm{H}\parallel[001]$, i.e., $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\geq 0)$, and those for $\bm{H}\parallel[110]$ with $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\leq 0)$, respectively.
+The representative double-$\bm{q}$ configurations in the quadrupole space and real space are also illustrated in Fig.
+8.
+Figure 6: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, and $c=0.5$.
+(a) $(b,J_{2},J_{3},J_{4})=(0.1,0.0,0.0,1.0)$, (b) $(0.05,0.0,0.0,1.0)$, and (c) $(0.03,0.5,0.3,0.4)$.
+Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, where double-$\bm{q}$ orders with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$ are assumed.
+For single-$\bm{q}$ orders, we use the convention of $\phi_{2}=0$.
+Thick lines represent phase boundaries, where “1st” represents a line of first-order phase transition.
+Dashed lines are boundaries among double-$\bm{q}$ configurations.
+Low $T$ regime is not shown since the Landau expansion is not valid in general.
 Figure 6(a) shows the phase diagram for $J_{1}=1$, $c=0.5$, and $b=0.1$.
 The exchange coupling constants are the same as used in Fig.
-3: $(J_{2},J_{3},J_{4})=(0.0,0.0,1.0)$.
+3: $(J_{2},J_{3},J_{4})=(0,0,1)$.
 Except for the low-field region, the phase diagram is similar to those shown in Fig.
@@
 At ${\bm{H}}=0$, the orthogonal double-$\bm{q}$ states with $\phi_{1}=\phi_{2}$ discussed in Sec.
-IV.1 are realized and degenerate in the angle $\theta=0$.
-As ${\bm{H}}$ increases, $\phi_{1}$ becomes larger than $\phi_{2}$.
+IV.1 are realized.
+As $H$ increases, $\phi_{1}$ becomes larger than $\phi_{2}$.
 Since the phase I is the single-$\bm{q}$ order with $v_{\rm L}=0$, the symmetry breaking from the phase I to the phase I2q is caused by the emergence of $\bm{\phi}_{2}=(0,\phi_{2})$ through the second-order transition.
@@
 8(c) and 8(f), where $\bm{\phi}_{1}=\tfrac{1}{\sqrt{2}}\phi_{1}(1,1)$ and $\bm{\phi}_{2}=\tfrac{1}{\sqrt{2}}\phi_{2}(-1,1)$ with $\phi_{1}=\phi_{2}$.
-This configuration is stabilized by the nontrivial coupling discussed in the last part of Sec.
+This configuration is stabilized by the fifth-order coupling discussed in the last part of Sec.
 IV.1.
@@
 The overall phase diagram, however, is unchanged.
-As an example for more general cases, we show the phase diagram for $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ in Fig.
-6(c).
-For this parameter set, the second-order coefficient $a_{\rm X}$ of the X point fields almost vanish at $T=0$ as shown in Fig.
-1(c), the renormalization effects due to the X point fields grows rapidly at low temperatures.
-This causes the strong stabilization of the phase III2q in Fig.
-6(c), showing reentrant behavior as a function of $H$.
-Interestingly, there is a thin region of the phase II2q.
-The phase boundary between the phase II2q and II is of the second order.
-Similarly, that between the small region of phase I2q and III is also second order.
-We note that the boundaries indicated by the dashed lines are not symmetry breaking phase boundary since the symmetry of the three double-$\bm{q}$ phases are the same as is evident from the real space configurations in Figs.
-8(d)–8(f).
-See also the discussion in Appendix C.
-For $\bm{H}\parallel[110]$, a phase I2q appears at low field region as shown in Fig.
-7.
-In contrast to the phase I2q for $\bm{H}\parallel[001]$, the quadrupole configuration is $\bm{\phi}_{1}=(0,\phi_{1})$ and $\bm{\phi}_{2}=(\phi_{2},0)$ with $\phi_{1}>\phi_{2}$.
-This is natural since the neighboring single-$\bm{q}$ phase is the phase IV.
-The sharp increase of the critical field between I2q and IV phases are due to the fact that $a_{\rm X}$ is close to 0 in (c).
-Figure 6: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, and $c=0.5$.
-(a) $(b,J_{2},J_{3},J_{4})=(0.1,0.0,0.0,1.0)$, (b) $(0.05,0.0,0.0,1.0)$, and (c) $(0.03,0.5,0.3,0.4)$.
-Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, where double-$\bm{q}$ orders with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$ are assumed.
-For single-$\bm{q}$ orders, we use the convention of $\phi_{2}=0$.
-Thick lines represent phase boundaries, where “1st” represents a line of first-order phase transition.
-Dashed lines are boundaries among double-$\bm{q}$ configurations.
-Low $T$ regime is not shown since the Landau expansion is not valid in general.
 Figure 7: $T$$H$ phase diagram for $\bm{H}\parallel[110]$, $J_{1}=1.0$, and $c=0.5$.
@@
 Arrows indicate the order parameters $\bm{\phi}_{1}$ (L1), $\bm{\phi}_{2}$ (L2), and $\bm{\phi}_{0}$ $(\Gamma)$.
-Four small circles represent quadrupole moments at the A sublattice real-space positions $\bm{r}=(0,0,0),(1,0,0),(\tfrac{1}{2},\frac{1}{2},0)$, and $(\tfrac{1}{2},\frac{1}{2},1)$.
+Four small circles represent the quadrupole moments at the $A$ sublattice real-space positions $\bm{r}=(0,0,0),(1,0,0),(\tfrac{1}{2},\frac{1}{2},0)$, and $(\tfrac{1}{2},\frac{1}{2},1)$.
 Real space L point quadrupole configurations for (d) $(\theta_{1},\theta_{2})=(0,\pi/2)$ and $\phi_{1}=\sqrt{3}\phi_{2}$ in the phase I2q, (e) $(\theta_{1},\theta_{2})=(\pi/24,13\pi/24)$ and $\phi_{1}=1.2\phi_{2}$ in the phase II2q, and (f) $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$, and $\phi_{1}=\phi_{2}$ in the phase III2q.
 Note that only L point quadrupole components without those at the $\Gamma$ point are shown.
-Skyblue and gray arrows represent the directions of ${\bm{k}}_{1}$ and $-{\bm{k}}_{2}$.
 For guide to eyes, some of the nearest-neighbor bonds are connected by white bonds.
-Circles are the A sublattice atoms in the diamond structure and their colors represent the sublattice (red, green, blue, white).
+Circles are the $A$ sublattice atoms and their colors represent the sublattice (red, green, blue, white).
 For clarity, the $B$ sublattice atoms are not drawn.
-Colored bonds running [01¯$1]$ 1 ] directions connects the A sublattice sites with identical quadrupole moments.
+Colored bonds running [01¯$1]$ 1 ] directions connects the $A$ sublattice sites with the same quadrupole moments.
+As an example of more general cases, we show the phase diagram for $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ in Fig.
+6(c).
+For this parameter set, the second-order coefficient $a_{\rm X}$ of the X point fields almost vanish at $T=0$ as shown in Fig.
+1(c), the renormalization effects due to the X point fields grows rapidly at low temperatures.
+This causes the strong stabilization of the phase III2q in Fig.
+6(c), showing reentrant behavior as a function of $H$.
+Note, however, the Landau theory is not applicable to such low temperature region far away from $T\sim T_{c}$.
+There is a thin region of the phase II2q.
+The phase boundary between the phase II2q and II is of the second order.
+Similarly, that between the small region of phase I2q and III is also second order.
+We note that the boundaries indicated by the dashed lines are not symmetry breaking phase boundary since the symmetry of the three double-$\bm{q}$ phases are the same as is evident from the real space configurations in Figs.
+8(d)–8(f).
+See also the discussion in Appendix C.
+For $\bm{H}\parallel[110]$, a phase I2q appears at low-field region as shown in Fig.
+7.
+In contrast to the phase I2q for $\bm{H}\parallel[001]$, the quadrupole configuration is $\bm{\phi}_{1}=(0,\phi_{1})$ and $\bm{\phi}_{2}=(\phi_{2},0)$ with $\phi_{1}>\phi_{2}$.
+This is natural since the neighboring single-$\bm{q}$ phase is the phase IV.
+The sharp increase of the critical field between I2q and IV phases in Fig.
+7(c) is due to the fact that $a_{\rm X}$ is close to zero.
 V Discussions
@@
 (2017).
-The analysis of this study is fully based on this fact.
+The analysis of this study is based on this experimental fact.
 Namely, we have assumed that the leading order parameter is the quadrupole moment at the L points.
-Constructing the symmetry-allowed free energy consisting of them and the related degrees of freedom, we have analyzed the phase diagram for $\bm{H}\parallel[001]$ in addition to [110].
+Constructing the symmetry-allowed free energy consisting of them and the related degrees of freedom, we have analyzed the phase diagram for $\bm{H}\parallel[001]$ in addition to that for $\bm{H}\parallel[110]$.
 Our analysis demonstrates that a $2z^{2}-x^{2}-y^{2}$ configuration with the ordering vector at the L point is stable at high-$T$ and up to intermediate magnetic field strength for $\bm{H}\parallel[001]$.
@@
 We have shown that single- and double-$\bm{q}$ quadrupole orders are realized in our study as a phenomenological model for PrIr2Zn20.
-The orthogonal double-$\bm{q}$ orders are favored by the effective fourth-order terms due to the X point fields, while the quadruple ones are not.
+The orthogonal double-$\bm{q}$ orders are favored by the effective fourth-order terms arising from the renormalization due to the X point fields, while the quadruple ones are not.
 Although we are interested in the orbital orders at the L points in this paper, quadruple-$\bm{q}$ orders at incommensurate ordering vectors along [111] directions has been studied in cubic magnets, showing hedgehog type topological magnetic structure.Okumura et al.
 (2020, 2022) In principle, such exotic phases are also possible for the orbital model discussed in this paper when the exchange coupling favors incommensurate ordering vectors by introducing interaction frustration or anisotropic interaction ignored in this studyHattori et al.
+(2024).
+In this sense, it is interesting to explore possibilities of incommensurate orbital orders as the candidates for the high-$T$ and intermediate field phase instead of the phase I proposed in this paper as studied in magnetic ordersD.
+Bergman and Balents (2007); S.
+Gao and Rüegg (2017); Rosales et al.
 (2022).
@@
 (25).
-Here, we concentrate on the sublattice independent orders $\bm{\phi}(\bm{r})$ with $\Phi^{\alpha}=\Phi$ and $\theta^{\alpha}_{n}=\theta_{n}$ since only such orders appear in this paper.
-Thus, one rewrites Eq.
+Here, we concentrate on the $A$ sublattice moments, since those for the $B$ sublattice are obtained by ${\boldsymbol{\phi}}_{n}\to{\boldsymbol{\phi}}_{n}(-1)^{p_{n}}$.
+Thus, one can rewrite Eq.
 (25) as
-$\bm{\phi}(\bm{r})=\mathbb{E}_{\mathcal{N}}(\bm{r})\cdot\begin{pmatrix}\bm{\phi}_{1}\\ \bm{\phi}_{2}\\ \cdots\end{pmatrix},\ \bm{\phi}_{n}=\frac{\Phi_{n}}{\sqrt{\mathcal{N}}}\begin{pmatrix}\cos\theta_{n}\\ \sin\theta_{n}\end{pmatrix},$ (29)
+$\bm{\phi}^{A}(\bm{r})=\mathbb{E}_{\mathcal{N}}(\bm{r})\cdot\begin{pmatrix}\bm{\phi}_{1}\\ \bm{\phi}_{2}\\ \cdots\end{pmatrix},\ \bm{\phi}_{n}=\frac{\Phi_{n}}{\sqrt{\mathcal{N}}}\begin{pmatrix}\cos\theta_{n}\\ \sin\theta_{n}\end{pmatrix},$ (29)
 where $n=1,2,\cdots,\mathcal{N}$ and $\mathbb{E}_{\mathcal{N}}$ is an ${\mathcal{N}}$ dimensional vector
@@
 For quadruple-$\bm{q}$ orders, the primitive translational vectors are as twice large as those for the original fcc lattice: (1,1,0), (1,0,1), and (0,1,1).
-The unit cell contains eight sites for the quadruple-$\bm{q}$ orders and sixteen when the B-sublattice is included.
+The unit cell contains eight sites for the quadruple-$\bm{q}$ orders and sixteen when the $B$-sublattice sites are included.
 Figure 9: (a) Double-$\bm{q}$ configuration $\mathbb{E}_{2}$ at the $A$ sublattice.
@@
 8, we here examine the real space configuration of the symmetric quadruple-$\bm{q}$ orders discussed in Sec.
-IV.1, although they are not stabilized in the $T$$H$ phase diagram in Fig.
-6.
+IV.1, although they are not stabilized in the $T$$H$ phase diagram in this paper.
 It is evident that the configurations with “collinear” types such as $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,\pi)$ are partially ordered states.
@@
 They are more important than other modes at general wavenumbers since, for example, $\bm{q}_{1}+\bm{q}_{2}+\bm{q}_{\rm X}\equiv\bf{0}$ leads to cubic mode coupling with the two different fields at the L points.
-The free energy for the X points fields are sublattice diagonal and $F_{\rm X}=\sum_{\alpha={\rm A,B}}F_{\rm X}^{\alpha}$, where
+The free energy for the X points fields are sublattice diagonal and $F_{\rm X}=\sum_{\alpha=A,B}F_{\rm X}^{\alpha}$, where
 $F^{\alpha}_{\rm X}=$ $\frac{1}{2}\sum_{\ell={\rm X,Y,Z}}\bm{Q}^{\alpha}_{\ell}\begin{bmatrix}a_{\rm X}-6bu_{0}^{\alpha}&6bv_{0}^{\alpha}\\ 6bv_{0}^{\alpha}&a_{\rm X}+6bu_{0}^{\alpha}\end{bmatrix}\bm{Q}^{\alpha}_{\ell}$
@@
 $a_{\rm X}\equiv T-4J_{2}+6J_{4},$ (33)
-$\begin{split}&C_{{\rm X}1}^{\alpha}=u_{1}^{\alpha}u_{2}^{\alpha}-v_{1}^{\alpha}v_{2}^{\alpha},\quad C_{{\rm X}2}^{\alpha}=u_{3}^{\alpha}u_{4}^{\alpha}-v_{3}^{\alpha}v_{4}^{\alpha},\\ &S_{{\rm X}1}^{\alpha}=u_{1}^{\alpha}v_{2}^{\alpha}+v_{2}^{\alpha}u_{1}^{\alpha},\quad S_{{\rm X}2}^{\alpha}=u_{3}^{\alpha}v_{4}^{\alpha}+v_{3}^{\alpha}u_{4}^{\alpha},\\ &C_{{\rm Y}1}^{\alpha}=u_{1}^{\alpha}u_{4}^{\alpha}-v_{1}^{\alpha}v_{4}^{\alpha},\quad C_{{\rm Y}2}^{\alpha}=u_{2}^{\alpha}u_{3}^{\alpha}-v_{2}^{\alpha}v_{3}^{\alpha},\\ &S_{{\rm Y}1}^{\alpha}=u_{1}^{\alpha}v_{4}^{\alpha}+v_{1}^{\alpha}u_{4}^{\alpha},\quad S_{{\rm Y}2}^{\alpha}=u_{2}^{\alpha}v_{3}^{\alpha}+v_{2}^{\alpha}u_{3}^{\alpha},\\ &C_{{\rm Z}1}^{\alpha}=u_{1}^{\alpha}u_{3}^{\alpha}-v_{1}^{\alpha}v_{3}^{\alpha},\quad C_{{\rm Z}2}^{\alpha}=u_{2}^{\alpha}u_{4}^{\alpha}-v_{2}^{\alpha}v_{4}^{\alpha},\\ &S_{{\rm Z}1}^{\alpha}=u_{1}^{\alpha}v_{3}^{\alpha}+v_{1}^{\alpha}u_{3}^{\alpha},\quad S_{{\rm Z}2}^{\alpha}=u_{2}^{\alpha}v_{4}^{\alpha}+v_{2}^{\alpha}u_{4}^{\alpha}.\end{split}$ (34)
+$\begin{split}&C_{\ell 1}^{\alpha}=u_{m}^{\alpha}u_{n}^{\alpha}-v_{m}^{\alpha}v_{n}^{\alpha},\quad C_{\ell 2}^{\alpha}=u_{m^{\prime}}^{\alpha}u_{n^{\prime}}^{\alpha}-v_{m^{\prime}}^{\alpha}v_{n^{\prime}}^{\alpha},\\ &S_{\ell 1}^{\alpha}=u_{m}^{\alpha}v_{n}^{\alpha}+v_{m}^{\alpha}u_{n}^{\alpha},\quad S_{\ell 2}^{\alpha}=u_{m^{\prime}}^{\alpha}v_{n^{\prime}}^{\alpha}+v_{m^{\prime}}^{\alpha}u_{n^{\prime}}^{\alpha},\end{split}$ (34)
+where $(\ell,m,n)=({\rm X},1,2),\ ({\rm Y},1,4),\ ({\rm Z},1,3)$, and $(\ell,m^{\prime},n^{\prime})=({\rm X},3,4),\ ({\rm Y},2,3),\ ({\rm Z},2,4)$.
 By minimizing $F_{\rm X}$ with respect to the X points fields, we obtain their stationary values as
@@
 and
-$\delta F_{1}^{\alpha}=$ $-\frac{108b^{3}}{a_{\rm X}^{2}}\Bigg{\{}\sum_{m>n}\Big{[}u_{0}^{\alpha}\cos(2\theta^{\alpha}_{mn})+v_{0}^{\alpha}\sin(2\theta^{\alpha}_{mn})\Big{]}(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}$
-$+6\Big{(}u_{0}^{\alpha}\cos\bar{\theta}^{\alpha}+v_{0}^{\alpha}\sin\bar{\theta}^{\alpha}\Big{)}\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4}\Bigg{\}},$ (37)
-where $\theta_{mn}\equiv\theta_{m}+\theta_{n}$ and $\bar{\theta}^{\alpha}\equiv\theta_{1}^{\alpha}+\theta_{2}^{\alpha}+\theta_{3}^{\alpha}+\theta_{4}^{\alpha}$.
+$\delta F_{1}^{\alpha}=$ $-\frac{108b^{3}}{a_{\rm X}^{2}}\Bigg{[}\sum_{m>n}\cos(2\theta^{\alpha}_{mn}-\theta_{0}^{\alpha})(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}\phi_{0}^{\alpha}$
+$+6\cos(\bar{\theta}^{\alpha}-\bar{\theta}_{\alpha})\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4}\phi_{0}^{\alpha}\Bigg{]},$ (37)
+where $\bm{\phi}_{0}^{\alpha}=(u_{0}^{\alpha},v_{0}^{\alpha})=\phi^{\alpha}_{0}(\cos\theta^{\alpha}_{0},\sin\theta^{\alpha}_{0})$ and $\theta_{mn}\equiv\theta_{m}+\theta_{n}$ and $\bar{\theta}^{\alpha}\equiv\theta_{1}^{\alpha}+\theta_{2}^{\alpha}+\theta_{3}^{\alpha}+\theta_{4}^{\alpha}$.
 Equation (36) renormalizes $c^{\prime}$ and $c^{\prime\prime\prime}$ as
@@
 Appendix C Free energy for the orthogonal double-q configuration
-[…]
+We show the free energy for the orthogonal double-$\bm{q}$ orders with $\bm{\phi}_{2}\perp\bm{\phi}_{1}$: $\bm{\phi}_{1}=\phi_{1}(\cos\theta,\sin\theta),\bm{\phi}_{2}=\phi_{2}(\cos(\theta+\tfrac{\pi}{2}),\sin(\theta+\tfrac{\pi}{2}))=\phi_{2}(-\sin\theta,\cos\theta)$, with the $\Gamma$ point mode $\bm{\phi}_{0}=\phi_{0}(\cos\theta_{0},\sin\theta_{0})$.
+Here, we assume $\bm{\phi}_{1}=\bm{\phi}_{1}^{A}=\bm{\phi}_{1}^{B}$ and $\bm{\phi}_{2}=\bm{\phi}_{2}^{A}=-\bm{\phi}_{2}^{B}$ as discussed in the main text.
+The free energy is given as
 $F=$ $\sum_{n=1,2}\Big{[}(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big{]}+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
@@
 $+\left(c+\frac{c^{\prime}}{2}-\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\cos\theta_{0}\right)\phi_{1}^{4}+F_{\Gamma}.$ (41)
-One can also see that the deviation $\delta\phi=\phi_{1}-\phi_{2}\neq 0$ from the phase III2q accompanies changes in the angle as $\theta=\pi/4+\epsilon$.
+Although we will not discuss in detail, we can also see that the deviation $\delta\phi=\phi_{1}-\phi_{2}\neq 0$ from the phase III2q accompanies changes in the angle as $\theta=\pi/4+\epsilon$.
 The quadratic form of $\delta\phi$ and $\epsilon$ describe the stability of the phase III2q and when one of the eigenvalues vanishes a second-order transition takes place toward the phase II2q.
@@
 $+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos\theta_{0}+F_{\Gamma}.$ (42)
-Rotating $\theta$ leads to the transition to the phase II2q. Assuming $|\theta|\ll 1$ and $\theta_{0}=0$ in Eq. (39), we obtain the quadratic part of $F$ in $\theta$ as
-$F=$ $\Big{[}12\lambda_{b}(\phi_{1}^{2}-\phi_{2}^{2})-\lambda_{c}^{(2)}\phi_{0}(\phi_{1}^{2}-2\phi_{2}^{2})-\frac{1728b^{3}}{a_{\rm X}^{2}}\phi^{2}_{1}\phi^{2}_{2}\Big{]}\phi_{0}\theta^{2}.$ (43)
-Thus, a second-order transition to the phase II2q takes place when []delimited-[][$\cdots$][ ⋯ ] in Eq. (43) vanishes.
+Rotating $\theta$ leads to the transition to the phase II2q. This accompanies also rotation in $\theta_{0}$ accordingly.

### v4

@@
-Orbital orders under magnetic fields in cubic PrIr2Zn20
+L-point quadrupole order under magnetic field in cubic PrIr2Zn20
 Abstract
-We study orbital orders in quadrupolar heavy-fermion compound PrIr2Zn20 under magnetic fields on the basis of the Landau theory.
-Assuming $E_{g}$ orbital (electric quadrupolar) orders in the cubic symmetry with the ordering wavenumber at the L points in the face-centered cubic lattice Brillouin zone as observed experimentally, we construct the Landau free energy and analyze it.
-We find that the unidentified high-temperature ordered phase under the magnetic field $\bm{H}\parallel[001]$ reported earlier can be interpreted as a consequence of the internal rotation in the orbital moments of f electrons at the Pr site.
+We study quadrupole orders in heavy-fermion compound PrIr2Zn20 under magnetic fields on the basis of the Landau theory.
+Assuming $E_{g}$ electric quadrupolar orders in the cubic symmetry with the ordering wavenumber at the L points in the face-centered cubic lattice Brillouin zone as observed experimentally, we construct a Landau free energy and analyze the resulting free energy.
+We find that the unidentified high-temperature ordered phase under the magnetic field $\bm{H}\parallel[001]$ reported earlier arises from the rotation of the quadrupole moments of f electrons on the Pr site.
 We also discuss the phase diagram for other magnetic-field directions and possible double-$\bm{q}$ quadrupolar orders in this system.
-pacs:
-Valid PACS appear here
 I Introduction
-Strongly correlated electron systems with orbital moments have attracted significant attention in recent years.
-Tokura and Nagaosa (2000) Various multipolar orders appear in d- and f-electron-based materials and have been intensively studiedMydosh and Oppeneer (2014); Kuramoto et al.
-(2009); Chen et al.
-(2010).
-Among them, electric quadrupole orders are the most typical and extensively investigatedShiina et al.
-(1997); Onimaru et al.
-(2005); Chubukov et al.
-(2016).
-In addition, spin nematic orders have been also studied as well.Andreev and Grishchuk (1984); Läuchli et al.
-(2006); Tsunetsugu and Arikawa (2006); Jiang et al.
-(2023) The nature of orbital degrees of freedom leads to spatially anisotropic interactions, making the symmetry in orbital space generally anisotropic.
+Strongly correlated electron systems with orbital moments have attracted significant attention in recent years [1, 2].
+Various multipolar orders appear in d- and f-electron-based materials and have been intensively studied [3, 4, 5].
+Among them, electric quadrupole orders are the most typical and extensively investigated ones [6, 7, 8].
+In addition, spin nematic orders have also been studied [9, 10, 11, 12].
+The nature of orbital degrees of freedom leads to spatially anisotropic interactions, and thus, the symmetry in orbital space is generally anisotropic.
 Each type of orbital order has unique properties, and understanding these properties remains a key issue in condensed matter theory and experiment.
 […]
-Among the 1-2-20 compounds, PrIr2Zn20 is a prototypical material, exhibiting quadrupole ordering below 1 K and undergoing a phase transition to a superconducting state at 0.1 KOnimaru et al.
-(2011); Ishii et al.
-(2011); Onimaru et al.
-(2016).
-The early neutron scattering experimentIwasa et al.
-(2017) under an applied magnetic field $\bm{H}\parallel[\bar{1}10]$ suggested an ordering vector of $\bm{q}=(\pi,\pi,\pi)$, corresponding to the L point at the first Brillouin zone boundary.
+Among the 1-2-20 compounds, PrIr2Zn20 is a prototypical material, exhibiting quadrupole ordering below 1 K and undergoing a phase transition to a superconducting state at 0.1 K [31, 32, 33].
+The early neutron scattering experiment [34] under an applied magnetic field $\bm{H}\parallel[\bar{1}10]$ suggested an ordering wave number of $\bm{q}=(\pi,\pi,\pi)$, corresponding to the L point at the first Brillouin zone boundary.
 Here, we have set the lattice constant to unity.
-The analysis of induced magnetic moments indicates an orbital order of the $x^{2}-y^{2}$ type for $\bm{H}\parallel[\bar{1}10]$.
-Recent reinvestigations of the temperature–magnetic field phase diagram reveal strong anisotropy and highlight the existence of the high-temperature phaseKittaka et al.
-(2024) reported in earlier studiesOnimaru et al.
-(2011) across a wide range of magnetic field directions from [001] to [111], while it is absent for $\bm{H}\parallel[110]$.
+The analysis of induced magnetic moments indicated a quadrupole order of the $x^{2}$-$y^{2}$ type for $\bm{H}\parallel[\bar{1}10]$.
+Recent reinvestigations of the temperature–magnetic field phase diagram have revealed strong anisotropy and highlight the existence of the high-temperature phase [35] reported in earlier studies [31] across a wide range of magnetic field directions from [001] to [111], while it is absent for $\bm{H}\parallel[110]$.
 In this paper, we analyze a phenomenological Landau model for PrIr2Zn20, incorporating $E_{g}$ quadrupole moments at the L points.
-We elucidate the nature of $E_{g}$ orbital ordering under magnetic fields and its dependence on field direction.
-We examine the order parameter suggested by the neutron scattering experiments for $\bm{H}\parallel[\bar{1}10]$, construct phase diagrams under applied magnetic fields, and compare them with experimental results.
-The ordering vectors at the L points reveal unique multiple-$\bm{q}$ orders, which differs from the previous studies about quadrupole orders at the X point ordering vectors in the face-center cubic lattice.Tsunetsugu et al.
-(2021); Hattori et al.
-(2023)
+We elucidate the nature of $E_{g}$ quadrupole ordering under magnetic fields and its dependence on field directions.
+We examine the order parameter suggested by the neutron scattering experiments for $\bm{H}\parallel[\bar{1}10]$, construct phase diagrams under applied magnetic fields, and compare them with the experimental results.
+The quadrupole orders at the L points reveal unique multiple-$\bm{q}$ orders, which differ from those in the previous studies on multiple-$\bm{q}$ quadrupole orders at the X-point ordering wavenumbers in the face-center-cubic (fcc) lattice [36, 37].
 This paper is organized as follows.
 In Sec.
-II, we construct the Landau free energy for the $E_{g}$ quadrupole moments at wave vectors $\bm{q}=(\pi,\pi,\pi)$ and $(0,0,0)$.
+II, we construct the Landau free energy for the $E_{g}$ quadrupole moments at wavenumbers $\bm{q}$$(\pi,\pi,\pi)$ and $(0,0,0)$.
 In Sec.
-III, we present the results of the Landau analysis for single-$q$ orbital orders, outlining the fundamental properties of the system.
-We then discuss multiple-$q$ orbital orders and the stability of various ordered phases in Sec.
+III, we present the results of the Landau analysis for single-$\bm{q}$ quadrupole orders, outlining the fundamental properties of the system.
+We then discuss multiple-$\bm{q}$ quadrupole orders and the stability of various ordered phases in Sec.
 IV.
@@
 II Model
-We start to introduce the exchange interactions between the $E_{g}$ orbital moments in Sec.
+We begin by introducing the exchange interactions between the $E_{g}$ quadrupole moments in Sec.
 II.1 and the Landau free energy for the order parameters at the L points in Sec.
 II.2.
-Mode coupling terms related to the order parameter at the L points and other degrees of freedom are discussed in Sec.
-II.5
-Figure 1: Fourier transform of the exchange interaction for $J_{1}=1$ and (a) $(J_{2},J_{3},J_{4})=(0.0,0.0,0.0)$, (b) $(0.0,0.0,1.0)$, and (c) $(0.5,0.3,0.4)$.
-(d) The first Brillouin zone of the fcc lattice.
+Mode-coupling terms related to the order parameter at the L points and other degrees of freedom are discussed in Sec.
+II.5.
+Throughout this paper, we do not consider the $T_{xyz}$ octupolar degrees of freedom, since any octupolar orders have not been experimentally reported in PrIr2Zn20 so far.
+Although the excited CEF states possess various multipole moments such as magnetic dipole, octupole, and $T_{2}$-type quadrupole moments, we will ignore them since they are expected to play a minor role in PrIr2Zn20.
+Figure 1: (a) Crystal structure of the diamond structure formed by Pr ions in PrIr2Zn20.
+AA ($B$) sublattice sites are illustrated in white (red) spheres.
+The unit cell is taken as two sites along the [111] direction as indicated by shaded oval in green.
+Exchange interactions $J_{1}\sim J_{4}$ are also indicated.
+(b) Quadrupole moments $u_{\bm{r}}^{\alpha}$ and $v_{\bm{r}}^{\alpha}$ are schematically shown, where the different color represent the sign of moments.
+Figure 2: (a) The first Brillouin zone of the fcc lattice.
 The four L points are indicated by filled red circles and other symmetric points are indicated by filled black circles.
+(b)–(d) Fourier transform of the exchange interactions with $J_{1}=1$.
+(b) $(J_{2},J_{3},J_{4})=(0.0,0.0,0.0)$, (c) $(0.0,0.0,1.0)$, and (d) $(0.5,0.3,0.4)$.
 II.1 Exchange interaction
+Local moments of f electrons on the Pr sites, which form the diamond structure as shown in Fig.
+1(a), tend to order at low temperatures.
+This is usually caused by Ruderman-Kittel-Kasuya-Yosida interaction due to the presence of the itinerant electrons in heavy-fermion compounds in general.
+To discuss their ordering, it is reasonable to integrate the itinerant electrons and obtain the effective exchange model for the localized f electron moments.
+Since the local moments play a dominant role in the symmetry breaking orders in these compounds, we use the localized model as a minimal one to understand the nature of their symmetry breakings.
+Although it is desirable to obtain the information of the exchange interactions from the excitation spectra by inelastic x-ray or neutron scattering experiments, there is no data available at present.
+Thus, we use a phenomenological and minimal model which realizes the L point order in this paper.
 Let us denote $E_{g}$ quadrupole moments at the position $\bm{r}=(x,y,z)$ in the diamond structure as $\bm{\phi}_{\bm{r}}^{\alpha}=(u_{\bm{r}}^{\alpha},v_{\bm{r}}^{\alpha})$, where $\alpha=A$ or $B$ is the sublattice index.
-We take one of the $A$ sublattice site at the origin $\bm{r}=(x,y,z)=(0,0,0)$ and its pair $B$ sublattice site in the unit cell at $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$, where we set the lattice constant $a=1$.
-$\bm{\phi}_{\bm{r}}^{\alpha}$ form a two-dimensional vector field as a classical variable.
-The two components are $u$ $\sim$ $2z^{2}$--$x^{2}$--$y^{2}$ and $v$ $\sim$ $\sqrt{3}$($x^{2}$--$y^{2}$).
+See Fig.
+1(b).
+We take one of the $A$-sublattice site at the origin $\bm{r}=(x,y,z)=(0,0,0)$ and its pair $B$ sublattice site in the unit cell at $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$, where we set the lattice constant $a=1$.
+$\bm{\phi}_{\bm{r}}^{\alpha}$ forms a two-dimensional vector field as a classical variable.
+The two components are $u$ $\sim$ $2z^{2}$-$x^{2}$-$y^{2}$ and $v$ $\sim$ $\sqrt{3}$($x^{2}$-$y^{2}$).
 The Fourier transform in the wavenumber $\bm{k}$} space is defined as
-${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}=\begin{bmatrix}u^{\alpha}_{{\bm{k}}}\\ v^{\alpha}_{{\bm{k}}}\end{bmatrix}=\frac{1}{\sqrt{N}}\sum_{\bm{r}}e^{-i{\bm{k}}\cdot\bm{r}}\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}.$ (1)
+${\bm{\phi}}_{{\bm{k}}}^{\alpha}=\begin{bmatrix}u^{\alpha}_{{\bm{k}}}\\ v^{\alpha}_{{\bm{k}}}\end{bmatrix}=\frac{1}{\sqrt{N}}\sum_{\bm{r}}e^{-i{\bm{k}}\cdot\bm{r}}\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}$.
+(1)
 Here, the summation for $\bm{r}$} runs over the $A$-sublattice sites and $N$ is the number of the unit cell.
 The exchange Hamiltonian is in general written as
-$H=\sum_{{\bm{k}}}[u^{A}_{-{\bm{k}}},v^{A}_{-{\bm{k}}},u^{B}_{-{\bm{k}}},v^{B}_{-{\bm{k}}}]\begin{bmatrix}H^{AA}_{{\bm{k}}}&H^{AB}_{{\bm{k}}}\\ {H^{AB}_{{\bm{k}}}}^{\dagger}&H^{AA}_{{\bm{k}}}\end{bmatrix}\begin{bmatrix}u^{A}_{{\bm{k}}}\\ v^{A}_{{\bm{k}}}\\ u^{B}_{{\bm{k}}}\\ v^{B}_{{\bm{k}}}\end{bmatrix},$ (2)
-where $H^{\alpha\beta}_{{\bm{k}}}$ is a $2\times 2$ matrix.
-We focus on the orders with the ordering wavenumber at the L points: ${\bm{k}}_{1}\equiv(\pi,\pi,\pi)$, ${\bm{k}}_{2}\equiv(-\pi,\pi,\pi)$, ${\bm{k}}_{3}\equiv(-\pi,-\pi,\pi)$, and ${\bm{k}}_{4}\equiv(\pi,-\pi,\pi)$.
-Since the exchange interactions at the L points are isotropic due to the symmetry, we retain such exchange interactions while neglect anisotropic ones.
-Although the anisotropic ones are generally present and can influence the ordering vectors and the excitation energy of quadrupole wave, it has no effect on the transition temperature for the order at the L points.
+${\mathcal{H}}=\sum_{{\bm{k}}}[u^{A}_{-{\bm{k}}},v^{A}_{-{\bm{k}}},u^{B}_{-{\bm{k}}},v^{B}_{-{\bm{k}}}]\begin{bmatrix}{\mathcal{H}}^{AA}_{{\bm{k}}}&{\mathcal{H}}^{AB}_{{\bm{k}}}\\ {{\mathcal{H}}^{AB}_{{\bm{k}}}}^{\dagger}&{\mathcal{H}}^{AA}_{{\bm{k}}}\end{bmatrix}\begin{bmatrix}u^{A}_{{\bm{k}}}\\ v^{A}_{{\bm{k}}}\\ u^{B}_{{\bm{k}}}\\ v^{B}_{{\bm{k}}}\end{bmatrix}$, (2)
+where ${\mathcal{H}}^{\alpha\beta}_{{\bm{k}}}$ is a $2\times 2$ matrix.
+Such model has been analyzed in incommensurate magnetic orders in the diamond structure [38] and the fcc lattice as a part of the diamond structure [39].
+We focus on the orders with the ordering wavenumber at the L points: ${\bm{k}}_{1}\equiv(\pi,\pi,\pi)$, ${\bm{k}}_{2}\equiv(-\pi,\pi,\pi)$, ${\bm{k}}_{3}\equiv(-\pi,-\pi,\pi)$, and ${\bm{k}}_{4}\equiv(\pi,-\pi,\pi)$ as shown in Fig.
+2(a).
+Since the exchange interactions at the L points are isotropic due to the symmetry, we retain such exchange interactions while neglecting anisotropic ones in the following.
+Although the anisotropic ones are generally present and can influence the ordering vectors and the excitation energy of quadrupole wave, they have no effect on the transition temperature for the order at the L points.
 Thus, a model without anisotropic interactions can be a good starting model as a minimal one.
@@
 The $2\times 2$ blocks in Eq. (2) are given by
-$H^{AA}_{{\bm{k}}}$ $=[J_{2}\gamma_{2}({\bm{k}})+J_{4}\gamma_{4}({\bm{k}})]\hat{1},$ (3)
-$H^{AB}_{{\bm{k}}}$ $=\frac{1}{2}[J_{1}\gamma_{1}({\bm{k}})+J_{3}\gamma_{3}({\bm{k}})]\hat{1}.$ (4)
-Here, $\hat{1}$ is a 2$\times$2 identity matrix.
-$J_{i}$’s are isotropic coupling constants.
+${\mathcal{H}}^{AA}_{{\bm{k}}}$ $=[J_{2}\gamma_{2}({\bm{k}})+J_{4}\gamma_{4}({\bm{k}})]\hat{1}$, (3)
+${\mathcal{H}}^{AB}_{{\bm{k}}}$ $=\frac{1}{2}[J_{1}\gamma_{1}({\bm{k}})+J_{3}\gamma_{3}({\bm{k}})]\hat{1}$. (4)
+Here, $\hat{1}$ is a $\times2$ identity matrix.
+$J_{i}$’s are isotropic coupling constants between the $i$th neighbor sites as indicated in Fig.
+1(a).
 The expressions for $\gamma$’s are listed in Table 1.
 Note that $\gamma_{1}({\bm{k}})$ and $\gamma_{3}({\bm{k}})$ are complex.
-From the neutron experimental data, the ordering patterns are consistent with those expected for $J_{1}>0$ Iwasa et al.
-(2017).
-For describing the symmetry breaking at the L point modulation within our model, it is important to check which interactions are important.
+From the neutron experimental data, the ordering patterns are consistent with those expected for $J_{1}>0$ [34].
+For describing the symmetry breaking at the L-point modulation within our model, it is important to identify relevant interactions.
 As will be explained later, four special wavenumbers at the $\Gamma$ point and the X points $\{(2\pi,0,0),(0,2\pi,0),(0,0,2\pi)\}$ are important other than the L point.
 First, among the same sub-lattice interactions, we note that the next-nearest interaction $\gamma_{2}$ vanishes at the four L points, while the fourth neighbor one $\gamma_{4}$ is finite as shown in Table 1.
-This means that positive $J_{4}$ favors the order at the L points while does not those at the $\Gamma$ and X points.
+This means that positive $J_{4}$ favors the order at the L points while does not favor those at the $\Gamma$ and X points.
 For the inter sublattice ones, the nearest-neighbor interaction $\gamma_{1}$ and the third-neighbor one $\gamma_{3}$ are finite, while both vanish at the X points.
@@
 (2), one obtains the eigenvalues $J({\bm{k}})$.
-Figure 1 shows $J({\bm{k}})$ for typical parameter sets with $J_{1}=1$ as a unit of energy.
+Figures 2(b)–2(d) show $J({\bm{k}})$ for typical parameter sets with $J_{1}=1$ as a unit of energy.
 For the nearest-neighbor model with $J_{4}=0$ in Fig.
-1(a), $J({\bm{k}})$ is smallest at the $\Gamma$ point, leading to a Néel type antifero quadrupole order.
+2(b), $J({\bm{k}})$ is smallest at the $\Gamma$ point, leading to a Néel type antiferoic quadrupole order.
 When $J_{4}$ is sufficiently large or $J_{2}$ is positive and not too large, the $J({\bm{k}})$ has minima at the L points as shown in Figs.
-1(b) and 1(c), and these cases are situations we analyze in this paper.
+2(c) and 2(d), and these cases are situations we analyze in this paper.
+For $J_{1}$-$J_{2}$ model, it is known that the frustration leads to various incommensurate orders in the diamond lattice Heisenberg magnet [40].
+Since there is no experimental information about the exchange coupling constants, we try to understand the basic phase diagram for the quadrupole orders in PrIr2Zn20 [31, 32, 35], focusing on the parameter sets that lead to the orders at the L points.
 Table 1: exchange interaction coefficients.
 The values of $\gamma({\bm{k}})$’s at ${\bm{k}}=(0,0,0),(0,0,2\pi)$, ${\bm{k}}_{1}=(\pi,\pi,\pi)$, and the other L points are also shown.
-Here, we take the unit cell convention of $A$: (0,0,0) and $B$: $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$ and thus, the ${\bm{k}}_{1}$ is apparently inequivalent to the remaining ${\bm{k}}_{2,3,4}$.
+Here, we take the unit cell convention of AA: (0,0,0) and $B$: $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$ and thus, the ${\bm{k}}_{1}$ is inequivalent to the remaining ${\bm{k}}_{2,3,4}$.
 $\eta_{{\bm{k}}}\equiv e^{i(k_{x}+k_{y}+k_{z})/4}$, $\zeta^{\pm}_{{\bm{k}}}\equiv\pm 2(\cos\frac{k_{x}}{2}+\cos\frac{k_{y}}{2}+\cos\frac{k_{z}}{2})-3$.
-The mean-field second-order transition temperatures $T_{c}$’s are $T_{c}=2|J_{1}-3J_{3}|+6J_{4}$ for the L points, $T_{c}=4|J_{1}+3J_{3}|-12J_{2}-6J_{4}$ for the $\Gamma$ points, and $T_{c}=4J_{2}-6J_{4}$ for the X points.
-notation expression $\Gamma(0,0,0)$ X(0,0,$2\pi$) L($\pi,\pi,\pi$) L(others)
-$\gamma_{2}({\bm{k}})$ $2\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{z}}{2}\cos\frac{k_{x}}{2}\right)$ $6$ $-2$ 00 00
-$\gamma_{4}({\bm{k}})$ $\cos k_{x}+\cos k_{y}+\cos k_{z}$ $3$ $3$ $-3$ $-3$
-$\gamma_{1}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\right)$ $4$ 00 $-2$ $2$
-$\gamma_{3}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}\zeta^{+}_{{\bm{k}}}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\zeta^{-}_{{\bm{k}}}\right)$ $12$ 00 $6$ $-6$
+The mean-field second-order transition temperatures $T_{c}$’s are $T_{c}=2|J_{1}-3J_{3}|+6J_{4}$ for the L points, $T_{c}=4|J_{1}+3J_{3}|-12J_{2}-6J_{4}$ for the $\Gamma$ point, and $T_{c}=4J_{2}-6J_{4}$ for the X points.
+[…]
+$\gamma_{2}({\bm{k}})$ $2\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{z}}{2}\cos\frac{k_{x}}{2}\right)$ 66 2-2 0 0
+$\gamma_{4}({\bm{k}})$ $\cos k_{x}+\cos k_{y}+\cos k_{z}$ 33 33 3-3 3-3
+$\gamma_{1}({\bm{k}})$ $\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\right) 44$ 0 2-2 22
+$\gamma_{3}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}\zeta^{+}_{{\bm{k}}}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\zeta^{-}_{{\bm{k}}}\right)$ 1212 0 66 6-6
 II.2 Landau Expansion
-We now focus on the order parameter at the four equivalent L points.
+We focus on the order parameter at the four equivalent L points.
 Note that ${\bm{k}}_{1}+{\bm{k}}_{2}+{\bm{k}}_{3}+{\bm{k}}_{4}\equiv{\bf 0}$ and that the fields at the L points are real since ${\bm{k}}_{n}\equiv-{\bm{k}}_{n}$.
-The Landau’s free energy $F$ for the quadrupole moments is written as
-$F$ $=\sum_{{\bm{k}}}[{\boldsymbol{\phi}}^{A}_{-{\bm{k}}},{\boldsymbol{\phi}}^{B}_{-{\bm{k}}}]\left[\frac{T}{2}+\hat{J}({\bm{k}})\right]\begin{bmatrix}{\boldsymbol{\phi}}^{A}_{{\bm{k}}}\\ {\boldsymbol{\phi}}^{B}_{{\bm{k}}}\end{bmatrix}$
+The Landau’s free energy $\mathcal{F}$} for the quadrupole moments is written as
+${\mathcal{F}}$ $=\sum_{{\bm{k}}}[{\bm{\phi}}^{A}_{-{\bm{k}}},{\bm{\phi}}^{B}_{-{\bm{k}}}]\left[\frac{T}{2}+\hat{J}({\bm{k}})\right]\begin{bmatrix}{\bm{\phi}}^{A}_{{\bm{k}}}\\ {\bm{\phi}}^{B}_{{\bm{k}}}\end{bmatrix}$
 $+\sum_{\alpha}\Big[-b\sum_{{\bm{k}},{\bm{k}}^{\prime},{\bm{k}}^{\prime\prime}}u^{\alpha}_{{\bm{k}}}\Big(u^{\alpha}_{{\bm{k}}^{\prime}}u^{\alpha}_{{\bm{k}}^{\prime\prime}}-3v^{\alpha}_{{\bm{k}}^{\prime}}v^{\alpha}_{{\bm{k}}^{\prime\prime}}\Big)\delta_{{\bm{k}}+{\bm{k}}^{\prime}+{\bm{k}}^{\prime\prime},{\bm{G}}}$
-$+\frac{c}{4}\sum_{{\bm{k}},{\bm{k}}^{\prime},{\bm{k}}^{\prime\prime},{\bm{k}}^{\prime\prime\prime}}({\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}}\cdot{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime}})({\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime\prime}})\delta_{{\bm{k}}+{\bm{k}}^{\prime}+{\bm{k}}^{\prime\prime}+{\bm{k}}^{\prime\prime\prime},{\bm{G}}}\Big{]}.$ (5)
+$+\frac{c}{4}\sum_{{\bm{k}},{\bm{k}}^{\prime},{\bm{k}}^{\prime\prime},{\bm{k}}^{\prime\prime\prime}}({\bm{\phi}}^{\alpha}_{{\bm{k}}}\cdot{\bm{\phi}}^{\alpha}_{{\bm{k}}^{\prime}})({\bm{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime}}\cdot{\bm{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime\prime}})\delta_{{\bm{k}}+{\bm{k}}^{\prime}+{\bm{k}}^{\prime\prime}+{\bm{k}}^{\prime\prime\prime},{\bm{G}}}\Big].$ (5)
 Here, $\hat{J}({\bm{k}})$ is the $4\times 4$ matrix in Eq.
-(2) with its eigenvalue $J({\bm{k}})$ and $T$ is the temperature and $\bm{G}$ is a reciprocal lattice vector.
+(2) with its eigenvalue $J({\bm{k}})$, $T$ is the temperature, and $\bm{G}$ is a reciprocal lattice vector.
 We have taken into account local cubic and fourth-order terms only, which arise from local anisotropy such as CEF potential.
+We note that the cubic term $b$ arises from the effects of the excited states and relatively smaller than $c$ as estimated by using the actual CEF scheme [21].
 II.3 L point free energy
-We now concentrate on the L point components ${\boldsymbol{\phi}}^{\alpha}_{n}=(u_{n}^{\alpha},v_{n}^{\alpha})\equiv{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}(n=1,2,3,4)$ with the short handed notation, we find the part of the free energy consisting of only ${\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}$’s is
-$F_{\rm L}$ $=\frac{a_{\rm L}}{2}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}+J_{\rm L}\sum_{n}(-1)^{p_{n}}{\boldsymbol{\phi}}^{A}_{n}\cdot{\boldsymbol{\phi}}^{B}_{n}+\frac{c}{4}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{4}$
-$+\frac{c^{\prime}}{4}\sum_{\alpha}\sum_{m<n}|{\boldsymbol{\phi}}^{\alpha}_{m}|^{2}|{\boldsymbol{\phi}}^{\alpha}_{n}|^{2}+\frac{c^{\prime\prime}}{4}\sum_{\alpha}\sum_{m<n}({\boldsymbol{\phi}}_{m}^{\alpha}\cdot{\boldsymbol{\phi}}_{n}^{\alpha})^{2}$
-$+\frac{c^{\prime\prime\prime}}{4}\sum_{\alpha}\sum_{(m,n,m^{\prime},n^{\prime})}({\boldsymbol{\phi}}^{\alpha}_{m}\cdot{\boldsymbol{\phi}}^{\alpha}_{n})({\boldsymbol{\phi}}^{\alpha}_{m^{\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{n^{\prime}}).$ (6)
-Note that $b$ terms are absent owing to the wavnumber conservation and $(m,n,m^{\prime},n^{\prime})=(1,2,3,4)$, $(1,3,2,4)$, and $(1,4,2,3)$ in the last term in the right-hand side.
-The factor $p_{n}(n=1,2,3,4)$ represents the sign of exchange interaction between the $A$ and $B$ sublattices for four kinds of the L points.
+We concentrate on the L point components ${\bm{\phi}}^{\alpha}_{n}=(u_{n}^{\alpha},v_{n}^{\alpha})\equiv{\bm{\phi}}^{\alpha}_{{\bm{k}}_{n}}(n=1,2,3,4)$ with the short-handed notation, we find the part of the free energy consisting of only ${\bm{\phi}}^{\alpha}_{{\bm{k}}_{n}}$’s is
+${\mathcal{F}}_{\rm L}$ $=\frac{a_{\rm L}}{2}\sum_{\alpha,n}|{\bm{\phi}}_{n}^{\alpha}|^{2}+J_{\rm L}\sum_{n}(-1)^{p_{n}}{\bm{\phi}}^{A}_{n}\cdot{\bm{\phi}}^{B}_{n}+\frac{c}{4}\sum_{\alpha,n}|{\bm{\phi}}_{n}^{\alpha}|^{4}$
+$+\frac{c^{\prime}}{4}\sum_{\alpha}\sum_{m<n}|{\bm{\phi}}^{\alpha}_{m}|^{2}|{\bm{\phi}}^{\alpha}_{n}|^{2}+\frac{c^{\prime\prime}}{4}\sum_{\alpha}\sum_{m<n}({\bm{\phi}}_{m}^{\alpha}\cdot{\bm{\phi}}_{n}^{\alpha})^{2}$
+$+\frac{c^{\prime\prime\prime}}{4}\sum_{\alpha}\sum_{(m,n,m^{\prime},n^{\prime})}({\bm{\phi}}^{\alpha}_{m}\cdot{\bm{\phi}}^{\alpha}_{n})({\bm{\phi}}^{\alpha}_{m^{\prime}}\cdot{\bm{\phi}}^{\alpha}_{n^{\prime}}).$ (6)
+Note that $b$ terms are absent owing to the wavenumber conservation and $(m,n,m^{\prime},n^{\prime})=(1,2,3,4)$, $(1,3,2,4)$, and $(1,4,2,3)$ in the last term in the right-hand side.
+The factor $p_{n}(n=1,2,3,4)$ represents the sign of exchange interaction between the $A$ and $B$ sublattices for four types of L points.
 Since our unit cell consists of the bond parallel to ${\bm{k}}_{1}$, $p_{n}=0$ is distinct from the others $p_{2,3,4}=1$.
@@
 $c^{\prime}=2c,\quad c^{\prime\prime}=4c,\quad c^{\prime\prime\prime}=8c.$ (8)
-Note that ${\bm{k}}_{n}\equiv-{\bm{k}}_{n}$ and thus the ${\boldsymbol{\phi}}_{n}\cdot{\boldsymbol{\phi}}_{n}$ pair in the summation in $\sum_{{\bm{k}}}$ appears once.
-Since each of the fourth order terms are allowed by the cubic symmetry, we denote them as independent parameters.
-For discussing a single-$\bm{q}$ order at the L points, $c^{\prime}$, $c^{\prime\prime}$, and $c^{\prime\prime\prime}$ terms are irrelevant and the fourth-order free energy is simply given by
-$F_{L}^{1\bm{q}(4)}=\frac{c}{4}\sum_{\alpha}|{\boldsymbol{\phi}}^{\alpha}|^{4}.$ (9)
-When multiple-$\bm{q}$ states are considered, the coupling constants $c^{\prime}$, $c^{\prime\prime}$, and $c^{\prime\prime\prime}$ matter as will be discussed in Sec. V.
+Note that ${\bm{k}}_{n}\equiv-{\bm{k}}_{n}$ and thus the ${\bm{\phi}}_{n}\cdot{\bm{\phi}}_{n}$ pair in the summation in $\sum_{{\bm{k}}}$ appears once.
+Since the three fourth-order terms are independently allowed by the cubic symmetry, we denote them as independent parameters.
+For discussing a single-$\bm{q}$ order at the L points, $^{\prime}, c′′c^{\prime\prime}$, and $c^{\prime\prime\prime}$ terms are irrelevant and the fourth-order free energy is simply given by
+${\mathcal{F}}_{L}^{1\bm{q}(4)}=\frac{c}{4}\sum_{\alpha}|{\bm{\phi}}^{\alpha}|^{4}.$ (9)
+When multiple-$\bm{q}$ states are considered, the coupling constants $^{\prime}, c′′c^{\prime\prime}$, and $c^{\prime\prime\prime}$ matter as discussed in Sec. V.
 II.4 $\Gamma$ point free energy
-Under finite magnetic fields, the uniform quadrupole moment ${\boldsymbol{\phi}}_{0}^{\alpha}\equiv{\boldsymbol{\phi}}_{{\bm{k}}=\bf 0}^{\alpha}$ is induced via the effective field $\tilde{\bm{h}}$ in the second-order perturbation of the magnetic field $\bm{H}$.
-The free energy up to the fourth-order terms for the $\Gamma$ point modes are
-$F_{\Gamma}$ $=\sum_{\alpha}\Big{(}-\tilde{\bm{h}}\cdot{\boldsymbol{\phi}}_{0}^{\alpha}+\frac{a_{\Gamma 0}}{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}\Big{)}+J_{\Gamma 0}{\boldsymbol{\phi}}_{0}^{A}\cdot{\boldsymbol{\phi}}_{0}^{B}$
-$+\sum_{\alpha}\Big{\{}-bu_{0}^{\alpha}\Big{[}\big{(}u_{0}^{\alpha}\big{)}^{2}-3\big{(}v_{0}^{\alpha}\big{)}^{2}\Big{]}+\frac{c}{4}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{4}\Big{\}},$ (10)
+Under finite magnetic fields, the uniform quadrupole moment ${\bm{\phi}}_{0}^{\alpha}\equiv{\bm{\phi}}_{{\bm{k}}=\bf 0}^{\alpha}$ is induced via the effective field $\tilde{\bm{h}}$ in the second-order perturbation of the magnetic field $\bm{H}$.
+The free energy up to the fourth-order terms for the $\Gamma$-point modes are
+${\mathcal{F}}_{\Gamma}$ $=\sum_{\alpha}\Big(-\tilde{\bm{h}}\cdot{\bm{\phi}}_{0}^{\alpha}+\frac{a_{\Gamma 0}}{2}|{\bm{\phi}}_{0}^{\alpha}|^{2}\Big)+J_{\Gamma 0}{\bm{\phi}}_{0}^{A}\cdot{\bm{\phi}}_{0}^{B}$
+$+\sum_{\alpha}\Big\{-bu_{0}^{\alpha}\Big[\big(u_{0}^{\alpha}\big)^{2}-3\big(v_{0}^{\alpha}\big)^{2}\Big]+\frac{c}{4}|{\bm{\phi}}_{0}^{\alpha}|^{4}\Big\},$ (10)
 where
@@
 $\tilde{\bm{h}}=\begin{bmatrix}\tilde{h}_{u}\\ \tilde{h}_{v}\end{bmatrix}\equiv g\begin{bmatrix}2H_{z}^{2}-H_{x}^{2}-H_{y}^{2}\\ \sqrt{3}(H_{x}^{2}-H_{y}^{2})\end{bmatrix}$, (12)
-and $g=(g_{J}\mu_{\rm B})^{2}[7/(3E_{4})-1/E_{5}]$ with the excited $\Gamma_{4,5}$ CEF energy $E_{4,5}$.Hattori and Tsunetsugu (2014) $\mu_{\rm B}$ is the Bohr magneton and $g_{J}$ is the Lande’s $g$ factor for the multiplet with the angular momentum $J$; $g_{J}=4/5$ for the PrIr2Zn20 with $J=4$.
+and $g=(g_{J}\mu_{\rm B})^{2}[7/(3E_{4})-1/E_{5}]$ with the excited $\Gamma_{4,5}$ CEF energy $E_{4,5}$ [21].
+$\mu_{\rm B}$ is the Bohr magneton and $g_{J}$ is the Lande’s $g$ factor for the multiplet with the angular momentum $J$; $g_{J}=4/5$ for the PrIr2Zn20 with $J=4$.
 II.5 Mode coupling terms
-We now discuss mode coupling terms consisting of the fields at the L points and those at the $\Gamma$ points. The lowest-order contributions come from the cubic terms as
-$\delta F_{3}=$ $-3\lambda_{b}\sum_{\alpha,n}\Big{\{}u_{0}^{\alpha}\Big{[}\big{(}u_{n}^{\alpha}\big{)}^{2}-\big{(}v_{n}^{\alpha}\big{)}^{2}\Big{]}-v_{0}^{\alpha}(2u_{n}^{\alpha}v_{n}^{\alpha})\Big{\}}$
+We now discuss mode coupling terms consisting of the fields at the L points and those at the $\Gamma$ point. The lowest-order contributions come from the cubic terms as
+$\delta{\mathcal{F}}_{3}=$ $-3\lambda_{b}\sum_{\alpha,n}\Big\{u_{0}^{\alpha}\Big[\big(u_{n}^{\alpha}\big)^{2}-\big(v_{n}^{\alpha}\big)^{2}\Big]-v_{0}^{\alpha}(2u_{n}^{\alpha}v_{n}^{\alpha})\Big\}$
 $=-3\lambda_{b}\sum_{\alpha,n}\phi_{0}^{\alpha}\big(\phi_{n}^{\alpha}\big)^{2}\cos(\theta^{\alpha}_{0}+2\theta^{\alpha}_{n}).$ (13)
 Here, we have parameterized
-${\boldsymbol{\phi}}^{\alpha}_{\bf 0}$ $=\phi^{\alpha}_{0}\begin{bmatrix}\cos\theta^{\alpha}_{0}\\ \sin\theta^{\alpha}_{0}\end{bmatrix},\quad{\boldsymbol{\phi}}^{\alpha}_{n}=\phi^{\alpha}_{n}\begin{bmatrix}\cos\theta^{\alpha}_{n}\\ \sin\theta^{\alpha}_{n}\end{bmatrix},$ (14)
+${\bm{\phi}}^{\alpha}_{\bf 0}$ $=\phi^{\alpha}_{0}\begin{bmatrix}\cos\theta^{\alpha}_{0}\\ \sin\theta^{\alpha}_{0}\end{bmatrix},\quad{\bm{\phi}}^{\alpha}_{n}=\phi^{\alpha}_{n}\begin{bmatrix}\cos\theta^{\alpha}_{n}\\ \sin\theta^{\alpha}_{n}\end{bmatrix}$, (14)
 and the coupling constant $\lambda_{b}$ has been introduced as an independent parameter.
@@
 The next leading order terms are in the fourth order and given by
-$\delta F_{4}=\sum_{\alpha,n}\left[\frac{\lambda^{(1)}_{c}}{2}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}+\lambda_{c}^{(2)}({\boldsymbol{\phi}}_{n}^{\alpha}\cdot{\boldsymbol{\phi}}_{0}^{\alpha})^{2}\right].$ (15)
+$\delta{\mathcal{F}}_{4}=\sum_{\alpha,n}\left[\frac{\lambda^{(1)}_{c}}{2}|{\bm{\phi}}_{n}^{\alpha}|^{2}|{\bm{\phi}}_{0}^{\alpha}|^{2}+\lambda_{c}^{(2)}({\bm{\phi}}_{n}^{\alpha}\cdot{\bm{\phi}}_{0}^{\alpha})^{2}\right].$ (15)
 Again, the coupling constant $\lambda_{c}^{(1)}$ and $\lambda_{c}^{(2)}$ have been introduced as independent parameters.
@@
 (5), $\lambda^{(1)}_{c}=\lambda_{c}^{(2)}=c$.
-In principle, ${\boldsymbol{\phi}}_{0}^{\alpha}$ can couple with ${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}$ in the form of $\sim{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$ and ${\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$.
-In this paper, we retain the order parameter ${\boldsymbol{\phi}}_{n}^{\alpha}$ and the induced moment ${\boldsymbol{\phi}}_{0}^{\alpha}$ under magnetic fields since the other fields are not important for the phase transition with the L point fields.
-Some renormalization effects arising from fields at the X points will be discussed in Sec.
+In principle, ${\bm{\phi}}_{0}^{\alpha}$ can couple with ${\bm{\phi}}_{{\bm{k}}}^{\alpha}$ in the form of $\sim{\bm{\phi}}_{0}^{\alpha}{\bm{\phi}}_{\bm{k}}^{\alpha}{\bm{\phi}}_{-{\bm{k}}}^{\alpha}$ and ${\bm{\phi}}_{0}^{\alpha}{\bm{\phi}}_{0}^{\alpha}{\bm{\phi}}_{\bm{k}}^{\alpha}{\bm{\phi}}_{-{\bm{k}}}^{\alpha}$.
+In this paper, we retain the order parameter ${\bm{\phi}}_{n}^{\alpha}$ and the induced moment ${\bm{\phi}}_{0}^{\alpha}$ under magnetic fields since the other fields are not important for the phase transition with the L point fields.
+Some renormalization effects arising from fields at the X points are discussed in Sec.
 IV and Appendix B.
 III Results: single-q orders
-In this section, we carry out minimization of the free energy $F_{L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}$ [Eqs.
+In this section, we carry out minimization of the free energy ${\mathcal{F}}_{L}+{\mathcal{F}}_{\Gamma}+\delta{\mathcal{F}}_{3}+\delta{\mathcal{F}}_{4}$ [Eqs.
 (6), (10), (13), and (15)], assuming the single-$\bm{q}$ orders.
 The purpose here is to understand how mode coupling effects affect the order parameter configuration.
-The results in this section will become bases for understanding the full phase diagram shown in Sec.
+The results in this section will serve as the basis for understanding the full phase diagram shown in Sec.
 IV, where we will take into account multiple-$\bm{q}$ configurations.
-Since we are interested in the phase transition owing to the order parameters at the L point, we assume the coefficients of the quadratic terms in the free energy at the other wavenumber are always positive.
-We first discuss the candidates of the stable configurations by analyzing the third- and fourth-order terms in $F_{\rm L}$.
+Since we are interested in the phase transition owing to the order parameters at the L point, we assume the coefficients of the quadratic terms in the free energy at ${\bm{k}}\neq{\bm{k}}_{n}$ are always positive.
+We first discuss the candidates of the stable configurations by analyzing the third- and fourth-order terms in ${\mathcal{F}}_{\rm L}$.
 Then, we will show the numerical results of full temperature-field phase diagram.
 III.1 Qualitative analysis
-First, we notice that the inter-sublattice correlations are governed by $J_{\rm L}=-2J_{1}+6J_{3}$ for ${\boldsymbol{\phi}}_{n}^{\alpha}$ in Eq.
+First, we notice that the inter-sublattice correlations are governed by $J_{\rm L}=-2J_{1}+6J_{3}$ for ${\bm{\phi}}_{n}^{\alpha}$ in Eq.
 (6).
-To reproduce the quadrupole configuration propposed by the neutron scattering experimentsIwasa et al.
-(2017), one needs $J_{\rm L}<0$.
+To reproduce the quadrupole configuration propposed by the neutron scattering experiments [34], $J_{\rm L}<0$ is required.
 Thus, it is legitimate to set
-${\boldsymbol{\phi}}_{\rm L}\equiv{\boldsymbol{\phi}}_{1}^{A}={\boldsymbol{\phi}}_{1}^{B}=\begin{bmatrix}u_{\rm L}\\ v_{\rm L}\end{bmatrix}=\phi_{\rm L}\begin{bmatrix}\cos\theta_{\rm L}\\ \sin\theta_{\rm L}\end{bmatrix}.$ (16)
-Here, we have chosen one of the L points at ${\bm{k}}_{1}$ as a representative one since the other three fields at ${\bm{k}}_{n}(n=2,3,4)$ with ${\boldsymbol{\phi}}_{n}^{A}=-{\boldsymbol{\phi}}_{n}^{B}$ are equivalent and this is also the case even when $\bm{H}$ is finite.
+${\bm{\phi}}_{\rm L}\equiv{\bm{\phi}}_{1}^{A}={\bm{\phi}}_{1}^{B}=\begin{bmatrix}u_{\rm L}\\ v_{\rm L}\end{bmatrix}=\phi_{\rm L}\begin{bmatrix}\cos\theta_{\rm L}\\ \sin\theta_{\rm L}\end{bmatrix}$.
+(16)
+Here, we have chosen one of the L points at ${\bm{k}}_{1}$ as a representative one since the other three fields at ${\bm{k}}_{n}(n=2,3,4)$ with ${\bm{\phi}}_{n}^{A}=-{\bm{\phi}}_{n}^{B}$ are equivalent and this is also the case even when $\bm{H}$ is finite.
 See Table 1.
 For the modes at the $\Gamma$ point, which are gapped and just induced by magnetic fields or by the square of the order parameter at the L point, we set
-${\boldsymbol{\phi}}_{0}\equiv{\boldsymbol{\phi}}_{0}^{A}={\boldsymbol{\phi}}_{0}^{B}=\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}=\phi_{0}\begin{bmatrix}\cos\theta_{0}\\ \sin\theta_{0}\end{bmatrix}.$ (17)
-Second, we note that the single-$\bm{q}$ free energy $F_{\rm L}$ is isotropic.
-This means that the transition temperatures for $u_{\rm L}$ and $v_{\rm L}$ are the same at ${\bm{H}}={\bf 0}$ and $T_{c}=6J_{4}+2|J_{1}-3J_{3}|$ This isotropic nature, however, is destroyed by the mode coupling to the fields at the $\Gamma$ points through $\delta F_{3}$ and $\delta F_{4}$.
-When $T\lesssim T_{c}$, the induced ${\boldsymbol{\phi}}_{0}$ is sufficiently small, and thus, the third- and fourth-order terms in $F_{\Gamma}$ can be irrelevant.
-In this situation the relative directions between ${\boldsymbol{\phi}}_{n}$ and ${\boldsymbol{\phi}}_{0}$ is determined by $\delta F_{3}$ and $\delta F_{4}$ as
+${\bm{\phi}}_{0}\equiv{\bm{\phi}}_{0}^{A}={\bm{\phi}}_{0}^{B}=\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}=\phi_{0}\begin{bmatrix}\cos\theta_{0}\\ \sin\theta_{0}\end{bmatrix}$. (17)
+Second, we note that the single-$\bm{q}$ free energy ${\mathcal{F}}_{\rm L}$ is isotropic.
+This means that the transition temperatures for $u_{\rm L}$ and $v_{\rm L}$ are the same at ${\bm{H}}={\bf 0}$ and $T_{c}=6J_{4}+2|J_{1}-3J_{3}|$ .
+This isotropic nature, however, is destroyed by the mode coupling to the fields at the $\Gamma$ point through $\delta{\mathcal{F}}_{3}$ and $\delta{\mathcal{F}}_{4}$.
+When $T\lesssim T_{c}$, the induced ${\bm{\phi}}_{0}$ is sufficiently small, and thus, the third- and fourth-order terms in ${\mathcal{F}}_{\Gamma}$ can be irrelevant.
+In this situation the relative directions between ${\bm{\phi}}_{n}$ and ${\bm{\phi}}_{0}$ is determined by $\delta{\mathcal{F}}_{3}$ and $\delta{\mathcal{F}}_{4}$ as
 $\cos(\theta_{0}+2\theta_{\rm L})=1\ {\rm and}\ \cos(\theta_{0}-\theta_{\rm L})=0.$ (18)
@@
 See Fig.
-2(a).
-We should comment that the three directions of ${\boldsymbol{\phi}}_{0}$ do not coincide with those favored by the $b(>0)$ term in $F_{\Gamma}:-b\phi_{0}^{3}\cos(3\theta_{0})$.
-Since ${\boldsymbol{\phi}}_{0}$ is not the primary order parameter, the third-order term is expected to play a minor role and the configurations (19) are robust at least $\tilde{\bm{h}}={\bf 0}$.
+3(a).
+We should comment that the three directions of ${\bm{\phi}}_{0}$ do not coincide with those favored by the $b(>0)$ term in ${\mathcal{F}}_{\Gamma}:-b\phi_{0}^{3}\cos(3\theta_{0})$.
+Since ${\bm{\phi}}_{0}$ is not the primary order parameter, the third-order term is expected to play a minor role and the configurations (19) are robust at least $\tilde{\bm{h}}={\bf 0}$.
 Figure 3: Schematic single-$\bm{q}$ order parameter configurations for (a) $\bm{H}=0$}, (b) small $\bm{H}$}, (c) small $\bm{H}$} and low $T$, and (d) large $\bm{H}$}.
-Blue (red) arrows indicate ${\boldsymbol{\phi}}_{0}$(${\boldsymbol{\phi}}_{\rm L}$).
+Blue (red) arrows indicate ${\bm{\phi}}_{0}$(${\bm{\phi}}_{\rm L}$).
 Filled red circles represent the directions $\theta_{\rm L}$ given by Eq.
@@
 A configuration for large $\bm{H}$} in the phase II is shown in (e) and that for ${\bm{H}}\parallel[110]$ in (f).
+Figure 4: (a)–(c) TTHH phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, $J_{2}=J_{3}=0.0$, $J_{4}=1.0$, and $c=0.5$.
+(a) $b=0.1$, (b) $b=0.05$, and (c) $b=0.01$.
+Color map represents the L point order parameter angle $\theta_{\rm L}$.
+(d) TTHH phase diagram for $\bm{H}\parallel[110]$ with the same parameter set as in (b).
+The schematic L point order parameter configuration for (e) the phase I and for (f) the phases III and IV.
+For clarity, orbital shapes are drawn only for four sites.
+Red (white) spheres represent Pr sites with $\theta_{\rm L}=0\ (\pi)$ for (e) and $\theta_{\rm L}=\pi/2\ (-\pi/2)$ for (f), respectively.
+On shaded (111) planes, the order parameter $\bm{\phi}^{A}(\bm{r})$ is uniform and $\bm{\phi}^{B}(\bm{r})=\bm{\phi}^{A}(\bm{r})$.
+For the phase II, the L point order parameter is in between those in the phase I and III with $0<|\theta_{\rm L}|<\pi/2$.
 Now let us consider the cases for finite magnetic fields.
-For high symmetric directions $\bm{H}=\tfrac{H}{\sqrt{2}}(\sin\theta,\sin\theta,\sqrt{2}\cos\theta)$ with $20\leq\theta\leq\pi/$ 2, the effective field reads as $\tilde{h}_{u}=gH^{2}(3\cos^{2}\theta-1)$ and $\tilde{h}_{v}=0$.
+For high symmetric directions $\bm{H}=\tfrac{H}{\sqrt{2}}(\sin\Theta,\sin\Theta,\sqrt{2}\cos\Theta)$ with $0\leq\Theta\leq\pi/2$, the effective field reads as $\tilde{h}_{u}=gH^{2}(3\cos^{2}\Theta-1)$ and $\tilde{h}_{v}=0$.
 Note that $\tilde{\bm{h}}={\bf 0}$ at $\bm{H}\parallel[111]$ and $\tilde{h}_{u}>0$ for $\Theta<\tan^{-1}\sqrt{2}\equiv\Theta_{111}$, while $\tilde{h}_{u}<0$ for $\Theta>\Theta_{111}$.
@@
 In contrast, $\Delta\tilde{T}_{c}<0$ for $u_{0}<0$.
-As is clear from the expression of $\Delta\tilde{T}_{c}$, it is expected that there is no additional phase transition within the single-$\bm{q}$ ${\boldsymbol{\phi}}_{n}$ sector for $u_{0}<0$, i.e., $\theta_{0}>\theta_{111}$.
-In the following, we will concentrate on the cases for $u_{0}>0$ ($\theta_{0}<\theta_{111}$).
-For $u_{0}>0$, the linear coupling between ${\boldsymbol{\phi}}_{0}$ and $\tilde{h}_{u}$ can alter the stable configuration from the above six domains (19) since their energy gain is given by the second order in ${\boldsymbol{\phi}}_{0}$.
-Since $\Delta\tilde{T}_{c}>0$, the phase transition from the disordered phase occurs in the $u_{\rm L}$ component: $\theta_{n}=0$ or $\pi$.
-Here, the two choices 00 and $\pi$ represent a simple translation and we will consider $\theta_{\rm L}=0$ below.
-For small $\tilde{h}_{u}$, $\delta F_{3}$ is more important than $\delta F_{4}$.
-This means $\theta_{0}=0$ satisfying $02\theta_{\rm L}+\theta_{0}=$ 0 [Fig.
-2(b)].
-Such a collinear state possesses the free energy $F_{\rm col}$ and the fluctuations around this is given as
-$\delta F_{\rm col}$ $\simeq\lambda_{c}\phi_{0}\phi_{\rm L}^{2}\begin{bmatrix}\theta_{0}&\theta_{\rm L}\end{bmatrix}\hat{A}\begin{bmatrix}\theta_{0}\\ \theta_{\rm L}\end{bmatrix},$ (21)
-$\hat{A}$ $=\begin{bmatrix}(\lambda_{c}\phi_{\rm L}^{2})^{-1}(\tilde{h}_{u}+9b\phi_{0}^{2})+\frac{1}{2}\phi_{0}^{*}-2\phi_{0}&\phi_{0}^{*}+2\phi_{0}\\ \phi_{0}^{*}+2\phi_{0}&2\phi_{0}^{*}-2\phi_{0}\end{bmatrix}.$ (22)
+As is clear from the expression of $\Delta\tilde{T}_{c}$, it is expected that there is no additional phase transition within the single-$\bm{q}$ ${\bm{\phi}}_{n}$ sector for $u_{0}<0$, i.e., $\Theta>\Theta_{111}$.
+In the following, we will concentrate on the cases for $u_{0}>0$ ($\Theta<\Theta_{111}$).
+For $u_{0}>0$, the linear coupling between ${\bm{\phi}}_{0}$ and $\tilde{h}_{u}$ can alter the stable configuration from the above six domains (19) since their energy gain is given by the second order in ${\bm{\phi}}_{0}$.
+Since $\Delta\tilde{T}_{c}>0$, the phase transition from the disordered phase occurs in the $u_{\rm L}$ component: $\theta_{\rm L}=0$ or $\pi$.
+Here, the two choices 0 and $\pi$ represent a simple translation and we will consider $\theta_{\rm L}=0$ below.
+For small $\tilde{h}_{u}$, $\delta{\mathcal{F}}_{3}$ is more important than $\delta{\mathcal{F}}_{4}$.
+This means $\theta_{0}=0$ satisfying $2\theta_{\rm L}+\theta_{0}=0$ [Fig.
+3(b)].
+Such a collinear state possesses the free energy ${\mathcal{F}}_{\rm col}$ and the fluctuations around this is given as
+$\delta{\mathcal{F}}_{\rm col}$ $\simeq\lambda_{c}\phi_{0}\phi_{\rm L}^{2}\begin{bmatrix}\theta_{0}&\theta_{\rm L}\end{bmatrix}\hat{A}\begin{bmatrix}\theta_{0}\\ \theta_{\rm L}\end{bmatrix}$, (21)
+$\hat{A}$ $=\begin{bmatrix}(\lambda_{c}\phi_{\rm L}^{2})^{-1}(\tilde{h}_{u}+9b\phi_{0}^{2})+\frac{1}{2}\phi_{0}^{*}-2\phi_{0}&\phi_{0}^{*}+2\phi_{0}\\ \phi_{0}^{*}+2\phi_{0}&2\phi_{0}^{*}-2\phi_{0}\end{bmatrix}$.
+(22)
 Indeed, the instability to finite $\theta_{{\rm L},0}$ occurs when one of the eigenvalue of the matrix $\hat{A}$ in Eq. (22) becomes zero. This takes place when
@@
 Note that the factor $\phi_{0}^{*}-\phi_{0}$ of the denominator coincides with the factor in $\Delta T_{c}$.
-The configuration with $\theta_{0}\neq 0$ and $\theta_{\rm L}\neq 0$ gradually crossovers toward one for $\bm{H}=0$.
+The configuration with $\theta_{0}\neq 0$ and $\theta_{\rm L}\neq 0$ gradually crosses over to the configuration for $\bm{H}=0$.
 See Fig.
-2(c).
-When $\tilde{\bm{h}}$ is sufficiently large, $u_{0}$ grows and $\delta F_{4}$ plays an important role.
-Thus, ${\boldsymbol{\phi}}_{\rm L}$ is locked at $\theta_{\rm L}=\pm\pi/2$ as shown in Fig.
-2(d).
-This is because $\delta F_{4}>\delta F_{3}$ for large $\tilde{h}_{u}$.
+3(c).
+When $\tilde{\bm{h}}$ is sufficiently large, $u_{0}$ grows and $\delta{\mathcal{F}}_{4}$ plays an important role.
+Thus, ${\bm{\phi}}_{\rm L}$ is locked at $\theta_{\rm L}=\pm\pi/2$ as shown in Fig.
+3(d).
+This is because $\delta{\mathcal{F}}_{4}>\delta{\mathcal{F}}_{3}$ for large $\tilde{h}_{u}$.
 See Eq.
@@
 By the similar analysis to Eqs.
-(21)-(23), the boundary determined by the field below which $|\theta_{\rm L}|<\pi/2$ [Fig.
-2(e)] is given as
+(21)–(23), the boundary determined by the field below which $|\theta_{\rm L}|<\pi/2$ [Fig.
+3(e)] is given as
 $\tilde{h}_{u}^{c2}=27\lambda_{b}\frac{\phi_{0}}{\phi_{0}-\phi_{0}^{*}}\phi_{\rm L}^{2}-9b\phi_{0}^{2}.$ (24)
 For sufficiently large $\lambda_{b}$, the $\theta_{\rm L}=\pm\pi/2$ phase never appears because of $\tilde{T}_{c}^{v}<0$.
-Of course, for extremely large $\tilde{\bm{h}}$, the L point order vanishes trivially.
+Of course, for extremely large $\tilde{\bm{h}}$, the L-point order vanishes trivially.
 Finally, we briefly comment about the cases for $u_{0}<0$ ($\Theta>\Theta_{111}$) under finite magnetic fields.
 As we have noted, there is no phase transition once a quadrupole moment with $\theta_{\rm L}=\pm\pi/2$ is ordered as shown in Fig.
-2(f).
-This is because $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$ can minimize $\delta F_{3}$, $\delta F_{4}$, and $F_{\Gamma}$, simultaneously.
+3(f).
+This is because $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$ can minimize $\delta{\mathcal{F}}_{3}$, $\delta{\mathcal{F}}_{4}$, and ${\mathcal{F}}_{\Gamma}$, simultaneously.
 III.2 Phase diagram
 Now we show the results of numerical minimization of the single-$\bm{q}$ free energy (6), (10), (13), and (15).
-We will show the result for ${\bm{H}}\parallel[001]$, i.e., $\tilde{h}_{u}>0$ and then for ${\bm{H}}\parallel[110]$, $\tilde{h}_{u}<0$.
+We will discuss the result for ${\bm{H}}\parallel[001]$, i.e., $\tilde{h}_{u}>0$ and then for ${\bm{H}}\parallel[110]$, $\tilde{h}_{u}<0$.
 To understand the qualitative nature of the phase diagram, it is sufficient to discuss the cases for $J_{2}=J_{3}=0$.
-Figures 3(a)–3(c) show the $T$-$H$ phase diagram for $\bm{H}\parallel[001]$, where the color represents the angle for the L point moment $\theta_{\rm L}$: (a) $b=0.1$, (b) $b=0.05$ , and (c) $b=0.01$ with $c=0.5$ fixed.
+Figures 4(a)–4(c) show the TTHH phase diagram for $\bm{H}\parallel[001]$, where $H=|\bm{H}|$ and the color represents the angle for the L point moment $\theta_{\rm L}$: (a) $b=0.1$, (b) $b=0.05$, and (c) $b=0.01$ with $c=0.5$ fixed.
 As discussed in the previous section, the phase III is stabilized under high magnetic fields.
@@
 (23); there is no solution of the critical field $\tilde{h}_{u}^{c1}$ when $\phi_{\rm L}^{2}$, which is approximately $H$ independent, is large.
-Figure 3: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, $J_{2}=J_{3}=0.0$, $J_{4}=1.0$, and $c=0.5$.
-(a) $b=0.1$, (b) $b=0.05$, and (c) $b=0.01$.
-Color map represents the L point order parameter angle $\theta_{\rm L}$.
-For $\bm{H}\parallel[110]$, the situation becomes quite simpler.
-Figure 4(a) shows the $T$-$H$ phase diagram for $\bm{H}\parallel[110]$.
+For $\bm{H}\parallel[110]$, the situation becomes much simpler.
+Figure 4(d) shows the TTHH phase diagram for $\bm{H}\parallel[110]$.
 As discussed in the last part in Sec.
 III.1, there is only one stable phase (phase IV), where $\theta_{\rm L}=\pm\pi/2$ and $\theta_{0}=\pi$.
-Figures 4(b) and 4(c) show the real space quadrupole configurations with the ordering vector $\bm{k}_{1}$ for $\bm{H}\parallel[001]$ and $\bm{H}\parallel[110]$, respectively.
-For both cases, the quadrupole orders with $\bm{\phi}_{A}=\bm{\phi}_{B}$ are ferroic on the (111) plane, while antiferroic between the neighboring (111) planes as shown in different colors in Fig.
+For $b=0.1$ and $b=0.01$, the phase diagrams are similar and we do not show them.
+Figures 4(e) and 4(f) show the real space quadrupole configurations with the ordering wave number $\bm{k}_{1}$ for $\bm{H}\parallel[001]$ and $\bm{H}\parallel[110]$, respectively.
+For both cases, the quadrupole orders with $\bm{\phi}_{A}=\bm{\phi}_{B}$ are ferroic on the (111) planes, while antiferroic between the neighboring (111) planes as shown in Fig.
 4.
-When the ordering vector is $\bm{k}_{n}(n=2,3,4)$, the quadrupole configurations are ferroic on the plane perpendicular to ${\bm{k}}_{n}$ with $\bm{\phi}_{A}=-\bm{\phi}_{B}$.
-Figure 4: (a) $T$$H$ phase diagram for $\bm{H}\parallel[110]$, $J_{1}=1.0$, $J_{2}=J_{3}=0$, $J_{4}=1$, $c=0.5$, and $b=0.05$.
-Color map represents the L point order parameter angle $\theta_{\rm L}$.
-Single-$\bm{q}$ order with ${\bm{k}}={\bm{k}}_{1}$ for (b) the phase I and (c) the phase IV.
+Namely the quadrupole moments at the sites indicated by white spheres and those by red ones have opposite sign.
+When the ordering wave number is $\bm{k}_{n}(n=2,3,4)$, the quadrupole configurations are ferroic on the plane perpendicular to ${\bm{k}}_{n}$ with $\bm{\phi}_{A}=-\bm{\phi}_{B}$.
 IV Results: multiple-q orders
@@
 We consider double-$\bm{q}\ (\mathcal{N}=2)$ and quadruple-$\bm{q}\ (\mathcal{N}=4)$ configurations, whose real space moments are given by
-${\boldsymbol{\phi}}^{\alpha}(\bm{r})=\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}=\frac{\Phi_{\alpha}}{\sqrt{\mathcal{N}}}\sum_{n=1}^{\mathcal{N}}e^{i{\bm{k}}_{n}\cdot\bm{r}}\begin{bmatrix}\cos\theta_{n}^{\alpha}\\ \sin\theta_{n}^{\alpha}\\ \end{bmatrix},$ (25)
-where $\bm{r}$ is labeled by the $A$ sublattice position as before and we have assumed that the magnitude for each ${\bm{k}}_{n}$ is the same $\Phi_{\alpha}/\sqrt{\mathcal{N}}(>0)$ for simplicity (symmetric states).
+${\bm{\phi}}^{\alpha}(\bm{r})=\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}=\frac{\Phi_{\alpha}}{\sqrt{\mathcal{N}}}\sum_{n=1}^{\mathcal{N}}e^{i{\bm{k}}_{n}\cdot\bm{r}}\begin{bmatrix}\cos\theta_{n}^{\alpha}\\ \sin\theta_{n}^{\alpha}\\ \end{bmatrix}$, (25)
+where $\bm{r}$ is labeled by the $A$-sublattice position as before and we have assumed that the magnitude for each ${\bm{k}}_{n}$ is the same $\Phi_{\alpha}/\sqrt{\mathcal{N}}~(>0)$ for simplicity (symmetric states).
 For $n=1$ $\bm{\phi}^{A}_{1}=\bm{\phi}^{B}_{1}$, while, for $n=2,3$ and $4$, $\bm{\phi}^{A}_{n}=-\bm{\phi}^{B}_{n}$.
@@
 For double-$\bm{q}$ states, the fourth-order terms are
-$F_{\rm L}^{2\bm{q}(4)}=\sum_{\alpha}\Phi_{\alpha}^{4}\left[\frac{c}{8}+\frac{c^{\prime}}{16}+\frac{c^{\prime\prime}}{16}(c^{\alpha}_{12})^{2}\right].$ (26)
+${\mathcal{F}}_{\rm L}^{2\bm{q}(4)}=\sum_{\alpha}\Phi_{\alpha}^{4}\left[\frac{c}{8}+\frac{c^{\prime}}{16}+\frac{c^{\prime\prime}}{16}(c^{\alpha}_{12})^{2}\right].$ (26)
 Here, $c^{\alpha}_{mn}\equiv\cos(\theta^{\alpha}_{m}-\theta^{\alpha}_{n})$.
@@
 For quadruple-$\bm{q}$ states,
-$F_{\rm L}^{4\bm{q}(4)}=$ $\sum_{\alpha}\Phi_{\alpha}^{4}\Bigg{[}\frac{c}{16}+\frac{3c^{\prime}}{32}+\frac{c^{\prime\prime}}{64}\sum_{m<n}(c^{\alpha}_{mn})^{2}\Bigg{]}.$ (27)
+${\mathcal{F}}_{\rm L}^{4\bm{q}(4)}=$ $\sum_{\alpha}\Phi_{\alpha}^{4}\Bigg[\frac{c}{16}+\frac{3c^{\prime}}{32}+\frac{c^{\prime\prime}}{64}\sum_{m<n}(c^{\alpha}_{mn})^{2}\Bigg].$ (27)
 Note that the $c^{\prime\prime\prime}$ terms cancel when summing up the terms for $\alpha=A$ and $B$, since $\phi^{A}_{n}=(-1)^{p_{n}}\phi^{B}_{n}$.
-The minimum value of $F_{\rm L}^{4\bm{q}(4)}$ for $c^{\prime\prime}>0$ is realized with $\sum_{m<n}(c^{\alpha}_{mn})^{2}=2$, where two orthogonal pairs with the relative angle between the two is arbitrary, e.g., $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})=(\theta^{\alpha}+\pi/2,\theta^{\alpha},\varphi^{\alpha}+\pi/2,\varphi^{\alpha})$.
+The minimum value of ${\mathcal{F}}_{\rm L}^{4\bm{q}(4)}$ for $c^{\prime\prime}>0$ is realized with $\sum_{m<n}(c^{\alpha}_{mn})^{2}=2$, where two orthogonal pairs with the relative angle between the two is arbitrary, e.g., $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})=(\theta^{\alpha}+\pi/2,\theta^{\alpha},\varphi^{\alpha}+\pi/2,\varphi^{\alpha})$.
 For $c^{\prime\prime}<0$, the configurations for the minimum free energy is, e.g., $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})=(\theta,\theta,\theta,\theta)$ with $\sum_{m<n}(c^{\alpha}_{mn})^{2}=6$.
@@
 9 in Appendix A.
-Figure 6 shows the phase diagram at $T=T_{c}$ on the $c^{\prime\prime}$$c^{\prime}$ plane for $c>0$.
+Figure 5 shows the phase diagram at $T=T_{c}$ on the $^{\prime\prime}cc^{\prime}$ plane for $c>0$.
 Note that there is no $c^{\prime\prime\prime}$ dependence.
@@
 The conditions given by Eq.
-(8) just correspond to the point $(c^{\prime}/c,c^{\prime\prime}/c)=(2,4)$, where the single-$\bm{q}$ and double-$\bm{q}$ are degenerate.
-As long as the local fourth-order terms are dominant, quadrupole-$\bm{q}$ orders are unlikely.
+(8) just correspond to the point $(c^{\prime}/c,c^{\prime\prime}/c)=(2,4)$, where the single-$\bm{q}$ and double-$\bm{q}$ orders are degenerate.
+As long as the local fourth-order terms are dominant, quadruple-$\bm{q}$ orders are unlikely.
 Thus, the $c^{\prime}$ term is important when one compares the free energy for the single- and double-$\bm{q}$ orders.
-Figure 5: Phase diagram at $T=T_{c}$ in the $c^{\prime}$$c^{\prime\prime}$ plane for $c>0$.
-Filled circle indicate the bare parameter set $c^{\prime}=2c$ and $c^{\prime\prime}=4c$ given by Eq.
+Figure 5: Phase diagram at $T=T_{c}$ in the $^{\prime}c′′c^{\prime\prime}$ plane for $c>0$.
+A filled circle indicates the bare parameter set $c^{\prime}=2c$ and $c^{\prime\prime}=4c$ given by Eq.
 (8).
@@
 For quadruple-$\bm{q}$ states, their representative configurations for $(\theta^{\alpha}_{1},\theta^{\alpha}_{2},\theta^{\alpha}_{3},\theta^{\alpha}_{4})$ are indicated.
-So far, we have treated the fourth-order coefficients $c$’s as a phenomenological parameters.
+So far, we have treated the fourth-order coefficients $c$’s as phenomenological parameters.
 Here, we discuss the renormalization owing to modes other than $\bm{\phi}_{0}$ and $\bm{\phi}_{n}$.
@@
 (38).
-Furthermore, the fifth-order couplings between the $\Gamma$ and the L points fields appear after integrating out the X point fields as $\sim\phi_{0}\phi_{m}^{2}\phi_{n}^{2}$ and $\sim\phi_{0}\phi_{1}\phi_{2}\phi_{3}\phi_{4}$.
-They play an important role under magnetic fields $\bm{H}$ since $\phi_{0}$ is induced by $\bm{H}$, and thus, these terms renormalize $c$’s with varying their magnitude depending on $\bm{H}$.
+Furthermore, the fifth-order couplings between the $\Gamma$- and the L-points fields appear after integrating out the X-point fields as $\sim\phi_{0}\phi_{m}^{2}\phi_{n}^{2}$ and $\sim\phi_{0}\phi_{1}\phi_{2}\phi_{3}\phi_{4}$.
+They play an important role under magnetic fields $\bm{H}$ since $\phi_{0}$ is induced by $\bm{H}$.
+Thus, these terms renormalize $c$’s with varying their magnitude depending on $\bm{H}$.
 The suppression of $c^{\prime}$ and $c^{\prime\prime\prime}$ due to the renormalization with retaining $c^{\prime\prime}>0$ suggests that the orthogonal double-$\bm{q}$ states are more stable.
@@
 As we will see in Sec.
-IV.2, the fifth-order terms generated by the integration of the X points fields choose some particular $\theta$ among degenerate orthogonal double-$\bm{q}$ configurations.
+IV.2, the fifth-order terms generated by the integration of the X-points fields choose some particular $\theta_{n}$’s among degenerate orthogonal double-$\bm{q}$ configurations.
 IV.2 Phase diagram
-Now, we discuss $T$$H$ phase diagram ($H\equiv|\bm{H}|$) taking into account multiple-$\bm{q}$ orders in addition to the single-$\bm{q}$ ones.
-We numerically minimize the free energy $F_{\rm L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}+\delta F$, where $\delta F$ is due to the integration of the X point fields as discussed in Appendix B and the others are given in Eqs.
+Now, we discuss TTHH phase diagram ($H=|\bm{H}|$) taking into account multiple-$\bm{q}$ orders in addition to the single-$\bm{q}$ ones.
+We numerically minimize the free energy ${\mathcal{F}}_{\rm L}+{\mathcal{F}}_{\Gamma}+\delta{\mathcal{F}}_{3}+\delta{\mathcal{F}}_{4}+\delta{\mathcal{F}}_{\rm{XL}}$, where $\delta{\mathcal{F}}_{\rm{XL}}$ is due to the integration of the X-point fields as discussed in Appendix B and the others are given in Eqs.
 (6), (10), (13), and (15).
-For simplify the analysis, we assume ${\boldsymbol{\phi}}_{n}^{B}={\boldsymbol{\phi}}_{n}^{A}(-1)^{p_{n}}$ valid for $J_{\rm L}<0$ and do not take into account quadruple-$\bm{q}$ orders, since their energy is much higher than single- and double-$\bm{q}$ states for the conventional fourth-order terms [Eq.
+To simplify the analysis, we assume ${\bm{\phi}}_{n}^{B}={\bm{\phi}}_{n}^{A}(-1)^{p_{n}}$ valid for $J_{\rm L}<0$ and do not take into account quadruple-$\bm{q}$ orders, since their energy is much higher than single- and double-$\bm{q}$ states for the conventional fourth-order terms [Eq.
 (8)] as shown in Fig.
-6.
-Figures 6 and 7 show phase diagrams for $\bm{H}\parallel[001]$, i.e., $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\geq 0)$, and those for $\bm{H}\parallel[110]$ with $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\leq 0)$, respectively.
-The representative double-$\bm{q}$ configurations in the quadrupole space and real space are also illustrated in Fig.
-8.
-Figure 6: $T$$H$ phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, and $c=0.5$.
+5.
+Figures 6(a)–6(d) show the phase diagrams for $\bm{H}\parallel[001]$, i.e., $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\geq 0)$, $(J_{2},J_{3},J_{4})=(0.0,0.0,1.0)$ with (a) $b=0.1$ and (b) $b=0.05$.
+For $\bm{H}\parallel[110]$ with $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\leq 0)$, the corresponding phase diagrams are shown in Figs.
+6(c) and 6(d).
+One can notice that there are two double-$\bm{q}$ phases labeled by I2q and II2q, in addition to the single-$\bm{q}$ phases I, II, III, and IV as already discussed above.
+Figure 6: TTHH phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, and $c=0.5$.
 (a) $(b,J_{2},J_{3},J_{4})=(0.1,0.0,0.0,1.0)$, (b) $(0.05,0.0,0.0,1.0)$, and (c) $(0.03,0.5,0.3,0.4)$.
@@
 The exchange coupling constants are the same as used in Fig.
-3: $(J_{2},J_{3},J_{4})=(0,0,1)$.
+4: $(J_{2},J_{3},J_{4})=(0.0,0.0,1.0)$.
 Except for the low-field region, the phase diagram is similar to those shown in Fig.
-3, showing the single-$\bm{q}$ orders I, II, and III.
-An orthogonal double-$\bm{q}$ configuration I2q with $\bm{\phi}_{{\bm{k}}_{1}}=(\phi_{1},0)$ and $\bm{\phi}_{{\bm{k}}_{2}}=(0,\phi_{2})$ is stabilized for finite magnetic fields at high $T$ region below the phase I.
-Here, we take a domain with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$.
+4, showing the single-$\bm{q}$ orders I, II, and III.
+An orthogonal double-$\bm{q}$ configuration I2q with $\bm{\phi}_{{\bm{k}}_{1}}=(\phi_{1},0)$ and $\bm{\phi}_{{\bm{k}}_{2}}=(0,\phi_{2})$ is stabilized for finite magnetic fields, touching its high-$T$ side boundary with the phase I.
 The configuration of I2q state in the quadrupole space is shown in Fig.
-8(a) and its real-space one is illustrated in Fig.
-8(b).
-At ${\bm{H}}=0$, the orthogonal double-$\bm{q}$ states with $\phi_{1}=\phi_{2}$ discussed in Sec.
+7(a) and its real-space one is illustrated in Fig.
+7(b), where a domain with the ordering wave numbers ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$ is shown.
+At $H=0$, the orthogonal double-$\bm{q}$ states with $\phi_{1}=\phi_{2}$ discussed in Sec.
 IV.1 are realized.
@@
 In this sense, $\phi_{1}>\phi_{2}$ is quite natural.
-In the free energy, this is due to the $\delta F_{3}$, which is minimized when the L point field is parallel to $\bm{\phi}_{0}=(u_{0},0)$, while maximized when it is perpendicular.
+In the free energy, this is due to the $\delta{\mathcal{F}}_{3}$, which is minimized when the L point field is parallel to $\bm{\phi}_{0}=(u_{0},0)$, while maximized when it is perpendicular.
 As for the phase boundary between the phase I2q and the phase II, it is of first-order, while the other phase boundaries are second-order.
-At lower $T$, the induced $\Gamma$ point moment $\bm{\phi}_{0}$ favors the configuration III2q shown in Figs.
-8(c) and 8(f), where $\bm{\phi}_{1}=\tfrac{1}{\sqrt{2}}\phi_{1}(1,1)$ and $\bm{\phi}_{2}=\tfrac{1}{\sqrt{2}}\phi_{2}(-1,1)$ with $\phi_{1}=\phi_{2}$.
+[…]
+There is another double-$\bm{q}$ order labeled by II2q in Figs.
+6(a) and 6(b).
+The quadrupole configurations in the phase II2q are “rotated” from those in the phase I2q with keeping $\bm{\phi}_{1}\perp\bm{\phi}_{2}$ as shown in Fig.
+7(b).
+For smaller $b=0.05$ [Fig.
+5(b)], the phase I, II, and the double-$\bm{q}$ orders shrink, while the phase III expands.
+The overall phase diagram, however, is unchanged.
+For $\bm{H}\parallel[110]$, the phase I2q appears at low-field region as shown in Figs.
+6(c) and 6(d), while the phase II2q does not appear.
+In contrast to the phase I2q for $\bm{H}\parallel[001]$, the quadrupole configuration is $\bm{\phi}_{1}=(0,\phi_{1})$ and $\bm{\phi}_{2}=(\phi_{2},0)$ with $\phi_{1}>\phi_{2}$.
+This is natural since the neighboring single-$\bm{q}$ phase is the phase IV.
+The fact that $\phi_{1}$ is larger than $\phi_{2}$ explains why there is no II2q for $\bm{H}\parallel[110]$.
+The coupling $\lambda_{c}^{(2)}$ in Eq.
+(15) favors $\bm{\phi}_{n}\perp\bm{\phi}_{0}$.
+This suggests that the configuration in the phase I2q for $\bm{H}\parallel[110]$ is stable since the larger $\bm{\phi}_{1}$ is perpendicular to $\bm{\phi}_{0}$, while parallel in the phase I2q for $\bm{H}\parallel[110]$.
+Figure 8: TTHH phase diagram for $J_{1}=1.0$, $c=0.5$ and $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ for (a) $\bm{H}\parallel[001]$ and (b) $\bm{H}\parallel[110]$.
+Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, similarly to Fig.
+6.
+(c) Schematic orthogonal double-$\bm{q}$ order parameter configurations in the phase III2q for $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$ with $\phi_{1}=\phi_{2}$ (a) in the uuvv space and (b) in the real space.
+Colored bonds represent the uniform alignment of $\bm{\phi}^{A}(\bm{r})$ similarly to those in Figs.
+7(b) and 7(d).
+As an example of more general cases, we show the phase diagram for $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ in Fig.
+8.
+At low temperatures, the induced $\Gamma$-point moment $\bm{\phi}_{0}$ favors the configuration III2q shown in Figs.
+8(c) and 8(d), where $\bm{\phi}_{1}=\tfrac{1}{\sqrt{2}}\phi_{1}(1,1)$ and $\bm{\phi}_{2}=\tfrac{1}{\sqrt{2}}\phi_{2}(-1,1)$ with $\phi_{1}=\phi_{2}$.
 This configuration is stabilized by the fifth-order coupling discussed in the last part of Sec.
@@
 (39).
-In Fig.
-6(a), an intermediate phase II2q toward the phase III2q appears at low temperatures.
-The quadrupole configurations in the phase II2q are “rotated” from those in the phase I2q with keeping $\bm{\phi}_{1}\perp\bm{\phi}_{2}$ as shown in Fig.
-8(b).
-For smaller $b=0.05$ [Fig.
-6(b)], the phase I, II, and the double-$\bm{q}$ orders shrink, while the phase III expands.
-The overall phase diagram, however, is unchanged.
-Figure 7: $T$$H$ phase diagram for $\bm{H}\parallel[110]$, $J_{1}=1.0$, and $c=0.5$.
-(a) $(b,J_{2},J_{3},J_{4})=(0.1,0.0,0.0,1.0)$, (b) $(0.05,0.0,0.0,1.0)$, and (c) $(0.03,0.5,0.3,0.4)$, which correspond to those in Fig.
-7(a), 7(b), and 7(c), respectively.
-Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, similarly to Fig.
-6
-Figure 8: Schematic orthogonal double-$\bm{q}$ order parameter configurations for (a) I2q: $(\theta_{1},\theta_{2})=(0,\pi/2)$ with $\phi_{1}>\phi_{2}$, (b) II2q: $(\theta_{1},\theta_{2})=(\theta,\theta+\pi/2)$ with $\phi_{1}>\phi_{2}$, and (c) III2q: $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$ and $\phi_{1}=\phi_{2}$.
-Arrows indicate the order parameters $\bm{\phi}_{1}$ (L1), $\bm{\phi}_{2}$ (L2), and $\bm{\phi}_{0}$ $(\Gamma)$.
-Four small circles represent the quadrupole moments at the $A$ sublattice real-space positions $\bm{r}=(0,0,0),(1,0,0),(\tfrac{1}{2},\frac{1}{2},0)$, and $(\tfrac{1}{2},\frac{1}{2},1)$.
-Real space L point quadrupole configurations for (d) $(\theta_{1},\theta_{2})=(0,\pi/2)$ and $\phi_{1}=\sqrt{3}\phi_{2}$ in the phase I2q, (e) $(\theta_{1},\theta_{2})=(\pi/24,13\pi/24)$ and $\phi_{1}=1.2\phi_{2}$ in the phase II2q, and (f) $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$, and $\phi_{1}=\phi_{2}$ in the phase III2q.
-Note that only L point quadrupole components without those at the $\Gamma$ point are shown.
-For guide to eyes, some of the nearest-neighbor bonds are connected by white bonds.
-Circles are the $A$ sublattice atoms and their colors represent the sublattice (red, green, blue, white).
-For clarity, the $B$ sublattice atoms are not drawn.
-Colored bonds running [01¯$1]$ 1 ] directions connects the $A$ sublattice sites with the same quadrupole moments.
-As an example of more general cases, we show the phase diagram for $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ in Fig.
-6(c).
-For this parameter set, the second-order coefficient $a_{\rm X}$ of the X point fields almost vanish at $T=0$ as shown in Fig.
-1(c), the renormalization effects due to the X point fields grows rapidly at low temperatures.
+For this parameter set, the second-order coefficient $a_{\rm X}$ of the X-point fields almost vanish at $T=0$ as shown in Fig.
+2(d), the renormalization effects due to the X point fields grows rapidly at low temperatures.
 This causes the strong stabilization of the phase III2q in Fig.
-6(c), showing reentrant behavior as a function of $H$.
-Note, however, the Landau theory is not applicable to such low temperature region far away from $T\sim T_{c}$.
-There is a thin region of the phase II2q.
+8(a) for $\bm{H}\parallel[001]$, showing reentrant behavior as a function of $H$.
+For $\bm{H}\parallel[110]$, the sharp increase of the critical field between I2q and IV phases in Fig.
+8(b) is also due to the fact that $a_{\rm X}$ is close to zero.
+Note that the Landau theory is not applicable to such low temperature region far away from $T\sim T_{c}$.
+There is a thin region of the phase II2q for $\bm{H}\parallel[001]$.
 The phase boundary between the phase II2q and II is of the second order.
@@
 We note that the boundaries indicated by the dashed lines are not symmetry breaking phase boundary since the symmetry of the three double-$\bm{q}$ phases are the same as is evident from the real space configurations in Figs.
-8(d)–8(f).
+7(b), 7(d), and 8(d).
 See also the discussion in Appendix C.
-For $\bm{H}\parallel[110]$, a phase I2q appears at low-field region as shown in Fig.
-7.
-In contrast to the phase I2q for $\bm{H}\parallel[001]$, the quadrupole configuration is $\bm{\phi}_{1}=(0,\phi_{1})$ and $\bm{\phi}_{2}=(\phi_{2},0)$ with $\phi_{1}>\phi_{2}$.
-This is natural since the neighboring single-$\bm{q}$ phase is the phase IV.
-The sharp increase of the critical field between I2q and IV phases in Fig.
-7(c) is due to the fact that $a_{\rm X}$ is close to zero.
 V Discussions
+So far, we have shown how the L point quadrupole orders evolve under magnetic fields by using the Landau analysis.
+In this section, we will compare these results with the experimental data of PrIr2Zn20 and the related compounds.
+We also discuss the similarity between other quadrupole orders in Pr-based 1-2-20 compounds such as the Néel type antiferroic orders and those at the X points from theoretical point of view.
+In the last part of this section, we will comment on the future perspectives toward full understanding of the quadrupole orders in PrIr2Zn20.
 V.1 Comparison with the experimental data of PrIr2Zn20
-PrIr2Zn20 shows several phase transitions under magnetic field at low temperatures as reported in the early study.Onimaru et al.
-(2011) However, the order parameters have not been identified except for the $x^{2}-y^{2}$ type quadrupole order with the ordering vector at the L point for $\bm{H}\parallel[110]$ by the neutron scattering experimentIwasa et al.
-(2017).
+PrIr2Zn20 shows several phase transitions under magnetic field at low temperatures as reported in the early study [31].
+However, the order parameters have not been identified except for the $x^{2}$-$y^{2}$ type quadrupole order with the ordering wave number at the L point for $\bm{H}\parallel[110]$ by the neutron scattering experiment [34].
 The analysis of this study is based on this experimental fact.
@@
 Constructing the symmetry-allowed free energy consisting of them and the related degrees of freedom, we have analyzed the phase diagram for $\bm{H}\parallel[001]$ in addition to that for $\bm{H}\parallel[110]$.
-Our analysis demonstrates that a $2z^{2}-x^{2}-y^{2}$ configuration with the ordering vector at the L point is stable at high-$T$ and up to intermediate magnetic field strength for $\bm{H}\parallel[001]$.
-This suggests that the $2z^{2}-x^{2}-y^{2}$ order can be a promising candidate for the high-$T$ phase found for $\bm{H}\parallel[001]$ in the experimental worksOnimaru et al.
-(2011); Kittaka et al.
-(2024).
+Our analysis demonstrates that a $2z^{2}$-$x^{2}$-$y^{2}$ configuration with the ordering wave number at the L point is stable at high-$T$ and up to intermediate magnetic field strength for $\bm{H}\parallel[001]$.
+This suggests that the $2z^{2}$-$x^{2}$-$y^{2}$ order can be a promising candidate for the high-$T$ phase found for $\bm{H}\parallel[001]$ in the experimental works [31, 35].
 As for the order parameter at zero field, the single-$\bm{q}$ configuration is not stable while the double-$\bm{q}$ I2q is realized with $\phi_{1}=\phi_{2}$.
@@
 Thus, as increasing $\bm{H}\parallel[001]$, the double-$\bm{q}$, the single-$\bm{q}$ phase II with both $u_{\rm L}$ and $v_{\rm L}$ finite, and then the phase III with only $v_{\rm L}$ are realized.
-The trace of the phase boundaries between them have not been clearly observed in the experiments.
-However, there are several anomalies in the ultrasonic experimentsIshii et al.
-(2011).
+The trace of the phase boundaries between them has not been clearly observed in the experiments.
+However, there are several anomalies in the ultrasonic experiments [32].
 We hope the relation between our phase diagram and these anomalies can be clarified in future experimental studies.
 In addition to this, there is ambiguity about the stability of the phase III, whether the phase III extends toward the higher field as shown in Fig.
-3(c) or occupies only small area [Fig.
-3(a)].
+4(c) or occupies only small area [Fig.
+4(a)].
 This can be controlled by the cubic coupling $b$.
-For $\bm{H}\parallel[110]$, there is no high-$T$ phase in contrast to the case for $\bm{H}\parallel[001]$ in the experiments.Onimaru et al.
-(2011); Kittaka et al.
-(2024) This point is consistent with our results.
-Other anomalies observed in the ultrasound experimentsIshii et al.
-(2011) similar to those for $\bm{H}\parallel[001]$ are also interesting to examine in more detail in the future works.
-Since the anomalies are tiny and thus the comparison between the experiments and the present theory is not straightforward, while there is a phase boundary between the phase I2q and IV in Fig.
-7.
-We also note that the phase diagram as a function of magnetic field from $\bm{H}\parallel[001]$ to $\bm{H}\parallel[110]$ is described in a function of angle $\Theta$, where $\bm{H}=\tfrac{1}{2}H(\sin\Theta,\sin\Theta,2\cos\Theta)$.
+For $\bm{H}\parallel[110]$, there is no high-$T$ phase in contrast to the case for $\bm{H}\parallel[001]$ in the experiments [31, 35].
+This point is consistent with our results.
+Other anomalies observed in the ultrasound experiments [32] similar to those for $\bm{H}\parallel[001]$ are also interesting to examine in more detail in the future works.
+Since the anomalies are tiny and thus the comparison between the experiments and the present theory is not straightforward, while there is a phase boundary between the phase I2q and IV in Figs.
+6(c), 6(d), and 8(b), irrespectively to the detail of the parameters.
+We also note that the phase diagram as a function of magnetic field from $\bm{H}\parallel[001]$ to $\bm{H}\parallel[110]$ is described in a function of angle $\Theta$, where $\bm{H}=\tfrac{1}{\sqrt{2}}H(\sin\Theta,\sin\Theta,\sqrt{2}\cos\Theta)$.
 In Eq.
@@
 $H_{c}(\Theta)=\frac{\sqrt{2}H_{c}(0)}{\sqrt{3\cos^{2}\Theta-1}}.$ (28)
-Such behavior can be one of strong evidences for quadrupole orders. We hope the future experimental studies on the phase diagram at finite fields clarify this point.
-V.2 Future perspectives
+Such behavior can be one of strong evidences for quadrupole orders and indeed have been observed recently [41].
+V.2 Comparison with quadrupole orders at the $\Gamma$ point and the X points
+In the previous studies, quadrupole orders in Pr-based 1-2-20 compounds have been theoretically analyzed in several types of ordering wavenumbers.
+In the early stage, simple Néel type quadrupole orders [21, 23] are analyzed and TTHH phase diagrams show similar order parameter switching as discussed in the analyses of single-$\bm{q}$ sates shown in Fig.
+4.
+This is because the form of the Landau’s free energy for both cases are almost the same when the multiple-$\bm{q}$ configurations are neglected.
+For PrTi2Al20, the experiments clarify the order is ferroic [20, 24, 25] and its unusual behavior in the quadrupole switching under magnetic fields is analyzed by field-induced interactions unique to the quadrupole system [24].
+In this study such induced interactions are not taken into account.
+We do not consider they are important in PrIr2Zn20 since they require specific electronic properties such as existence of small pocket Fermi surfaces.
+For PrV2Al20, there are two transitions at zero magnetic field.
+The high-temperature transition is considered to be a quadrupole order, while ferroic octupole moments emerge below the lower transition temperature [42, 27].
+The Landau analysis for a model including the octupole moment in addition to the quadrupole moments [28] and Monte Carlo simulations with extended interactions have been done [29, 26].
+These theories explain the finite octupole moment as a consequence of symmetry breaking of ferroic octupole order.
+The lack of microscopic observation of both quadrupole and octupole moments (even their ordering wavenumbers) prevents theoretical investigation from further analyses.
+As an alternative approach, it has been pointed out that if the high-temperature transition is that into triple-$\bm{q}$ quadrupole order at the X points suggested by the band structure calculation [43], the lower temperature transition is expected to be triple-$\bm{q}$ octupole order at the same wavenumbers with induced ferroic octupole moments [30].
+The physics of the L point quadrupole order is distinct from the ferroic and the X-point quadrupole moments in particular in the absence of cubic invariant consisting solely of the order parameters since the sum of three ordering wavenumbers lead to ${\bm{k}}_{1}+{\bm{k}}_{2}+{\bm{k}}_{3}\neq{\bm{0}}$.
+Although the Néel type order parameters do not possess such cubic invariant, they have trivially no multiple-$\bm{q}$ orders.
+In this sense, the present results including double-$\bm{q}$ quadrupole orders and the phase transitions from/to the single-$\bm{q}$ orders are unique properties for the L point quadrupole systems.
+V.3 Future perspectives
 We have shown that single- and double-$\bm{q}$ quadrupole orders are realized in our study as a phenomenological model for PrIr2Zn20.
-The orthogonal double-$\bm{q}$ orders are favored by the effective fourth-order terms arising from the renormalization due to the X point fields, while the quadruple ones are not.
-Although we are interested in the orbital orders at the L points in this paper, quadruple-$\bm{q}$ orders at incommensurate ordering vectors along [111] directions has been studied in cubic magnets, showing hedgehog type topological magnetic structure.Okumura et al.
-(2020, 2022) In principle, such exotic phases are also possible for the orbital model discussed in this paper when the exchange coupling favors incommensurate ordering vectors by introducing interaction frustration or anisotropic interaction ignored in this studyHattori et al.
-(2024).
-In this sense, it is interesting to explore possibilities of incommensurate orbital orders as the candidates for the high-$T$ and intermediate field phase instead of the phase I proposed in this paper as studied in magnetic ordersD.
-Bergman and Balents (2007); S.
-Gao and Rüegg (2017); Rosales et al.
-(2022).
+The orthogonal double-$\bm{q}$ orders are favored by the effective fourth-order terms arising from the renormalization due to the X-point fields, while the quadruple ones are not.
+Although we are interested in the orders at the L points in this paper, quadruple-$\bm{q}$ orders at incommensurate ordering wave numbers along [111] directions has been studied in cubic magnets, showing hedgehog type topological magnetic structure [44, 45].
+In principle, such exotic phases are also possible for the orbital model discussed in this paper when the exchange coupling favors incommensurate ordering wave numbers by introducing interaction frustration or anisotropic interaction ignored in this study [46].
+In this sense, it is interesting to explore possibilities of incommensurate orders as the candidates for the high-$T$ and intermediate field phase instead of the phase I proposed in this paper as studied in magnetic orders [40, 47, 48].
 For quantitative comparison between the theoretical and experimental results, it is also needed to incorporate the actual crystal electric field states in the real materials.
 They vary under magnetic fields and thus the parameters in the free energy can change.
-For theoretical or computational aspects, it is also interesting to take into account the fluctuation effects by using e.g., classical Monte Carlo simulation employed in the previous studies on Pr-based materials.Hattori and Tsunetsugu (2016); Freyer et al.
-(2018, 2020) The thermal fluctuation may alter the mean-field phase diagram and to check their stability is important issue as one of our future problems.
+Recent ultrasonic experiments [49] detected rotational effects and its microscopic analysis is also interesting.
+For theoretical or computational aspects, it is also interesting to take into account the fluctuation effects by using e.g., classical Monte Carlo simulation employed in the previous studies on Pr-based materials [23, 29, 26].
+The thermal fluctuation may alter the mean-field phase diagram and to check their stability is important issue as one of our future problems [50].
+For example, our analysis of mode-mode coupling is restricted to those with the X-point modes via the cubic terms.
+This may cause overestimation of double-$\bm{q}$ orders.
+In this sense, the numerical simulation can clarify the effects of these fluctuations.
+The influence of the non-Fermi liquid properties [20, 33, 51, 18] on the symmetry breaking, which is completely ignored in this study, is also nontrivial and important subject.
 VI Summary
@@
 We also propose that the unidentified high-temperature phase for $\bm{H}\parallel[001]$ in the experiments is a collinear $2z^{2}$-$x^{2}$-$y^{2}$ order and the orthogonal double-$\bm{q}$ configurations are stable at low fields.
-These results offers further experimental examination of the order parameters in PrIr2Zn20 and related compounds.
+These results offer further experimental examination of the order parameters in PrIr2Zn20 and related compounds.
 Acknowledgment
-The authors thank T. Onimaru T. Ishitobi, and S. Kittaka for fruitful discussion. This work was supported by JSPS KAKENHI (Grant Nos. JP23K20824 and JP23H04869).
+The authors thank T. Onimaru, T. Ishitobi, and S. Kittaka for fruitful discussion. This work was supported by JSPS KAKENHI (Grant No. JP23K20824, No. JP23H04866, and No. JP23H04869).
+Figure 9: (a) Double-$\bm{q}$ configuration $\mathbb{E}_{2}$.
+Colored thick tubes connect the same $\mathbb{E}_{2}$’s.
+Thin bonds indicate the nearest-neighbor bonds of diamond structure.
+Arrows indicate the direction of ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$.
+(b) Quadruple-$\bm{q}$ configuration $\mathbb{E}_{4}$, where only the $A$-sublattice sites are shown by spheres.
+(c) Symmetric quadruple-$\bm{q}$ configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and (d) for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$.
+For (d), there are four kinds of ferroic chains.
+Two of them are similar to the double-$\bm{q}$ configuration for $x=n$ ($n:$ integers) planes with the $2z^{2}$-$x^{2}$-$y^{2}$ orbital, while the other two chains for $x=n+\tfrac{1}{2}$ plane with the $x^{2}$-$y^{2}$ orbital run along ${\bm{k}}_{1}+{\bm{k}}_{2}$ direction.
 Appendix A Multiple-q configuration
@@
 (25).
-Here, we concentrate on the $A$ sublattice moments, since those for the $B$ sublattice are obtained by ${\boldsymbol{\phi}}_{n}\to{\boldsymbol{\phi}}_{n}(-1)^{p_{n}}$.
+Here, we concentrate on the $A$-sublattice moments, since those for the $B$ sublattice are obtained by ${\bm{\phi}}_{n}\to{\bm{\phi}}_{n}(-1)^{p_{n}}$.
 Thus, one can rewrite Eq.
@@
 Here, we have assumed, without loss of generality, double-$\bm{q}$ orders with ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$.
-Since $\bm{r}$ is the position in the fcc lattice, there are four choices for $\mathbb{E}_{2}(\bm{r}):(\pm 1,\pm 1)$, and eight choices for $\mathbb{E}_{4}(\bm{r}):\pm(1,1,1,1),\ \pm(1,1,-1,-1),\ \pm(1,-1,1,-1)$, and ±$\pm(-1,1,1,-1)$.
+Since $\bm{r}$ is the position in the fcc lattice, there are four choices for $\mathbb{E}_{2}(\bm{r}):(\pm 1,\pm 1)$, and eight choices for $\mathbb{E}_{4}(\bm{r}):\pm(1,1,1,1),\ \pm(1,1,-1,-1),\ \pm(1,-1,1,-1)$, and $\pm(-1,1,1,-1)$.
 In Fig.
 9, $\mathbb{E}_{\mathcal{N}}(\bm{r})$’s are illustrated and the colors distinguish the vectors $\mathbb{E}_{2,4}$.
-Note that only the $A$-sublattice positions in the diamond structure are shown.
-The unit cell for the double-$\bm{q}$ orders contains four $A$ sublattice sites, each of which forms ferroic chain along the direction parallel to ${\bm{k}}_{3}+{\bm{k}}_{4}\parallel(0,-1,1)$ as shown in Fig.
+Note that only the $A$-sublattice positions in the diamond structure are shown in Fig.
+9(b).
+The unit cell for the double-$\bm{q}$ orders contains four $A$-sublattice sites, each of which forms ferroic chain along the direction parallel to ${\bm{k}}_{3}+{\bm{k}}_{4}\parallel(0,-1,1)$ as shown in Fig.
 9(a).
 The primitive translation vector along this direction is unchanged, while those perpendicular to it are modified to $\frac{1}{2}(2,1,1)$ and $\frac{1}{2}(2,-1,-1)$.
-For quadruple-$\bm{q}$ orders, the primitive translational vectors are as twice large as those for the original fcc lattice: (1,1,0), (1,0,1), and (0,1,1).
+For quadruple-$\bm{q}$ orders, the primitive translation vectors are as twice large as those for the original fcc lattice: (1,1,0), (1,0,1), and (0,1,1).
 The unit cell contains eight sites for the quadruple-$\bm{q}$ orders and sixteen when the $B$-sublattice sites are included.
-Figure 9: (a) Double-$\bm{q}$ configuration $\mathbb{E}_{2}$ at the $A$ sublattice.
-Tubes connect the same $\mathbb{E}_{2}$’s.
-Thin lines indicates the NN bonds of diamond structure.
-Arrows indicates the direction of ${\bm{k}}_{1}$ and $-{\bm{k}}_{2}$.
-(b) Quadruple-$\bm{q}$ configuration $\mathbb{E}_{4}$.
-(c) Symmetric quadruple-$\bm{q}$ configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and (d) for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$.
-For (d), there are four kinds of ferroic chains.
-Two of them are similar to the double-$\bm{q}$ configuration for $x=2n$($n:$ integers) planes with the $2z^{2}-x^{2}-y^{2}$ orbital, while the other two chains for $x=2n+1$ plane with the $x^{2}-y^{2}$ orbital run along ${\bm{k}}_{1}+{\bm{k}}_{2}$ direction.
 Since several double-$\bm{q}$ configurations are shown in Fig.
-8, we here examine the real space configuration of the symmetric quadruple-$\bm{q}$ orders discussed in Sec.
-IV.1, although they are not stabilized in the $T$$H$ phase diagram in this paper.
+7, we here examine the real space configuration of the symmetric quadruple-$\bm{q}$ orders discussed in Sec.
+IV.1, although they are not stabilized in the TTHH phase diagram in this paper.
 It is evident that the configurations with “collinear” types such as $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,\pi)$ are partially ordered states.
 For example, the quadrupole moments vanish at the sites where $\sum_{n}e^{i{\bm{k}}_{n}\cdot\bm{r}}=0$ for the former as shown in Fig.
-9(c).
+9(c); $\bm{\phi}^{A}(\bm{r})={\bf 0}$ for four fcc sites and $\bm{\phi}^{B}(\bm{r})=-\tfrac{1}{2}\bm{\phi}^{A}(\bm{0})$.
 As an example of “non-collinear” case, Fig.
 9(d) illustrates the configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$.
+This is also a partially ordered state, where $\bm{\phi}^{B}(\bm{r})=\bm{0}$ at two of the $B$ sublattice sites within the cubic unit cell shown in Fig.
+9(d), while the other two have finite moments $|\bm{\phi}^{B}(\bm{r})|=\sqrt{2}|\bm{\phi}^{A}(\bm{0})|$.
 Appendix B Renormalization effects due to the fields at the X points
@@
 They are more important than other modes at general wavenumbers since, for example, $\bm{q}_{1}+\bm{q}_{2}+\bm{q}_{\rm X}\equiv\bf{0}$ leads to cubic mode coupling with the two different fields at the L points.
-The free energy for the X points fields are sublattice diagonal and $F_{\rm X}=\sum_{\alpha=A,B}F_{\rm X}^{\alpha}$, where
-$F^{\alpha}_{\rm X}=$ $\frac{1}{2}\sum_{\ell={\rm X,Y,Z}}\bm{Q}^{\alpha}_{\ell}\begin{bmatrix}a_{\rm X}-6bu_{0}^{\alpha}&6bv_{0}^{\alpha}\\ 6bv_{0}^{\alpha}&a_{\rm X}+6bu_{0}^{\alpha}\end{bmatrix}\bm{Q}^{\alpha}_{\ell}$
+The free energy for the X-points fields are sublattice diagonal and ${\mathcal{F}}_{\rm X}=\sum_{\alpha=A,B}{\mathcal{F}}_{\rm X}^{\alpha}$, where
+${\mathcal{F}}^{\alpha}_{\rm X}=$ $\frac{1}{2}\sum_{\ell={\rm X,Y,Z}}\bm{Q}^{\alpha}_{\ell}\begin{bmatrix}a_{\rm X}-6bu_{0}^{\alpha}&6bv_{0}^{\alpha}\\ 6bv_{0}^{\alpha}&a_{\rm X}+6bu_{0}^{\alpha}\end{bmatrix}\bm{Q}^{\alpha}_{\ell}$
 $-6b\sum_{\ell={\rm X,Y,Z}}\sum_{i=1}^{2}\Big[Q_{\ell,z}^{\alpha}C_{\ell i}^{\alpha}-Q_{\ell,x}^{\alpha}S_{\ell i}^{\alpha}\Big]+\cdots.$ (32)
-Here, the fourth-order terms and the cubic term for three X points fields are neglected since the modes at the X points are not critical.
-$a_{\rm X}$ and $C$’s and $S$’s consisting of two L points variables are given as
+Here, the fourth-order terms and the cubic term for three X-point fields are neglected since the modes at the X points are not critical.
+$a_{\rm X}$ and $C$’s and $S$’s consisting of two L-point variables are given as
 $a_{\rm X}\equiv T-4J_{2}+6J_{4},$ (33)
@@
 where $(\ell,m,n)=({\rm X},1,2),\ ({\rm Y},1,4),\ ({\rm Z},1,3)$, and $(\ell,m^{\prime},n^{\prime})=({\rm X},3,4),\ ({\rm Y},2,3),\ ({\rm Z},2,4)$.
-By minimizing $F_{\rm X}$ with respect to the X points fields, we obtain their stationary values as
+By minimizing ${\mathcal{F}}_{\rm X}$ with respect to the X-point fields, we obtain their stationary values as
 $Q_{\ell,z}^{\alpha}=\frac{6b}{a_{\rm X}}\sum_{i=1}^{2}C^{\alpha}_{{\ell}i},\quad Q^{\alpha}_{\ell,x}=-\frac{6b}{a_{\rm X}}\sum_{i=1}^{2}S^{\alpha}_{{\ell}i}.$ (35)
-Substituting them into Eq. (32), we obtain up to the linear order in $\phi_{0}$: $\delta F=\sum_{\alpha={\rm A,B}}\delta F^{\alpha}_{0}+\delta F^{\alpha}_{1}+\cdots$ as
-$\delta F^{\alpha}_{0}=$ $-\frac{18b^{2}}{a_{\rm X}}\sum_{m>n}(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}$
+Substituting them into Eq.
+(32), we obtain up to the linear order in $\phi_{0}$: $\delta{\mathcal{F}}_{\rm{XL}}=\sum_{\alpha={\rm A,B}}\delta{\mathcal{F}}^{\alpha}_{\rm{XL}0}+\delta{\mathcal{F}}^{\alpha}_{\rm{XL}1}+\cdots$ as
+$\delta{\mathcal{F}}^{\alpha}_{\rm{XL}0}=$ $-\frac{18b^{2}}{a_{\rm X}}\sum_{m>n}(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}$
 $-\frac{36b^{2}}{a_{\rm X}}(c_{12}^{\alpha}c_{34}^{\alpha}+c_{13}^{\alpha}c_{24}^{\alpha}+c_{14}^{\alpha}c_{23}^{\alpha})\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4},$ (36)
 and
-$\delta F_{1}^{\alpha}=$ $-\frac{108b^{3}}{a_{\rm X}^{2}}\Bigg{[}\sum_{m>n}\cos(2\theta^{\alpha}_{mn}-\theta_{0}^{\alpha})(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}\phi_{0}^{\alpha}$
-$+6\cos(\bar{\theta}^{\alpha}-\bar{\theta}_{\alpha})\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4}\phi_{0}^{\alpha}\Bigg{]},$ (37)
+$\delta{\mathcal{F}}_{\rm{XL}1}^{\alpha}=$ $-\frac{108b^{3}}{a_{\rm X}^{2}}\Bigg[\sum_{m>n}\cos(2\theta^{\alpha}_{mn}-\theta_{0}^{\alpha})(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}\phi_{0}^{\alpha}$
+$+6\cos(\bar{\theta}^{\alpha}-\theta_{0}^{\alpha})\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4}\phi_{0}^{\alpha}\Bigg],$ (37)
 where $\bm{\phi}_{0}^{\alpha}=(u_{0}^{\alpha},v_{0}^{\alpha})=\phi^{\alpha}_{0}(\cos\theta^{\alpha}_{0},\sin\theta^{\alpha}_{0})$ and $\theta_{mn}\equiv\theta_{m}+\theta_{n}$ and $\bar{\theta}^{\alpha}\equiv\theta_{1}^{\alpha}+\theta_{2}^{\alpha}+\theta_{3}^{\alpha}+\theta_{4}^{\alpha}$.
@@
 $c^{\prime}\to c^{\prime}-\frac{72b^{2}}{a_{\rm X}},\quad c^{\prime\prime\prime}\to c^{\prime\prime\prime}-\frac{144b^{2}}{a_{\rm X}}.$ (38)
-The terms appearing in $\delta F_{1}^{\alpha}$ are emergent fifth-order coupling constants between the mode at the $\Gamma$ and the L points.
+The terms appearing in $\delta{\mathcal{F}}_{\rm{XL}1}^{\alpha}$ are emergent fifth-order coupling constants between the mode at the $\Gamma$ and the L points.
 Since we have kept terms up to the fourth orders in the original free energy in Eq.
 (6), cases for large $108b^{3}/a_{\rm X}$ are beyond the applicable range of the approximation.
+We also note that the last terms in Eqs.
+(36) and (37) vanish when the sublattice ($\alpha$) summations are carried outwith assuming $\bm{\phi}_{n}^{B}=(-1)^{p_{n}}\bm{\phi}_{n}^{A}$.
+See the discussions in Secs.
+III.1 and IV.1.
 Appendix C Free energy for the orthogonal double-q configuration
-We show the free energy for the orthogonal double-$\bm{q}$ orders with $\bm{\phi}_{2}\perp\bm{\phi}_{1}$: $\bm{\phi}_{1}=\phi_{1}(\cos\theta,\sin\theta),\bm{\phi}_{2}=\phi_{2}(\cos(\theta+\tfrac{\pi}{2}),\sin(\theta+\tfrac{\pi}{2}))=\phi_{2}(-\sin\theta,\cos\theta)$, with the $\Gamma$ point mode $\bm{\phi}_{0}=\phi_{0}(\cos\theta_{0},\sin\theta_{0})$.
+We show the free energy for the orthogonal double-$\bm{q}$ orders with $\bm{\phi}_{2}\perp\bm{\phi}_{1}$: $\bm{\phi}_{1}=\phi_{1}(\cos\theta,\sin\theta),\bm{\phi}_{2}=\phi_{2}(\cos(\theta+\tfrac{\pi}{2}),\sin(\theta+\tfrac{\pi}{2}))=\phi_{2}(-\sin\theta,\cos\theta)$, with the $\Gamma$-point mode $\bm{\phi}_{0}=\phi_{0}(\cos\theta_{0},\sin\theta_{0})$.
 Here, we assume $\bm{\phi}_{1}=\bm{\phi}_{1}^{A}=\bm{\phi}_{1}^{B}$ and $\bm{\phi}_{2}=\bm{\phi}_{2}^{A}=-\bm{\phi}_{2}^{B}$ as discussed in the main text.
 The free energy is given as
-$F=$ $\sum_{n=1,2}\Big{[}(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big{]}+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
+${\mathcal{F}}_{\rm or}=$ $\sum_{n=1,2}\Big[(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big]+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
 $-6\lambda_{b}\phi_{0}(\phi_{1}^{2}-\phi_{2}^{2})\cos(2\theta+\theta_{0})$
 $+2\lambda_{c}^{(2)}\phi_{0}^{2}\Big[\phi_{1}^{2}\cos^{2}(\theta-\theta_{0})+\phi_{2}^{2}\sin^{2}(\theta-\theta_{0})\Big]$
-$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos(4\theta-\theta_{0})+F_{\Gamma},$ (39)
-where $c^{\prime}=2c-72b^{2}/a_{\rm X},\ \lambda_{b}=b,\ \lambda_{c}^{(1)}=\lambda_{c}^{(2)}=c$, $F_{\Gamma}$ is given by Eq.
+$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos(4\theta-\theta_{0})+{\mathcal{F}}_{\Gamma},$ (39)
+where $c^{\prime}=2c-72b^{2}/a_{\rm X},\ \lambda_{b}=b,\ \lambda_{c}^{(1)}=\lambda_{c}^{(2)}=c$, and ${\mathcal{F}}_{\Gamma}$ is given by Eq.
 (10).
@@
 For simplicity, we assume $\theta_{0}=0$, since $\tilde{h}$ induces $u_{0}$ under both $\bm{H}\parallel[001]$ and $[110]$.
-We treat $\phi_{0,1,2}$ as given parameters and the stationary condition of $F$ with respect to $\cos(2\theta)$ leads to
+We treat $\phi_{0,1,2}$ as given parameters and the stationary condition of ${\mathcal{F}}_{\rm or}$ with respect to $\cos(2\theta)$ leads to
 $\cos(2\theta)=\frac{6\lambda_{b}-\lambda_{c}^{(2)}\phi_{0}}{216b^{3}/a_{\rm X}^{3}}\left(\frac{\phi_{1}^{2}-\phi_{2}^{2}}{\phi_{1}^{2}\phi_{2}^{2}}\right).$ (40)
@@
 The free energy for the phase III2q is
-$F_{{\rm III}_{2q}}=$ $2\big{[}a_{\rm L}+J_{\rm L}+(\lambda_{c}^{(1)}+\lambda_{c}^{(2)})\phi_{0}^{2}\big{]}\phi_{1}^{2}$
-$+\left(c+\frac{c^{\prime}}{2}-\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\cos\theta_{0}\right)\phi_{1}^{4}+F_{\Gamma}.$ (41)
-Although we will not discuss in detail, we can also see that the deviation $\delta\phi=\phi_{1}-\phi_{2}\neq 0$ from the phase III2q accompanies changes in the angle as $\theta=\pi/4+\epsilon$.
-The quadratic form of $\delta\phi$ and $\epsilon$ describe the stability of the phase III2q and when one of the eigenvalues vanishes a second-order transition takes place toward the phase II2q.
+${\mathcal{F}}_{{\rm III}_{2q}}=$ $2\big[a_{\rm L}+J_{\rm L}+(\lambda_{c}^{(1)}+\lambda_{c}^{(2)})\phi_{0}^{2}\big]\phi_{1}^{2}$
+$+\left(c+\frac{c^{\prime}}{2}-\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\cos\theta_{0}\right)\phi_{1}^{4}+{\mathcal{F}}_{\Gamma}.$ (41)
+Although we do not discuss this in detail, we can also see that the deviation $\delta\phi=\phi_{1}-\phi_{2}\neq 0$ from the phase III2q accompanies changes in the angle as $\theta=\pi/4+\epsilon$.
+The quadratic form of $\delta\phi$ and $\epsilon$ describes the stability of the phase III2q.
+When one of the eigenvalues vanishes, a second-order transition takes place toward the phase II2q.
 For $\theta=0$, i.e., the phase I2q, the free energy is given by
-$F_{\rm{I}_{2q}}=$ $\sum_{n=1,2}\Big{[}(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big{]}+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
+${\mathcal{F}}_{{\rm{I}}_{2q}}=$ $\sum_{n=1,2}\Big[(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big]+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
 $-6\lambda_{b}\phi_{0}(\phi_{1}^{2}-\phi_{2}^{2})\cos\theta_{0}$
 $+2\lambda_{c}^{(2)}\phi_{0}^{2}\Big[\phi_{1}^{2}\cos^{2}\theta_{0}+\phi_{2}^{2}\sin^{2}\theta_{0}\Big]$
-$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos\theta_{0}+F_{\Gamma}.$ (42)
+$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos\theta_{0}+{\mathcal{F}}_{\Gamma}.$ (42)
 Rotating $\theta$ leads to the transition to the phase II2q. This accompanies also rotation in $\theta_{0}$ accordingly.